### v1

# All genus open mirror symmetry for the projective line

Abstract.
We prove that the Chekhov-Eynard-Orantin recursion on the mirror curve of $\mathbb{P}^{1}$ encodes all genus equivariant open Gromov-Witten invariants of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$. This result can be viewed as an all genus equivariant open mirror symmetry for $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$.
1. Introduction
1.1. Historical background and motivation
1.1.1. All genus closed mirror symmetry for the projective line
Mirror symmetry is a duality from string theory originally discovered by physicists. It says two dual string theories – type IIA and type IIB – on different Calabi-Yau 3-folds give rise to the same physics. Mathematicians became interested in this relationship around 1990 when Candelas, de la Ossa, Green, and Parkes [7] obtained a conjectural formula of the number of rational curves of arbitrary degree in the quintic 3-fold by relating it to period integrals of the quintic mirror.
By late 1990s mathematicians had established the foundation of Gromov-Witten (GW) theory as a mathematical theory of A-model topological closed strings. In this context, the genus $g$ free energy of the topological A-model is defined as a generating function of genus $g$ Gromov-Witten invariants. The mathematical aspect of genus zero closed mirror symmetry has been well-studied in many cases. Givental [22] and Lian-Liu-Yau [29] independently proved the genus zero mirror formula for the quintic Calabi-Yau 3-fold $Q$; later they extended their results to Calabi-Yau complete intersections in projective toric manifolds [23, 30, 31]. The genus-zero mirror theorem for toric Deligne-Mumford stacks is proved in [13, 11].
The higher genus mirror symmetry is much more subtle. Bershadsky-Cecotti-Ooguri-Vafa (BCOV) conjectured the genus-one and genus-two mirror formulae for the quintic 3-fold [1]. Combining the techniques of BCOV, results of Yamaguchi-Yau [35], and boundary conditions, Huang-Klemm-Quackenbush [27] proposed a mirror conjecture on $F_{g}^{Q}$ up to $g=51$. The BCOV genus-one mirror formula was first proved by A. Zinger in [40] using genus-one reduced Gromov-Witten theory, and later reproved in [28, 12] via quasimap theory and in [9] via MSP theory. The BCOV genus-two mirror formula was proved by Guo-Janda-Ruan [25] and Chang-Guo-Li [8]. The Yamaguchi–Yau’s finite generation and the holomorphic anomaly equation for quintic 3-folds were proved in [8] and [26].
The all genus closed mirror symmetry for the projective line is easier to study and there is a nice mathematical theory for the higher genus B-model. P. Dunin-Barkowski, N. Orantin, S. Shadrin and L. Spitz [14] relate the Chekhov-Eynard-Orantin topological recursion to the higher genus closed Gromov-Witten invariants of $\mathbb{P}^{1}$, proving the Norbury-Scott conjecture [34]. In [19], the above result is generalized to the equivariant setting. The main result of [19] relates the higher genus equivariant descendant Gromov-Witten potentials of $\mathbb{P}^{1}$ to the oscillatory integrals of Chekhov-Eynard-Orantin invariants of the mirror curve. This result can be viewed as an all genus equivariant closed mirror symmetry for $\mathbb{P}^{1}$. By taking the nonequivariant limit, the Norbury-Scott conjecture is recovered.
1.1.2. Open Gromov-Witten theory for the projective line
The open Gromov-Witten invariants are more difficult to study than the closed case. The all genus $S^{1}$-equivariant open Gromov-Witten theory of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ is studied by Buryak-Netser Zernik-Pandharipande-Tessler in [6]. A virtual localization formula is conjectured and proved in the genus zero case. The localization formula is crucial for the computation of open Gromov-Witten theory of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$. It is natural to ask whether one can extend the result in [19] to the open string sector. This will be introduced in Section 1.1.3.
1.1.3. All genus open mirror symmetry for the projective line
In this paper, we prove the all genus open mirror symmetry for $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$, extending the result in [19] to the open string sector. On A-model side, we consider the generating function of higher genus open Gromov-Witten invariants of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$. On B-model side, instead of taking the oscillatory integrals, we take the expansion of the Chekhov-Eynard-Orantin invariants under certain local coordinate at the puncture of the mirror curve. The main theorem (Theorem 1.1) of this paper states that this expansion coincides with the generating function of higher genus open Gromov-Witten invariants of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ under the mirror map.
We would like to remark that similar open mirror symmetry phenomenon appears in the case of toric Calabi-Yau 3-folds/3-orbifolds. Based on the work of Eynard-Orantin [15] and Mariño [33], Bouchard-Klemm-Mariño-Pasquetti (BKMP) [4, 5] proposed a new formalism of the topological B-model in terms of the Chekhov-Eynard-Orantin invariants of the mirror curve. BKMP conjectured a precise correspondence, known as the BKMP Remodeling Conjecture, between the local expansion of the Chekhov-Eynard-Orantin invariants at the puncture of the mirror curve and the generating function of open Gromov-Witten invariants of toric Calabi-Yau 3-folds/3-orbifolds. The mathematical study of the Remodeling Conjecture can be found, for example, in [3, 10, 16, 17, 18, 20, 21, 36, 37, 38, 39]. The difference in our case is that our target space $\mathbb{P}^{1}$ is compact and non-Calabi-Yau.
We emphasize the following feature of our main result. Since $\mathbb{P}^{1}$ and $\mathbb{R}\mathbb{P}^{1}$ are compact, the non-equivariant limit of the $S^{1}$-equivariant open Gromov-Witten invariants of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ is conjectured [6, Conjecture 1] to be equal to the non-equivariant open Gromov-Witten invariants of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$. These non-equivariant open Gromov-Witten invariants are conjectured to be defined geometrically via coherent boundary condition. In general, the computation of higher genus open Gromov-Witten invariants is quite difficult. The main result of our paper provides an effective algorithm of computing higher genus open Gromov-Witten invariants of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ recursively.
1.2. Statement of the main result
In this paper, we denote the complex projective line by $\mathbb{P}^{1}$. Let $t\in T=S^{1}$ act on $\mathbb{P}^{1}$ by
$t\cdot[z_{1},z_{2}]=[tz_{1},t^{-1}z_{2}].$
Let $\mathbb{C}[\mathsf{v}]=H_{T}^{*}(\mathrm{point};\mathbb{C})$ be the $T$-equivariant cohomology of a point. Let $p_{1}=[1,0]$ and $p_{2}=[0,1]$ be the $T$ fixed points. The $T$-equivariant cohomology of $\mathbb{P}^{1}$ is given by
$H^{*}_{T}(\mathbb{P}^{1};\mathbb{C})=\mathbb{C}[H,\mathsf{v}]/\langle(H+\mathsf{v}/2)(H-\mathsf{v}/2)\rangle$
where $\deg H=\deg\mathsf{v}=2$. Let
$L:=\{[e^{{\rm i}\varphi},e^{-{\rm i}\varphi}]\in\mathbb{P}^{1}:\varphi\in\mathbb{R}\}$
be the Lagrangian submanifold of $\mathbb{P}^{1}$, which is preserved by the $T$-action. By taking a Möbius transform, we can identify the pair $(\mathbb{P}^{1},L)$ with $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$.
In Section 3.5, we will define the generating function $F_{g,h}(\mathbf{t},Q;X_{1},\dots,X_{h})$ of genus $g$, $h$ boundary circles open Gromov-Witten invariants of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$. Here $\mathbf{t}=t^{0}1+t^{1}H$, $Q$ is the Novikov variable encoding the degree of the stable maps to $\mathbb{P}^{1}$, and $X_{1},\cdots,X_{h}$ are variables encoding the winding numbers (viewed as the open string coordinates).
In Section 4.1, we will study the mirror curve of $\mathbb{P}^{1}$. Consider the $T$-equivariant superpotential $W:\mathbb{C}^{*}\rightarrow\mathbb{C}$ defined as
$W(Y)=t^{0}+\frac{\mathsf{v}}{2}\log q+Y+\frac{q}{Y}-\mathsf{v}\log Y,$
where $q=e^{t^{1}}$. Let $X=e^{-x/\mathsf{v}}$, $Y=e^{y}$. Consider the mirror curve $C_{q}$ defined as follows:
$C_{q}=\{(x,y)\in\mathbb{C}^{2}:x=W(e^{y})\}.$
In Section 4.3, we will study the Chekhov-Eynard-Orantin invariants $\omega_{g,n}$ of the mirror curve $C_{q}$. Then we use $\omega_{g,n}$ to define the B-model potential $W_{g,n}(t^{0},q;X_{1},\dots,X_{n})$.
The following theorem is the main result of this paper:
Theorem 1.1 (= Theorem 5.6).
For any $n>0$ and $g\geq 0$, we have
$F_{g,n}(\mathbf{t},1;X_{1},\dots,X_{n})=(-1)^{g-1}W_{g,n}(t^{0},q;X_{1},\dots,X_{n}).$
1.3. Overview of the paper
In Section 2, we study the equivariant closed Gromov-Witten theory of $\mathbb{P}^{1}$ and give the graph sum formula for the descendant Gromov-Witten potential. In Section 3, we study the equivariant open Gromov-Witten theory of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ and define the A-model potential $F_{g,n}$. In Section 4, we study the mirror curve of $\mathbb{P}^{1}$ and the Chekhov-Eynard-Orantin topological recursion. Then we use the Chekhov-Eynard-Orantin invariants $\omega_{g,n}$ to define the B-model potential $W_{g,n}$. In Section 5, we prove the all genus open mirror symmetry for $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$, which is the main result of this paper.
Acknowledgements
The authors would like to thank Bohan Fang, Chiu-Chu Melissa Liu, and Song Yu for useful discussions. The second author is partially supported by the Natural Science Foundation of Beijing, China grant No. 1252008.
2. Equivariant closed Gromov-Witten theory of $\mathbb{P}^{1}$
2.1. Equivariant cohomology of $\mathbb{P}^{1}$
Let $t\in T=S^{1}$ act on $\mathbb{P}^{1}$ by
$t\cdot[z_{1},z_{2}]=[tz_{1},t^{-1}z_{2}].$
The $T$-equivariant cohomology of $\mathbb{P}^{1}$ is given by
$H^{*}_{T}(\mathbb{P}^{1};\mathbb{C})=\mathbb{C}[H,\mathsf{v}]/\langle(H+\mathsf{v}/2)(H-\mathsf{v}/2)\rangle.$
Let $p_{1}=[1,0]$ and $p_{2}=[0,1]$ be the $T$-fixed points. Then $H|_{p_{1}}=-\mathsf{v}/2$, $H|_{p_{2}}=\mathsf{v}/2$. The $T$-equivariant Poincaré dual of $p_{1}$ and $p_{2}$ are $H-\mathsf{v}/2$ and $H+\mathsf{v}/2$, respectively.
Let
$\phi_{1}:=-\frac{H-\mathsf{v}/2}{\mathsf{v}},\phi_{2}:=\frac{H+\mathsf{v}/2}{\mathsf{v}}\in H^{*}_{T}(\mathbb{P}^{1};\mathbb{C})\otimes_{\mathbb{C}[\mathsf{v}]}\mathbb{C}[\mathsf{v},\frac{1}{\mathsf{v}}].$
We have
$\phi_{1}+\phi_{2}=1,\quad\phi_{\alpha}\cup\phi_{\beta}=\delta_{\alpha\beta}\phi_{\alpha},$
$(\phi_{\alpha},\phi_{\beta})_{\mathbb{P}^{1},T}:=\int_{\mathbb{P}^{1}}\phi_{\alpha}\cup\phi_{\beta}=\delta_{\alpha\beta}\int_{\mathbb{P}^{1}}\phi_{\alpha}=\frac{\delta_{\alpha\beta}}{\Delta^{\alpha}},\quad\alpha,\beta\in\{1,2\},$
where
$\Delta^{1}=-\mathsf{v},\quad\Delta^{2}=\mathsf{v}.$
For convenience, we introduce the notations $R_{T}:=\mathbb{C}[\mathsf{v}]$ and ${S}_{T}:=\mathbb{C}[\mathsf{v},\frac{1}{\mathsf{v}}]$, and we let $\bar{S}_{T}$ be the minimal field extension of $S_{T}$ containing $\{\sqrt{\Delta^{\alpha}}:\alpha=1,2\}$. Then $\{\phi_{\alpha}:\alpha=1,2\}$ is a canonical basis of the semisimple Frobenius algebra
$(H^{*}_{T}(\mathbb{P}^{1};\mathbb{C})\otimes_{R_{T}}\bar{S}_{T},\cup,(\cdot,\cdot)_{\mathbb{P}^{1},T})$
over $\bar{S}_{T}$. Let
$\hat{\phi}_{\alpha}:=\sqrt{\Delta^{\alpha}}\phi_{\alpha}.$
Then $\{\hat{\phi}_{\alpha}:\alpha=1,2\}$ is the normalized canonical basis of the semisimple Frobenius algebra $H^{*}_{T}(\mathbb{P}^{1};\mathbb{C})\otimes_{R_{T}}\bar{S}_{T}$:
$\hat{\phi}_{\alpha}\cup\hat{\phi}_{\beta}=\delta_{\alpha\beta}\sqrt{\Delta^{\alpha}}\hat{\phi}_{\alpha},\quad(\hat{\phi}_{\alpha},\hat{\phi}_{\beta})_{\mathbb{P}^{1},T}=\delta_{\alpha\beta}.$
2.2. Equivariant Gromov-Witten invariants of $\mathbb{P}^{1}$
Let $E(\mathbb{P}^{1})$ denote the effective curve classes in $H_{2}(\mathbb{P}^{1};\mathbb{Z})$. Given nonnegative integers $g,n$ and an effective curve class $\beta\in E(\mathbb{P}^{1})$, let ¯$\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{1},\beta)$ be the moduli stack of genus $g$, $n$-pointed, degree $\beta$ stable maps to $\mathbb{P}^{1}$. Let $\mathrm{ev}_{i}:\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{1},\beta)\rightarrow\mathbb{P}^{1}$ be the evaluation map at the $i$-th marked point. The $T$-action on $\mathbb{P}^{1}$ induces an $T$-action on ¯$\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{1},\beta)$ and the evaluation map $\mathrm{ev}_{i}$ is $T$-equivariant.
For $i=1,\dots,n$, let $\mathbb{L}_{i}$ be the $i$-th tautological line bundle over ¯$\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{1},\beta)$ formed by the cotangent line at the $i$-th marked point. Define the $i$-th descendant class $\psi_{i}$ as
$\psi_{i}:=c_{1}(\mathbb{L}_{i})\in H^{2}(\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{1},\beta);\mathbb{Q}).$
Given $\gamma_{1},\dots,\gamma_{n}\in H^{*}_{T}(\mathbb{P}^{1};\mathbb{C})$ and nonnegative integers $a_{1},\dots,a_{n}$, we define genus-$g$, degree-$\beta$, $T$-equivariant descendant Gromov-Witten invariants of $\mathbb{P}^{1}$:
$\langle\tau_{a_{1}}(\gamma_{1})\dots\tau_{a_{n}}(\gamma_{n})\rangle_{g,n,\beta}^{\mathbb{P}^{1},T}:=\int_{[\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{1},\beta)]^{\mathrm{vir}}}\prod_{i=1}^{n}\psi_{i}^{a_{i}}\mathrm{ev}_{i}^{*}(\gamma_{i})\in\mathbb{C}[\mathsf{v}].$
The genus-$g$, degree-$\beta$, $T$-equivariant primary Gromov-Witten invariants of $\mathbb{P}^{1}$ is defined as
$\langle\gamma_{1}\dots\gamma_{n}\rangle_{g,n,\beta}^{\mathbb{P}^{1},T}:=\langle\tau_{0}(\gamma_{1})\dots\tau_{0}(\gamma_{n})\rangle_{g,n,\beta}^{\mathbb{P}^{1},T}.$
Define the Novikov ring
$\Lambda_{\mathrm{nov}}:=\mathbb{C}[\widehat{E(\mathbb{P}^{1})}]=\Bigg{\{}\sum_{\beta\in E(\mathbb{P}^{1})}c_{\beta}Q^{\beta}:c_{\beta}\in\mathbb{C}\Bigg{\}},$
where $Q$ is the Novikov variable. Let $\mathbf{t}=t^{0}1+t^{1}H$, we define the following double correlator:
$\tau_{a_{1}}(\gamma_{1}),\dots,\tau_{a_{n}}(\gamma_{n})\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}_{g,n}^{\mathbb{P}^{1},T}:=\sum_{\beta\in E(\mathbb{P}^{1})}\sum_{m=0}^{\infty}\frac{Q^{\beta}}{m!}\langle\tau_{a_{1}}(\gamma_{1}),\dots,\tau_{a_{n}}(\gamma_{n}),\mathbf{t}^{m}\rangle_{g,n+m,\beta}^{\mathbb{P}^{1},T}.⟨$
For $j=1,\dots,n$, introduce formal variables
${\bf{u}}_{j}={\bf{u}}_{j}(z)=\sum_{a\geq 0}(u_{j})_{a}z^{a}$
where $(u_{j})_{a}\in H^{*}_{T}(\mathbb{P}^{1})\otimes_{R_{T}}\bar{S}_{T}$. Define
${\bf u}_{1},\dots,{\bf u}_{n}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}_{g,n}^{\mathbb{P}^{1},T}=\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}{\bf u}_{1}(\psi),\dots,{\bf u}_{n}(\psi)\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}_{g,n}^{\mathbb{P}^{1},T}=\sum_{a_{1},\dots,a_{n}\geq 0}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}(u_{1})_{a_{1}}\psi^{a_{1}},\dots,(u_{n})_{a_{n}}\psi^{a_{n}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}_{g,n}^{\mathbb{P}^{1},T}.⟨$
Let $z_{1},\dots,z_{n}$ be formal variables and $\gamma_{1},\dots,\gamma_{n}\in H^{*}_{T}(\mathbb{P}^{1})\otimes_{R_{T}}\bar{S}_{T}$. Define
$\frac{\gamma_{1}}{z_{1}-\psi},\dots,\frac{\gamma_{n}}{z_{n}-\psi}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}_{g,n}^{\mathbb{P}^{1},T}=\sum_{a_{1},\dots,a_{n}\in\mathbb{Z}_{\geq 0}}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\gamma_{1}\psi^{a_{1}},\dots,\gamma_{n}\psi^{a_{n}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}_{g,n}^{\mathbb{P}^{1},T}\prod_{i=1}^{n}z_{i}^{-a_{i}-1}.⟨$
We use the conventions that
$\langle\frac{\gamma}{z-\psi}\rangle_{0,1,0}^{\mathbb{P}^{1},T}:=z\int_{\mathbb{P}^{1}}\gamma,$
$\langle\frac{\gamma_{1}}{z-\psi},\gamma_{2}\rangle_{0,2,0}^{\mathbb{P}^{1},T}:=\int_{\mathbb{P}^{1}}\gamma_{1}\cup\gamma_{2},$
$\langle\frac{\gamma_{1}}{z_{1}-\psi_{1}},\frac{\gamma_{2}}{z_{2}-\psi_{2}}\rangle_{0,2,0}^{\mathbb{P}^{1},T}:=\frac{1}{z_{1}+z_{2}}\int_{\mathbb{P}^{1}}\gamma_{1}\cup\gamma_{2}.$
2.3. Equivariant quantum cohomology and Frobenius structure
The $T$-equivariant quantum cohomology ring $QH_{T}^{*}(\mathbb{P}^{1})$ of $\mathbb{P}^{1}$ is defined by its genus-zero primary Gromov-Witten invariants. As a $\mathbb{C}[\mathsf{v}]$-module, $QH_{T}^{*}(\mathbb{P}^{1})=H^{*}_{T}(\mathbb{P}^{1})$. The ring structure is given by the quantum product $\star$:
$(\gamma_{1}\star\gamma_{2},\gamma_{3})_{\mathbb{P}^{1},T}=\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\gamma_{1},\gamma_{2},\gamma_{3}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}_{0,3}^{\mathbb{P}^{1},T}.$
In other words,
$\gamma_{1}\star\gamma_{2}=\gamma_{1}\cup\gamma_{2}+q\left(\int_{\mathbb{P}^{1}}\gamma_{1}\right)\left(\int_{\mathbb{P}^{1}}\gamma_{2}\right),$
where $\cup$ is the cup product in $H^{*}_{T}(\mathbb{P}^{1})$ and $q=Qe^{t^{1}}$. In particular,
$H\star H=\mathsf{v}^{2}/4+q.$
The $T$-equivariant quantum cohomology ring of $\mathbb{P}^{1}$ is given by
$QH^{*}_{T}(\mathbb{P}^{1};\mathbb{C})=\mathbb{C}[H,\mathsf{v},q]/\langle(H+\mathsf{v}/2)\star(H-\mathsf{v}/2)-q\rangle,$
where $\deg H=\deg\mathsf{v}=2$ and $\deg q=4$.
Let
$\phi_{1}(q)=\frac{1}{2}-\frac{H}{\mathsf{v}\sqrt{1+4q/\mathsf{v}^{2}}},\quad\phi_{2}(q)=\frac{1}{2}+\frac{H}{\mathsf{v}\sqrt{1+4q/\mathsf{v}^{2}}}.$
Then
$\phi_{\alpha}(q)\star\phi_{\beta}(q)=\delta_{\alpha\beta}\phi_{\alpha}(q),\quad(\phi_{\alpha}(q),\phi_{\beta}(q))_{\mathbb{P}^{1},T}=\frac{\delta_{\alpha\beta}}{\Delta^{\alpha}(q)},$
where
$\Delta^{1}(q)=-\mathsf{v}\sqrt{1+4q/\mathsf{v}^{2}},\quad\Delta^{2}(q)=-\Delta^{1}(q).$
Let $S:=\bar{S}_{T}[[q]]$. Then $\{\phi_{1}(q),\phi_{2}(q)\}$ is a canonical basis of the semisimple Frobenius algebra
$(H^{*}_{T}(\mathbb{P}^{1};\mathbb{C})\otimes_{R_{T}}S,\star,(\cdot,\cdot)_{\mathbb{P}^{1},T}).$
The normalized canonical basis is defined by
$\{\hat{\phi}_{\alpha}(q):=\sqrt{\Delta^{\alpha}(q)}\phi_{\alpha}(q):\alpha=1,2\}.$
They satisfy
$\hat{\phi}_{\alpha}(q)\star\hat{\phi}_{\beta}(q)=\delta_{\alpha\beta}\sqrt{\Delta^{\alpha}(q)}\hat{\phi}_{\alpha}(q),\quad(\hat{\phi}_{\alpha}(q),\hat{\phi}_{\beta}(q))_{\mathbb{P}^{1},T}=\delta_{\alpha\beta}.$
2.4. Canonical coordinates and the transition matrix
Let $\{t^{0},t^{1}\}$ be the flat coordinates with respect to the basis $\{T_{0}=1,T_{1}=H\}$, and let $\{u^{1},u^{2}\}$ be the canonical coordinates with respect to the basis $\{\phi_{1}(q),\phi_{2}(q)\}$. Then
$\frac{\partial}{\partial u^{1}}$ $=\frac{1}{2}\frac{\partial}{\partial t^{0}}+\frac{1}{\Delta^{1}(q)}\frac{\partial}{\partial t^{1}},$
$\frac{\partial}{\partial u^{2}}$ $=\frac{1}{2}\frac{\partial}{\partial t^{0}}+\frac{1}{\Delta^{2}(q)}\frac{\partial}{\partial t^{1}}.$
The canonical coordinates $u^{1}$ and $u^{2}$ are characterized by the above equations up to a constant in $\mathbb{C}[\mathsf{v},\frac{1}{\mathsf{v}}]$. We choose the canonical coordinates $(u^{1},u^{2})$ such that
$\lim_{q\rightarrow 0}(u^{1}-t^{0}+\mathsf{v}t^{1}/2)=0,\quad\lim_{q\rightarrow 0}(u^{2}-t^{0}-\mathsf{v}t^{1}/2)=0.$
We use Greek alphabet $\alpha\in\{1,2\}$ to represent canonical coordinates and the Latin alphabet $i\in\{0,1\}$ to represent flat coordinates. Let $\Psi=(\Psi_{i}^{\ \alpha})$ be the transition matrix between the flat and quantum normalized canonical bases:
$T_{i}=\sum_{\alpha\in\{1,2\}}\Psi_{i}^{\ \alpha}\hat{\phi}_{\alpha}(q).$
It can be rewritten as
$\frac{\partial}{\partial t^{i}}=\sum_{\alpha\in\{1,2\}}\Psi_{i}^{\ \alpha}\sqrt{\Delta^{\alpha}(q)}\frac{\partial}{\partial u^{\alpha}},$
which is equivalent to
$\frac{du^{\alpha}}{\sqrt{\Delta^{\alpha}(q)}}=\sum_{i\in\{0,1\}}dt^{i}\Psi_{i}^{\ \alpha}.$
Then
$\Psi_{0}^{\ \alpha}=\frac{1}{\sqrt{\Delta^{\alpha}(q)}},\quad\Psi_{1}^{\ \alpha}=\frac{\sqrt{\Delta^{\alpha}(q)}}{2}.$
Let $\Psi^{-1}=(\Psi^{-1})_{\alpha}^{\ i}$, so that
$\sum_{i\in\{0,1\}}(\Psi^{-1})_{\alpha}^{\ i}\Psi_{i}^{\ \beta}=\delta_{\alpha}^{\ \beta}.$
Then
[…]
2.5. The $\mathcal{S}$-operator
For any $a,b\in H^{*}_{T}(\mathbb{P}^{1};\bar{S}_{T})$, define the $\mathcal{S}$-operator by
$(a,\mathcal{S}(b))_{\mathbb{P}^{1},T}=(a,b)_{\mathbb{P}^{1},T}+\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}a,\frac{b}{z-\psi}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}_{0,2}^{\mathbb{P}^{1},T}.$
We consider several different (flat) bases for $H^{*}_{T}(\mathbb{P}^{1};\bar{S}_{T})$:
  1. (1)
    The natural basis: $T_{0}=1$ and $T_{1}=H$.
  2. (2)
    The basis dual to the natural basis: $T^{0}=H$ and $T^{1}=1$.
  3. (3)
    The classicial canonical basis $\{\phi_{1},\phi_{2}\}$.
  4. (4)
    The basis dual to the classicial canonical basis with respect to the $T$-equivariant Poincaré pairing: $\phi^{1}=\Delta^{1}\phi_{1}$, $\phi^{2}=\Delta^{2}\phi_{2}$.
  5. (5)
    The normalized classical basis $\hat{\phi}_{1}=\sqrt{\Delta^{1}}\phi_{1}$ and $\hat{\phi}_{2}=\sqrt{\Delta^{2}}\phi_{2}$, which is self-dual: $\hat{\phi}^{1}=\hat{\phi}_{1}$ and $\hat{\phi}^{2}=\hat{\phi}_{2}$.
We also consider several different bases for $H^{*}_{T}(\mathbb{P}^{1};\bar{S}_{T})\otimes_{\bar{S}_{T}}S$:
  1. (1)
    The quantum classical basis $\{\phi_{1}(q),\phi_{2}(q)\}$.
  2. (2)
    The basis dual to the quantun canonical basis with respect to the $T$-equivariant Poincaré pairing: $\phi^{1}(q)=\Delta^{1}(q)\phi_{1}(q)$, $\phi^{2}(q)=\Delta^{2}(q)\phi_{2}(q)$.
  3. (3)
    The normalized quantum classicial basis $\{\hat{\phi}_{1}(q),\hat{\phi}_{2}(q)\}$, which is self-dual.
We introduce some additional notations. For $\alpha,\beta\in\{1,2\}$, define
$S^{\alpha}_{\ \beta}=(\phi^{\alpha},\mathcal{S}(\phi_{\beta}))_{\mathbb{P}^{1},T},\quad S^{\hat{\underline{\alpha}}}_{\ \beta}=(\hat{\phi}_{\alpha}(q),\mathcal{S}(\phi_{\beta}))_{\mathbb{P}^{1},T},\quad S^{\hat{\underline{\alpha}}}_{\ \hat{\underline{\beta}}}=(\hat{\phi}_{\alpha}(q),\mathcal{S}(\hat{\phi}_{\beta}(q)))_{\mathbb{P}^{1},T}.$
For $\alpha,\beta\in\{1,2\}$ and $i\in\{0,1\}$, define
$S_{i}^{\ \hat{\alpha}}=(T_{i},\mathcal{S}(\hat{\phi}^{\alpha}))_{\mathbb{P}^{1},T}.$
For $a,b\in H^{*}_{T}(\mathbb{P}^{1};\bar{S}_{T})$, define
$S_{z}(a,b)$ $=(a,\mathcal{S}(b))_{\mathbb{P}^{1},T}$
$V_{z_{1},z_{2}}(a,b)$ $=\frac{(a,b)_{\mathbb{P}^{1},T}}{z_{1}+z_{2}}+\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\frac{a}{z_{1}-\psi_{1}},\frac{b}{z_{2}-\psi_{2}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}_{0,2}^{\mathbb{P}^{1},T}.$
We have the following identity:
$V_{z_{1},z_{2}}(a,b)=\frac{1}{z_{1}+z_{2}}\sum_{\alpha\in\{1,2\}}S_{z_{1}}(\hat{\phi}_{\alpha}(q),a)S_{z_{2}}(\hat{\phi}_{\alpha}(q),b).$
Remark 2.1.
Let $t^{\prime}=t^{1}H$ and $t^{\prime\prime}=t^{0}1$. By the divisor equation, we have
$\ (a,b)_{\mathbb{P}^{1},T}+\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}a,\frac{b}{z-\psi}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}_{0,2}^{\mathbb{P}^{1},T}$
$=$ $\ (a,be^{t^{\prime}/z})_{\mathbb{P}^{1},T}+\sum_{m=0}^{\infty}\sum_{\beta\in E(\mathbb{P}^{1})\atop{(\beta,m)\neq(0,0)}}\frac{Q^{\beta}e^{\int_{\beta}t^{\prime}}}{m!}\langle a,\frac{be^{t^{\prime}/z}}{z-\psi},(t^{\prime\prime})^{m}\rangle^{\mathbb{P}^{1},T}_{0,2+m,\beta}.$
The factor $e^{\int_{\beta}t^{\prime}}$ plays the role of $Q^{\beta}$, so the operator $\mathcal{S}(b)\big{|}_{Q=1}$ is well-defined.
2.6. Equivariant $J$-function
The $T$-equivariant $J$-function $J(z)$ is characterized by
$(J(z),a)_{\mathbb{P}^{1},T}=(1,\mathcal{S}(a)\big{|}_{Q=1})_{\mathbb{P}^{1},T},$
for any $a\in H^{*}_{T}(\mathbb{P}^{1};\bar{S}_{T})$. Equivalently,
$J(z)=1+\sum_{\alpha\in\{1,2\}}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}1,\frac{\hat{\phi}_{\alpha}}{z-\psi}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}_{0,2}^{\mathbb{P}^{1},T}\Big{|}_{Q=1}\hat{\phi}_{\alpha}.$
By the genus zero mirror theorem [22, 29],
$J(z)=e^{(t^{0}+t^{1}H)/z}\Bigg{(}1+\sum_{d=1}^{\infty}\frac{q^{d}}{\prod_{m=1}^{d}(H+\mathsf{v}/2+mz)\prod_{m=1}^{d}(H-\mathsf{v}/2+mz)}\Bigg{)},$
where $q=e^{t^{1}}$.
Let $J(z)=J^{1}\phi_{1}+J^{2}\phi_{2}$. Then for $\alpha=1,2$, we have
(1) $J^{\alpha}$ $=e^{(t^{0}+t^{1}\Delta^{\alpha}/2)/z}\sum_{d=0}^{\infty}\frac{q^{d}}{d!z^{d}}\frac{1}{\prod_{m=1}^{d}(\Delta^{\alpha}+mz)}$
$=e^{(t^{0}+t^{1}\Delta^{\alpha}/2)/z}\sum_{m=0}^{\infty}\Big{(}\frac{\sqrt{q}}{z}\Big{)}^{2m}\frac{\Gamma(\Delta^{\alpha}/z+1)}{m!\Gamma(\Delta^{\alpha}/z+m+1)}$
$=e^{t^{0}/z}z^{\Delta^{\alpha}/z}\Gamma(\Delta^{\alpha}/z+1)I_{\Delta^{\alpha}/z}\Big{(}\frac{2\sqrt{q}}{z}\Big{)},$
where the function $I_{\alpha}(x)$ is the modified Bessel function of first kind in Appendix A.
2.7. The $R$-matrix
Let $U$ denote the diagnoal matrix $\text{diag}(u^{1},u^{2})$. The results in [24] imply the following statement.
Theorem 2.2.
There exists a unique matrix power series $R(z)=\mathbf{1}+R_{1}z+R_{2}z^{2}+\dots$ satisfying the following properties:
  1. (1)
    The entries of each $R_{k}$ lie in $\bar{S}_{T}[[q,t^{0}]]$.
  2. (2)
    $\tilde{S}=\Psi R(z)e^{U/z}$ is a fundamental solution to the $T$-equivariant big quantum differential equation.
  3. (3)
    $R$ satisfies the unitary condition $R^{T}(-z)R(z)=\mathbf{1}$.
  4. (4)
    For $\alpha,\beta\in\{1,2\}$, we have
    $\lim_{q,t^{0}\rightarrow 0}R_{\alpha}^{\ \beta}(z)=\delta_{\alpha\beta}\exp\left(-\sum_{n=1}^{\infty}\frac{B_{2n}}{2n(2n-1)}\left(\frac{z}{\Delta^{\beta}}\right)^{2n-1}\right).$
2.8. The graph sum formula for descendant Gromov-Witten potential
In this subsection, we introduce the graph sum formula for the generating functions of descendant Gromov-Witten invariants. Given a connected graph $\Gamma$, we introduce the following notation.
  1. (1)
    $V(\Gamma)$ is the set of vertices in $\Gamma$.
  2. (2)
    $E(\Gamma)$ is the set of edges in $\Gamma$.
  3. (3)
    $H(\Gamma)$ is the set of half-edges in $\Gamma$.
  4. (4)
    $L^{o}(\Gamma)$ is the set of ordinary leaves in $\Gamma$. The ordinary leaves are ordered: $L^{o}(\Gamma)=\{l_{1},\dots,l_{n}\}$ where $n$ is the number of ordinary leaves.
  5. (5)
    $L^{1}(\Gamma)$ is the set of dilaton leaves in $\Gamma$. The dilaton leaves are unordered.
Withe the above notation, we introduce the following labels:
  1. (1)
    (genus) $g:V(\Gamma)\rightarrow\mathbb{Z}_{\geq 0}$.
  2. (2)
    (marking) $\beta:V(\Gamma)\rightarrow\{1,2\}$. This induces $\beta:L(\Gamma)=L^{o}(\Gamma)\cup L^{1}(\Gamma)\rightarrow\{1,2\}$, as follows: if $l\in L(\Gamma)$ is a leaf attached to a vertex $v\in V(\Gamma)$, define $\beta(l)=\beta(v)$.
  3. (3)
    (height) $k:H(\Gamma)\rightarrow\mathbb{Z}_{\geq 0}$.
Given an edge $e$, let $h_{1}(e)$ and $h_{2}(e)$ be the two half-edges associated to $e$. The order of the two half-edges does not affect the graph sum formula in this paper. Given a vertex $v\in V(\Gamma)$, let $H(v)$ denote the set of half-edges emanating from $v$. The valency of the vertex $v$ is equal to the cardinality of the set $H(v)$: ${\mathrm{val}}(v)=|H(v)|$. A labeled graph $\vec{\Gamma}=(\Gamma,g,\beta,k)$ is stable if
$02g(v)-2+{\mathrm{val}}(v)>$ 0
for all $v\in V(\Gamma)$.
Let ${\bf\Gamma}(\mathbb{P}^{1})$ denote the set of all stable labeled graphs $\vec{\Gamma}=(\Gamma,g,\beta,k)$. The genus of a stable labeled graph $\vec{\Gamma}$ is defined to be
$g(\vec{\Gamma}):=\sum_{v\in V(\Gamma)}g(v)+|E(\Gamma)|-|V(\Gamma)|+1=\sum_{v\in V(\Gamma)}(g(v)-1)+(\sum_{e\in E(\Gamma)}1)+1.$
Define
${\bf\Gamma}_{g,n}(\mathbb{P}^{1})=\{\vec{\Gamma}=(\Gamma,g,\beta,k)\in{\bf\Gamma}(\mathbb{P}^{1}):g(\vec{\Gamma})=g,|L^{o}(\Gamma)|=n\}.$
We assign weights to leaves, edges, and vertices of a labeled graph $\vec{\Gamma}\in{\bf\Gamma}(\mathbb{P}^{1})$ as follows.
  1. (1)
    Ordinary leaves. To each ordinary leaf $l_{j}\in L^{o}(\Gamma)$ with $\beta(l_{j})=\beta\in\{1,2\}$ and $k(l_{j})=k\in\mathbb{Z}_{\geq 0}$, we assign the following descendant weight:
    $(\mathcal{L}^{\bf u})^{\beta}_{k}(l_{j})=[z^{k}](\sum_{\alpha,\gamma\in\{1,2\}}\left(\frac{{\bf u}^{\alpha}_{j}(z)}{\sqrt{\Delta^{\alpha}(q)}}S^{\hat{\underline{\gamma}}}_{\ \hat{\underline{\alpha}}}(z)\right)_{+}R(-z)_{\gamma}^{\ \beta}).$
    where $(\cdot)_{+}$ means taking the nonnegative powers of $z$.
  2. (2)
    Dilaton leaves. To each dilaton leaf $l\in L^{1}(\Gamma)$ with $\beta(l)=\beta\in\{1,2\}$ and $2\leq k(l)=k\in\mathbb{Z}_{\geq 0}$, we assign
    $(\mathcal{L}^{1})^{\beta}_{k}(l)=[z^{k-1}](-\sum_{\alpha\in\{1,2\}}\frac{1}{\sqrt{\Delta^{\alpha}(q)}}R_{\alpha}^{\ \beta}(-z)).$
  3. (3)
    Edges. To an edge connecting a vertex marked by $\alpha\in\{1,2\}$ and a vertex marked by $\beta\in\{1,2\}$, and with heights $k$ and $l$ at the corresponding half-edges, we assign
    $\mathcal{E}^{\alpha,\beta}_{k,l}=[z^{k}w^{l}]\Big{(}\frac{1}{z+w}(\delta_{\alpha,\beta}-\sum_{\gamma\in\{1,2\}}R_{\gamma}^{\ \alpha}(-z)R_{\gamma}^{\ \beta}(-w))\Big{)}.$
  4. (4)
    Vertices. To a vertex $v$ with genus $g(v)=g\in\mathbb{Z}_{\geq 0}$ and with marking $\beta(v)=\beta$, with $n$ ordinary leaves and half-edges attached to it with heights $k_{1},\dots,k_{n}\in\mathbb{Z}_{\geq 0}$ and $m$ more dilaton leaves with heights $k_{n+1},\dots,k_{n+m}\in\mathbb{Z}_{\geq 0}$, we assign
    $\Big{(}\sqrt{\Delta^{\beta}(q)}\Big{)}^{2g(v)-2+{\mathrm{val}}(v)}\langle\tau_{k_{1}}\dots\tau_{k_{n+m}}\rangle_{g},$
    where $\langle\tau_{k_{1}}\dots\tau_{k_{n+m}}\rangle_{g}=\int_{\overline{\mathcal{M}}_{g,n+m}}\psi_{1}^{k_{1}}\dots\psi_{n+m}^{k_{n+m}}$.
We define the weight of a labeled graph $\vec{\Gamma}\in{\bf\Gamma}_{g,n}(\mathbb{P}^{1})$ to be
$\omega^{\bf u}_{A}(\vec{\Gamma})=$ $\prod_{v\in V(\Gamma)}\Big{(}\sqrt{\Delta^{\beta(v)}(q)}\Big{)}^{2g(v)-2+{\mathrm{val}}(v)}\langle\prod_{h\in H(v)}\tau_{k(h)}\rangle_{g(v)}\prod_{e\in E(\Gamma)}\mathcal{E}^{\beta(v_{1}(e)),\beta(v_{2}(e))}_{k(h_{1}(e)),k(h_{2}(e))}$
$\cdot\prod_{l\in L^{1}(\Gamma)}(\mathcal{L}^{1})^{\beta(l)}_{k(l)}(l)\prod_{j=1}^{n}(\mathcal{L}^{\bf u})^{\beta(l_{j})}_{k(l_{j})}(l_{j}).$
With the about definition of the weight of a labeled graph, we have the following theorem which expresses the $T$-equivariant descendant Gromov-Witten potential of $\mathbb{P}^{1}$ in terms of graph sum.
Theorem 2.3 ([24]).
Suppose that $02g-2+n>$ 0. Then
${\bf u}_{1},\dots,\mathbf{u}_{n}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}_{g,n}^{\mathbb{P}^{1},T}=\sum_{\vec{\Gamma}\in{\bf\Gamma}_{g,n}(\mathbb{P}^{1})}\frac{\omega^{\mathbf{u}}_{A}(\vec{\Gamma})}{|\mathrm{Aut}(\vec{\Gamma})|}.⟨$
3. Equivariant open Gromov-Witten theory of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$
3.1. Moduli spaces of stable maps to ($\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}$)
Let
$L:=\{[e^{{\rm i}\varphi},e^{-{\rm i}\varphi}]\in\mathbb{P}^{1}:\varphi\in\mathbb{R}\}$
be the Lagrangian submanifold of $\mathbb{P}^{1}$, which is preserved by the $T$-action. By taking a Möbius transform, we can identify the pair $(\mathbb{P}^{1},L)$ with $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$. Let $D_{1}$ and $D_{2}$ be the two hemispheres centered at $p_{1}$ and $p_{2}$ respectively. Then we have
$H_{2}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})=\mathbb{Z}[D_{1}]\oplus\mathbb{Z}[D_{2}].$
Sometimes, we identify the relative homology group $H_{2}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ to $\mathbb{Z}^{2}$ so that the generators $(1,0)$ and $(0,1)$ represent $D_{1}$ and $D_{2}$ respectively. The relative homology group $H_{2}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ could also be viewed as the extension of the homology groups $H_{2}(\mathbb{P}^{1})$ and $H_{1}(\mathbb{R}\mathbb{P}^{1})$. Consider the short exact sequence
[…]
For $d[\mathbb{P}^{1}]\in H_{2}(\mathbb{P}^{1})$, the map $\delta:H^{2}(\mathbb{P}^{1})\rightarrow H_{2}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ is defined by $\delta(d[\mathbb{P}^{1}])=(d,d)$. The connecting map $\partial:H_{2}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})\rightarrow H_{1}(\mathbb{R}\mathbb{P}^{1})$ is given by $\partial(a,b)=(b-a)[\mathbb{R}\mathbb{P}^{1}]$.
Let $(\Sigma,\partial\Sigma)$ be a prestable bordered Riemann surface, and let $\partial\Sigma=R_{1}\cup\dots\cup R_{h}$. The (small) genus $g(\Sigma)$ of $\Sigma$ is the genus of the closed Riemann surface obtained by capping each boundary component of $\Sigma$ with a disk. A genus-$g$ bordered Riemann surface with $h$ boundary components is called a Riemann surface of topological type $(g,h)$.
The double $\Sigma_{\mathbb{C}}$ of the prestable bordered Riemann surface $(\Sigma,\partial\Sigma)$ is obtained by gluing the surface $\Sigma$ and the conjugate surface $\bar{\Sigma}$ along the boundary $\partial\Sigma$ by the Schwartz reflection principle. The Riemann surface $\Sigma_{\mathbb{C}}$ is endowed with a canonical involution $b:\Sigma_{\mathbb{C}}\rightarrow\Sigma_{\mathbb{C}}$.
We concern about the maps $u:(\Sigma,\partial\Sigma)\rightarrow(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$. The degree of the map $u$ is the element
$\vec{d}=(d_{-},d_{+})=u_{*}[\Sigma]\in H_{2}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}).$
Let
$\mu_{i}[\mathbb{R}\mathbb{P}^{1}]=(u\mid_{R_{i}})_{*}[R_{i}]\in H_{1}(\mathbb{R}\mathbb{P}^{1})=\mathbb{Z}[\mathbb{R}\mathbb{P}^{1}],\quad i=1,\cdots,h.$
The number $\mu_{i}$ is called the $i$-th winding number. Let $\beta^{\prime}=u_{*}[\Sigma]\in H_{2}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ and let $\vec{\mu}=(\mu_{1},\dots,\mu_{h})\in\mathbb{Z}_{\neq 0}^{h}$. Then there exists $d\in\mathbb{Z}_{\geq 0}$ such that
$\beta^{\prime}=d[\mathbb{P}^{1}]+\sum_{i=1}^{h}\mu_{i}[\mathbb{R}\mathbb{P}^{1}].$
By the short exact sequence (2), we have
$d_{+}-d_{-}=\sum_{j=1}^{h}\mu_{j},$
$d=d_{+}-\sum_{\{j:\mu_{j}>0\}}\mu_{j}=d_{-}+\sum_{\{j:\mu_{j}<0\}}\mu_{j}.$
Let $(\Sigma,\partial\Sigma,x_{1},\dots,x_{n})$ be a prestable bordered Riemann surface with $n$ interior marked points, and consider stable maps:
$u:(\Sigma,\partial\Sigma,x_{1},\dots,x_{n})\rightarrow(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}).$
Let ¯$\overline{\mathcal{M}}_{(g,h),n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})$ be the moduli space of degree $\beta^{\prime}$ stable maps to $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ from type $(g,h)$ bordered Riemann surfaces with $n$ interior marked points, such that the winding numbers are given by $\mu_{i}\in\mathbb{Z}$.
3.2. Equivariant open Gromov-Witten invariants
Let $\gamma_{1},\dots,\gamma_{n}\in H^{*}_{T}(\mathbb{P}^{1};\mathbb{C})$, $\beta^{\prime}\in H_{2}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$. The $T$-action on $\mathbb{P}^{1}$ induces a $T$-action on the moduli space ¯$\overline{\mathcal{M}}_{(g,h),n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})$. Let $F:=\overline{\mathcal{M}}_{(g,h),n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})^{T}$ be the $T$-fixed locus. The evaluation map $\mathrm{ev}_{i}:\overline{\mathcal{M}}_{(g,h),n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})\to\mathbb{P}^{1}$ is $T$-equivariant.
For $i=1,\dots,n$, let $\mathbb{L}_{i}$ be the $i$-th tautological line bundle over ¯$\overline{\mathcal{M}}_{(g,h),n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})$ formed by the cotangent line at the $i$-th marked point. Define the $i$-th descendant class $\psi_{i}$ as
$\psi_{i}:=c_{1}(\mathbb{L}_{i})\in H^{2}(\overline{\mathcal{M}}_{(g,h),n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu});\mathbb{Q}).$
We choose a $T$-equivariant lift $\psi_{i}^{T}\in H^{2}_{T}(\overline{\mathcal{M}}_{(g,h),n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu});\mathbb{Q})$ of $\psi_{i}$. We define the $T$-equivariant open Gromov-Witten invariants of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$:
$\langle\tau_{a_{1}}(\gamma_{1})\dots\tau_{a_{n}}(\gamma_{n})\rangle^{\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1},T}_{g,\beta^{\prime},\vec{\mu}}:=\int_{[F]^{\mathrm{vir}}}\frac{\prod_{i=1}^{n}(\psi_{i}^{T})^{a_{i}}\mathrm{ev}_{i}^{*}(\gamma_{i})|_{F}}{e_{T}(N^{\mathrm{vir}})}\in\mathbb{C}[\mathsf{v},\frac{1}{\mathsf{v}}],$
where $[F]^{\mathrm{vir}}$ is the virtual fundamental class of $F$, and $N^{\mathrm{vir}}$ is the virtual normal bundle of $F$. Since $F$ is a compact orbifold without boundary, the above integral is well-defined.
3.3. Virtual localization formula
Definition 3.1 (Decorated graphs (a)).
Let $n\in\mathbb{Z}_{\geq 0}$ and $\beta^{\prime}=d[\mathbb{P}^{1}]+\sum_{i=1}^{h}\mu_{i}[\mathbb{R}\mathbb{P}^{1}]\in H_{2}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$. A genus $g$, $n$-pointed, degree $\beta^{\prime}$ decorated graph for $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ is a tuple $\vec{\Gamma}=(\Gamma,\vec{f},\vec{d},\vec{g},\vec{s})$ consisting of the following data.
  1. (1)
    $\Gamma$ is a compact, connected 1-dimensional CW complex. Let $V(\Gamma)\sqcup V_{\circ}(\Gamma)$ denote the set of vertices in $\Gamma$, denoted by $\bullet$ and $\circ$, respectively. The $\circ$ vertex is univalent. Let $E(\Gamma)$ denote the set of edges, where an edge $e$ is a line connecting two $\bullet$ vertices. Let $H(\Gamma)=\{h_{j}:j=1,\dots,h\}$ denote the set of half-edges, where a half-edge $h_{j}$ is a line connecting one $\bullet$ vertex and one $\circ$ vertex. We require the half-edges to be ordered. Let $F(\Gamma)$ be the set of flags:
    $\{(e,v)\in E(\Gamma)\times V(\Gamma):v\in e\}\cup\{(h_{j},v)\in H(\Gamma)\times V(\Gamma):v\in h_{j}\}.$
    For each $v\in V(\Gamma)$, let $F_{v}$ (resp. $E_{v}$, $H_{v}$) denote the flags (resp. edges, half-edges) attached to $v$, and let ${\mathrm{val}}(v)=|F_{v}|$ denote the number of flags incident to $v$.
  2. (2)
    The label map $\vec{f}:V(\Gamma)\rightarrow\{1,2\}$ labels each $\bullet$ with a number. If $v_{1},v_{2}\in V(\Gamma)$ are connected by an edge, we require $\vec{f}(v_{1})\neq\vec{f}(v_{2})$.
  3. (3)
    The degree map $\vec{d}:E(\Gamma)\cup H(\Gamma)\rightarrow\mathbb{Z}_{>0}$ sends an edge $e$ (resp. a half-edge $h_{j}$) to a positive integer $\vec{d}(e)=d_{e}$ (resp. $\vec{d}(h_{j})=d_{h_{j}}$).
  4. (4)
    The genus map $\vec{g}:V(\Gamma)\rightarrow\mathbb{Z}_{\geq 0}$ sends a vertex $v\in V(\Gamma)$ to a nonnegative integer $g_{v}$.
  5. (5)
    The marking map $\vec{s}:\{1,2,\dots,n\}\rightarrow V(\Gamma)$. For each $v\in V(\Gamma)$, define $S_{v}:=\vec{s}^{-1}(v)$, and $n_{v}=|S_{v}|$.
The data is required to satisfy the following conditions:
  • (i)
    (genus) $g=\sum_{v\in V(\Gamma)}g_{v}+|E(\Gamma)|-|V(\Gamma)|+1$.
  • (ii)
    (degree) $d=\sum_{e\in E(\Gamma)}d_{e}$.
  • (iii)
    (winding numbers) Let $h_{j}\in H(\Gamma)$ be the $j$-th half-edge, and let $v_{j}\in V(\Gamma)$ be its incident $\bullet$ vertex, then $\mu_{j}=(-1)^{\vec{f}(v_{j})}\vec{d}(h_{j})$.
Let $G_{g,n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})$ be the set of all decorated graphs $\vec{\Gamma}=(\Gamma,\vec{f},\vec{d},\vec{g},\vec{s})$ satisfying the above constraints.
Given a decorated graph $\vec{\Gamma}\in G_{g,n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})$. Let $V^{S}(\Gamma)$ be the set of stable vertices:
$V^{S}(\Gamma):=\{v\in V(\Gamma):2g_{v}-2+{\mathrm{val}}(v)+n_{v}>0\}.$
Let $\mathrm{Aut}(\vec{\Gamma})$ denote the group of automorphisms of $\vec{\Gamma}$, and let $A^{0}_{\vec{\Gamma}}$ be the group of covering automorphisms
$A^{0}_{\vec{\Gamma}}=\prod_{e\in E(\Gamma)}\mathbb{Z}_{\vec{d}(e)}\times\prod_{h_{j}\in H(\Gamma)}\mathbb{Z}_{\vec{d}(h_{j})}.$
The automorphism group $A_{\vec{\Gamma}}$ fits in the short exact sequence:
$1\rightarrow A^{0}_{\vec{\Gamma}}\rightarrow A_{\vec{\Gamma}}\rightarrow\mathrm{Aut}(\vec{\Gamma})\rightarrow 1$
We set
¯$\overline{\mathcal{M}}_{\vec{\Gamma}}:=\prod_{v\in V^{S}(\Gamma)}\overline{\mathcal{M}}_{g_{v},F_{v}\cup S_{v}},\quad F_{\vec{\Gamma}}:=[\overline{\mathcal{M}}_{\vec{\Gamma}}/A_{\vec{\Gamma}}].$
Every decorated graph $\vec{\Gamma}\in G_{g,n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})$ represents a topological type of the $T$-invariant stable open map $u:(\Sigma,\partial\Sigma,x_{1},\dots,x_{n})\rightarrow(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ in the following way:
  1. (1)
    Every $v\in V(\Gamma)$ represents a stable curve or a single point $\Sigma_{v}$ in the domain curve $\Sigma$. If $\Sigma_{v}$ is a stable curve, it is of genus-$g_{v}$ with marked points $S_{v}$. The map $u$ contracts $\Sigma_{v}$ to the fixed-point $p_{\vec{f}(v)}$.
  2. (2)
    Every half-edge $h_{j}\in H(\Gamma)$ represents a disk $D_{j}$ in the domain curve $\Sigma$. The map $u|_{D_{j}}$ is a standard degree $\vec{d}(h_{j})=|\mu_{j}|$ cover of disk branched at $p_{\vec{f}(v_{j})}$, where $v_{j}$ is the $\bullet$ vertex attached at $h_{j}$.
  3. (3)
    Every edge $e\in E(\Gamma)$ represents a sphere $\mathbb{P}^{1}\cong C\subset\Sigma$, so that $u|_{C}$ is a degree $\vec{d}(e)=d_{e}$ cover of $\mathbb{P}^{1}$ branched at $p_{1},p_{2}$.
Vice versa, given a $T$-invariant stable open map $u$, one can read off its decorated graph $\vec{\Gamma}$. Therefore, we can use the decorated graphs to describe the connected components of the $T$-fixed locus of the moduli space ¯$\overline{\mathcal{M}}_{(g,h),n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})$:
(3) ¯$\overline{\mathcal{M}}_{(g,h),n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})^{T}=\bigcup_{\vec{\Gamma}\in G_{g,n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})}F_{\vec{\Gamma}}.$
[…]
Figure 1. The left figure is an example of the graph $\vec{\Gamma}$in $G_{g,n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})$. The right figure is an example of stable map corresponding to $\vec{\Gamma}$.
Given $\vec{\Gamma}\in G_{g,n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})$, we introduce the following notations:
  • Let $J_{\pm}=\{j\in\{1,\dots,h\}:\pm\mu_{j}>0\}$.
  • (weight) We define
    ${\bf w}(p_{1})=-\mathsf{v},\quad{\bf w}(p_{2})=\mathsf{v},$
    For a flag $f\in F_{v}$, we define
    ${\bf w}_{f}:=\left\{\begin{aligned} &\frac{{\bf w}(p_{\vec{f}(v)})}{d_{e}},\quad f=(e,v),\\ &\frac{{\bf w}(p_{\vec{f}(v)})}{d_{h_{j}}},\quad f=(h_{j},v).\end{aligned}\right.$
  • (vertex contribution) Let $g\in\mathbb{Z}_{\geq 0}$,
    ${\bf h}(p_{i},g)=\frac{\Lambda^{\vee}_{g}({\bf w}(p_{i}))}{{\bf w}(p_{i})},$
    where $\Lambda^{\vee}_{g}(\mathsf{u}):=\sum_{i=0}^{g}(-1)^{i}\lambda_{i}\mathsf{u}^{g-i}$ and $\lambda_{i}$ is the $i$-th Hodge class on ¯$\overline{\mathcal{M}}_{g,n}$.
  • (edge contribution) For $d\in\mathbb{Z}_{>0}$, we define
    ${\bf h}(e,d)=\frac{(-1)^{d}d^{2d}}{(d!)^{2}\mathsf{v}^{2d}}.$
  • (disk factor) For $\mu\in\mathbb{Z}_{>0}$, we define
    $D^{1}(\mu)=(-1)^{\mu+1}\frac{\mu^{\mu-2}}{\mu!\mathsf{v}^{\mu-2}},\quad D^{2}(\mu)=\frac{\mu^{\mu-2}}{\mu!\mathsf{v}^{\mu-2}}.$
By the virtual localization formula in [6], we get the following proposition.
Proposition 3.2.
Let $\beta^{\prime}=d[\mathbb{P}^{1}]+\sum_{i=1}^{h}\mu_{i}[\mathbb{R}\mathbb{P}^{1}]\in H_{2}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$. Then for $\gamma_{1},\dots,\gamma_{n}\in H^{*}_{T}(\mathbb{P}^{1})$ and $g,a_{1},\dots,a_{n}\in\mathbb{Z}_{\geq 0}$, we have
$\langle\tau_{a_{1}}(\gamma_{1}),\dots,\tau_{a_{n}}(\gamma_{n})\rangle^{\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1},T}_{g,\beta^{\prime},\vec{\mu}}$
$=$ $\sum_{\vec{\Gamma}\in G_{g,n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})}\frac{1}{|\mathrm{Aut}(\vec{\Gamma})|}\prod_{e\in E(\Gamma)}\frac{{\bf h}(e,d_{e})}{d_{e}}\prod_{v\in V(\Gamma)}\Big{(}{\bf w}(p_{\vec{f}(v)})^{|E_{v}|}\prod_{i\in S_{v}}i^{*}_{p_{\vec{f}(v)}}\gamma_{i}\Big{)}$
$\cdot\prod_{j\in J_{-}}D^{1}(-\mu_{j})\prod_{j\in J_{+}}D^{2}(\mu_{j})\prod_{v\in V(\Gamma)}\int_{\overline{\mathcal{M}}_{g_{v},F_{v}\cup S_{v}}}\frac{{\bf h}(p_{\vec{f}(v)},g_{v})\prod_{i\in S_{v}}\psi_{i}^{a_{i}}}{\prod_{h_{j}\in H_{v}}{\bf w}_{(h_{j},v)}\prod_{f\in F_{v}}({\bf w}_{f}-\psi_{f})}.$
We use the following convention for the unstable integrals:
¯$\int_{\overline{\mathcal{M}}_{0,1}}\frac{1}{{\bf w}-\psi}={\bf w},\quad\int_{\overline{\mathcal{M}}_{0,2}}\frac{\psi_{2}^{a}}{{\bf w}-\psi_{1}}=(-{\bf w})^{a}$ $,\quad a\in\mathbb{Z}_{\geq 0},$
¯$\int_{\overline{\mathcal{M}}_{0,2}}\frac{1}{({\bf w}_{1}-\psi_{1})({\bf w}_{2}-\psi_{2})}=\frac{1}{{\bf w}_{1}+{\bf w}_{2}}$
3.4. Open/descendant correspondence
Given $u:(\Sigma,\partial\Sigma,{\bf x})\rightarrow(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ that represents a point $\xi\in\overline{\mathcal{M}}_{(g,h),n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})^{T}$, we have
$\Sigma=C\cup\bigcup_{j=1}^{h}D_{j},$
where $C$ is a closed nodal curve of genus $g$ with marked points $x_{1},\dots,x_{n}$, and $D_{j}$’s are holomorphic disks. $C$ and $D_{j}$ intersect at $y_{j}$. Let $\hat{u}=u|_{C}$ and $u_{j}=u|_{D_{j}}$. Then $\hat{u}:(C,{\bf x},{\bf y})\rightarrow\mathbb{P}^{1}$ represents a point $\hat{\xi}\in\overline{\mathcal{M}}_{g,n+h}(\mathbb{P}^{1},\beta)^{T}$, where the $(n+j)$-th marked point $x_{n+j}$ is $y_{j}$.
The following definition is introduced in [32, Definition 52].
Definition 3.3 (Decorated graphs (b)).
Define $G_{g,n+h}(\mathbb{P}^{1},\beta)$ to be the set of all decorated graphs $\hat{\Gamma}=(\Gamma,\vec{f},\vec{d},\vec{g},\vec{s})$ defined as follows. Let $n\in\mathbb{Z}_{\geq 0}$, $h\in\mathbb{Z}_{>0}$ and $\beta=d[\mathbb{P}^{1}]\in E(\mathbb{P}^{1})$. A genus $g$, $(n+h)$-pointed, degree $\beta$ decorated graph for $\mathbb{P}^{1}$ is a tuple $\hat{\Gamma}=(\Gamma,\vec{f},\vec{d},\vec{g},\vec{s})$ consisting of the following data.
  1. (1)
    $\Gamma$ is a compact, connected 1-dimensional CW complex. Let $V(\Gamma)$ denote the set of vertices in $\Gamma$, denoted by $\bullet$. Let $E(\Gamma)$ denote the set of edges, where an edge $e$ is a line connecting two $\bullet$ vertices. Let $F(\Gamma)$ be the set of flags:
    $\{(e,v)\in E(\Gamma)\times V(\Gamma):v\in e\}.$
    For each $v\in V(\Gamma)$, let $F_{v}$ (resp. $E_{v}$) denote the flags (resp. edges) attached to $v$, and let ${\mathrm{val}}(v)=|F_{v}|=|E_{v}|$ denote the number of flags (resp. edges) incident to $v$.
  2. (2)
    The label map $\vec{f}:V(\Gamma)\rightarrow\{1,2\}$ labels each $\bullet$ with a number. If $v_{1},v_{2}\in V(\Gamma)$ are connected by an edge, we require $\vec{f}(v_{1})\neq\vec{f}(v_{2})$.
  3. (3)
    The degree map $\vec{d}:E(\Gamma)\rightarrow\mathbb{Z}_{>0}$ sends an edge $e$ to a positive integer $\vec{d}(e)=d_{e}$.
  4. (4)
    The genus map $\vec{g}:V(\Gamma)\rightarrow\mathbb{Z}_{\geq 0}$ sends a vertex $v\in V(\Gamma)$ to a nonnegative integer $g_{v}$.
  5. (5)
    The marking map $\vec{s}:\{1,2,\dots,n+h\}\rightarrow V(\Gamma)$. For each $v\in V(\Gamma)$, define $S_{v}:=\vec{s}^{-1}(v)$, and $n_{v}=|S_{v}|$.
The data is required to satisfy the following conditions:
  • (i)
    (genus) $g=\sum_{v\in V(\Gamma)}g_{v}+|E(\Gamma)|-|V(\Gamma)|+1$.
  • (ii)
    (degree) $d=\sum_{e\in E(\Gamma)}d_{e}$.
Let $F_{\hat{\Gamma}}$ be the connected component corresponding to $\hat{\Gamma}$. We have
¯$\overline{\mathcal{M}}_{g,n+h}(\mathbb{P}^{1},\beta)^{T}=\bigcup_{\hat{\Gamma}\in G_{g,n+h}(\mathbb{P}^{1},\beta)}F_{\hat{\Gamma}}.$
By [32, Theorem 73], we get
Proposition 3.4.
Let $\beta=d[\mathbb{P}^{1}]\in E(\mathbb{P}^{1})$. Then for $\gamma_{1},\dots,\gamma_{n+h}\in H^{*}_{T}(\mathbb{P}^{1})$ and $g,a_{1},\dots,a_{n+h}\in\mathbb{Z}_{\geq 0}$, we have
$\langle\tau_{a_{1}}(\gamma_{1}),\dots,\tau_{a_{n+h}}(\gamma_{n+h})\rangle^{\mathbb{P}^{1},T}_{g,n+h,\beta}$
$=$ $\sum_{\hat{\Gamma}\in G_{g,n+h}(\mathbb{P}^{1},\beta)}\frac{1}{|\mathrm{Aut}(\hat{\Gamma})|}\prod_{e\in E(\Gamma)}\frac{{\bf h}(e,d_{e})}{d_{e}}\prod_{v\in V(\Gamma)}\Big{(}{\bf w}(p_{\vec{f}(v)})^{|E_{v}|}\prod_{i\in S_{v}}i^{*}_{p_{\vec{f}(v)}}\gamma_{i}\Big{)}$
$\cdot\prod_{v\in V(\Gamma)}\int_{\overline{\mathcal{M}}_{g_{v},E_{v}\cup S_{v}}}\frac{{\bf h}(p_{\vec{f}(v)},g_{v})\prod_{i\in S_{v}}\psi_{i}^{a_{i}}}{\prod_{e\in E_{v}}({\bf w}_{(e,v)}-\psi_{(e,v)})}.$
Let $\vec{\Gamma}$ and $\hat{\Gamma}$ be the decorated graphs corresponding to the map $u$ and $\hat{u}$ respectively. There is a map
(4) $G_{g,n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})\rightarrow G_{g,n+h}(\mathbb{P}^{1},\beta):\vec{\Gamma}\mapsto\hat{\Gamma}.$
by cutting all half-edges $h_{j}$ of $\vec{\Gamma}$, and then labeling the attached $\bullet$ vertex of $h_{j}$ with $n+j$. Because the half-edges $h_{j}$ are labeled, we have $\mathrm{Aut}(\vec{\Gamma})=\mathrm{Aut}(\hat{\Gamma})$.
By Proposition 3.2, Equation (4) and Proposition 3.4, we give a open/descendant correspondence theorem:
Theorem 3.5.
Let $\beta^{\prime}=d[\mathbb{P}^{1}]+\sum_{i=1}^{h}\mu_{i}[\mathbb{R}\mathbb{P}^{1}]\in H_{2}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$. Then for $\gamma_{1},\dots,\gamma_{n}\in H^{*}_{T}(\mathbb{P}^{1})$ and $g,a_{1},\dots,a_{n}\in\mathbb{Z}_{\geq 0}$, we have
$\langle\tau_{a_{1}}(\gamma_{1})\dots\tau_{a_{n}}(\gamma_{n})\rangle^{\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1},T}_{g,\beta^{\prime},\vec{\mu}}=\prod_{j\in J_{-}}D^{1}(-\mu_{j})\prod_{j\in J_{+}}D^{2}(\mu_{j})$
[¯$\cdot\int_{[\overline{\mathcal{M}}_{g,n+h}(\mathbb{P}^{1},\beta)]^{{\mathrm{vir}},T}}\frac{\prod_{i=1}^{n}\psi_{i}^{a_{i}}\mathrm{ev}_{i}^{*}(\gamma_{i})\prod_{j\in J_{-}}\mathrm{ev}_{n+j}^{*}\phi_{1}\prod_{j\in J_{+}}\mathrm{ev}_{n+j}^{*}\phi_{2}}{\prod_{j=1}^{h}\frac{\mathsf{v}}{\mu_{j}}(\frac{\mathsf{v}}{\mu_{j}}-\psi_{n+j})}.$
3.5. Generating function of open Gromov-Witten invariants
Consider the generating function of genus $g$, $h$ boundary circles open Gromov-Witten invariants of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$: let $\mathbf{t}=t^{0}1+t^{1}H$,
$F_{g,h}(\mathbf{t},Q;X_{1},\dots,X_{h})=\sum_{\beta\in E(\mathbb{P}^{1})\atop{l\geq 0}}\sum_{\vec{\mu}=(\mu_{1},\dots,\mu_{h})\atop{\mu_{i}\in\mathbb{Z}_{\neq 0}}}\frac{Q^{\beta}}{l!}\langle\mathbf{t}^{l}\rangle^{\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1},T}_{g,\beta^{\prime},\vec{\mu}}\prod_{i=1}^{h}X_{i}^{\mu_{i}}.$
Let
$\tilde{\xi}^{1}_{k}(X)=\sum_{d<0}(\frac{\mathsf{v}}{d})^{-(k+2)}D^{1}(-d)X^{d},$
$\tilde{\xi}^{2}_{k}(X)=\sum_{d>0}(\frac{\mathsf{v}}{d})^{-(k+2)}D^{2}(d)X^{d}.$
For $\alpha\in\{1,2\}$,
$\Big{(}\frac{1}{\mathsf{v}}X\frac{d}{dX}\Big{)}\tilde{\xi}^{\alpha}_{k}(X)=\tilde{\xi}^{\alpha}_{k+1}(X),$
$\tilde{\xi}^{\alpha}(z,X):=\sum_{k\in\mathbb{Z}_{\geq-2}}z^{k}\tilde{\xi}^{\alpha}_{k}(X).$
With the above notations and the Theorem 3.5, we have
Proposition 3.6.
  • (1)
    (disk invariants)
    $\ \ \ \ \ F_{0,1}(\mathbf{t},Q;X)$
    $=\sum_{d<0}S_{z}(1,\phi_{1})\Big{|}_{z=\frac{\mathsf{v}}{d}}D^{1}(-d)X^{d}+\sum_{d>0}S_{z}(1,\phi_{2})\Big{|}_{z=\frac{\mathsf{v}}{d}}D^{2}(d)X^{d}$
    $=\sum_{\alpha\in\{1,2\}}\Big{(}\frac{1}{\Delta^{\alpha}}\tilde{\xi}^{\alpha}_{-2}(X)+(\frac{t^{0}}{\Delta^{\alpha}}+\frac{t^{1}}{2})\tilde{\xi}^{\alpha}_{-1}(X)+\sum_{k\geq 0}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\phi_{\alpha}\psi^{k}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}_{0,1}^{\mathbb{P}^{1},T}\tilde{\xi}^{\alpha}_{k}(X)\Big{)}$
    $=[z^{-2}]\sum_{\alpha\in\{1,2\}}S_{z}(1,\phi_{\alpha})\tilde{\xi}^{\alpha}(z,X).$
  • (2)
    (annulus invariants)
    $\ \ \ \ \ F_{0,2}(\mathbf{t},Q;X_{1},X_{2})-F_{0,2}(0;X_{1},X_{2})$
    $=\sum_{k_{1},k_{2}\geq 0}\sum_{\alpha_{1},\alpha_{2}\in\{1,2\}}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\phi_{\alpha_{1}}\psi^{k_{1}},\phi_{\alpha_{2}}\psi^{k_{2}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}_{0,2}^{\mathbb{P}^{1},T}\tilde{\xi}_{k_{1}}^{\alpha_{1}}(X_{1})\tilde{\xi}_{k_{2}}^{\alpha_{2}}(X_{2})$
    $=[z_{1}^{-1}z_{2}^{-1}]\sum_{\alpha_{1},\alpha_{2}\in\{1,2\}}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\frac{\phi_{\alpha_{1}}}{z_{1}-\psi_{1}},\frac{\phi_{\alpha_{2}}}{z_{2}-\psi_{2}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}_{0,2}^{\mathbb{P}^{1},T}\tilde{\xi}^{\alpha_{1}}(z_{1},X_{1})\tilde{\xi}^{\alpha_{2}}(z_{2},X_{2}),$
    where
    $\frac{1}{\mathsf{v}}\Big{(}X_{1}\frac{\partial}{\partial X_{1}}+X_{2}\frac{\partial}{\partial X_{2}}\Big{)}F_{0,2}(0;X_{1},X_{2})=\sum_{\alpha\in\{1,2\}}\frac{1}{\Delta^{\alpha}}\tilde{\xi}^{\alpha}_{0}(X_{1})\tilde{\xi}^{\alpha}_{0}(X_{2}).$
    So we have
    $\ \ \ \ \ \frac{1}{\mathsf{v}}\Big{(}X_{1}\frac{\partial}{\partial X_{1}}+X_{2}\frac{\partial}{\partial X_{2}}\Big{)}F_{0,2}(\mathbf{t},Q;X_{1},X_{2})$
    $=[z_{1}^{0}z_{2}^{0}](z_{1}+z_{2})\sum_{\alpha_{1},\alpha_{2}\in\{1,2\}}V_{z_{1},z_{2}}(\phi_{\alpha_{1}},\phi_{\alpha_{2}})\tilde{\xi}^{\alpha_{1}}(z_{1},X_{1})\tilde{\xi}^{\alpha_{2}}(z_{2},X_{2}).$
  • (3)
    For $02g-2+n>$ 0,
    $\ \ \ \ \ F_{g,n}(\mathbf{t},Q;X_{1},\dots,X_{n})$
    $=\sum_{k_{1},\dots,k_{n}\geq 0}\sum_{\alpha_{1},\dots,\alpha_{n}\in\{1,2\}}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\phi_{\alpha_{1}}\psi^{k_{1}},\dots,\phi_{\alpha_{n}}\psi^{k_{n}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}_{g,n}^{\mathbb{P}^{1},T}\prod_{j=1}^{n}\tilde{\xi}_{k_{j}}^{\alpha_{j}}(X_{j})$
    $=[z_{1}^{-1}\dots z_{n}^{-1}]\sum_{\alpha_{1},\dots,\alpha_{n}\in\{1,2\}}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}\frac{\phi_{\alpha_{1}}}{z_{1}-\psi_{1}},\dots,\frac{\phi_{\alpha_{n}}}{z_{n}-\psi_{n}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}_{g,n}^{\mathbb{P}^{1},T}\prod_{j=1}^{n}\tilde{\xi}^{\alpha_{j}}(z_{j},X_{j}).$
Let $\vec{\Gamma}\in{\bf\Gamma}_{g,n}(\mathbb{P}^{1})$ be a labeled graph defined in Section 2.8. To each ordinary leaf $l_{j}\in L^{o}(\Gamma)$ with $\beta(l_{j})=\beta\in\{1,2\}$ and $k(l_{j})=k\in\mathbb{Z}_{\geq 0}$, we assign the following weight (open leaf)
(5) (~$(\tilde{\mathcal{L}}^{O})^{\beta}_{k}(l_{j})=[z^{k}](\sum_{\alpha,\gamma\in\{1,2\}}(\tilde{\xi}^{\alpha}(z,X_{j})S^{\hat{\underline{\gamma}}}_{\ \alpha})_{+}R(-z)_{\gamma}^{\ \beta}).$
We define a new weight $\widetilde{\omega}^{O}_{A}(\vec{\Gamma})$ of a labeled graph $\vec{\Gamma}\in{\bf\Gamma}_{g,n}(\mathbb{P}^{1})$ as
$\widetilde{\omega}^{O}_{A}(\vec{\Gamma})=$ $\prod_{v\in V(\Gamma)}\Big{(}\sqrt{\Delta^{\beta(v)}(q)}\Big{)}^{2g(v)-2+{\mathrm{val}}(v)}\langle\prod_{h\in H(v)}\tau_{k(h)}\rangle_{g(v)}\prod_{e\in E(\Gamma)}\mathcal{E}^{\beta(v_{1}(e)),\beta(v_{2}(e))}_{k(h_{1}(e)),k(h_{2}(e))}$
$\cdot\prod_{l\in L^{1}(\Gamma)}(\mathcal{L}^{1})^{\beta(l)}_{k(l)}(l)\prod_{j=1}^{n}(\tilde{\mathcal{L}}^{O})^{\beta(l_{j})}_{k(l_{j})}(l_{j}).$
Then by Theorem 2.3 and Proposition 3.6, we have the following graph sum formula for $F_{g,n}$.
Theorem 3.7.
For $02g-2+n>$ 0, we have
$F_{g,n}(\mathbf{t},Q;X_{1},\dots,X_{n})=\sum_{\vec{\Gamma}\in{\bf\Gamma}_{g,n}(\mathbb{P}^{1})}\frac{\widetilde{\omega}_{A}^{O}(\vec{\Gamma})}{|\mathrm{Aut}(\vec{\Gamma})|}.$
Definition 3.8.
We define
$F_{g,n}(\mathbf{t};X_{1},\dots,X_{n}):=F_{g,n}(\mathbf{t},1;X_{1},\dots,X_{n}).$
By Remark 2.1 and Proposition 3.6, the restriction of $F_{g,n}$ to $Q=1$ is well-defined.
Theorem 3.7 implies
Corollary 3.9.
Let $\omega_{A}^{O}(\vec{\Gamma})=\widetilde{\omega}_{A}^{O}(\vec{\Gamma})\big{|}_{Q=1}$. For $02g-2+n>$ 0, we have
$F_{g,n}(\mathbf{t};X_{1},\dots,X_{n})=\sum_{\vec{\Gamma}\in{\bf\Gamma}_{g,n}(\mathbb{P}^{1})}\frac{\omega_{A}^{O}(\vec{\Gamma})}{|\mathrm{Aut}(\vec{\Gamma})|}.$
Convention 3.10.
From this point onwards, we use the specialization $Q=1$, $q=e^{t^{1}}$ in the calculation.
3.6. Explicit formula of $F_{0,1}$
Following from Proposition 3.6, the generating function $F_{0,1}$ is given by
$F_{0,1}(\mathbf{t};X)=\sum_{d>0}\Big{(}J_{1}(-\mathsf{v}/d)D^{1}(d)X^{-d}+J_{2}(\mathsf{v}/d)D^{2}(d)X^{d}\Big{)},$
where $J_{\alpha}(z):=(J(z),\phi_{\alpha})_{\mathbb{P}^{1},T}$ are the components of the $J$-function in Section 2.6.
By Equation (1), for $d>0$
$J^{1}(-\mathsf{v}/d)$ $=-\mathsf{v}J_{1}(-\mathsf{v}/d)=e^{-dt^{0}/\mathsf{v}}(-\mathsf{v}/d)^{d}\Gamma(d+1)I_{d}(-2\sqrt{q}d/\mathsf{v}),$
$J^{2}(\mathsf{v}/d)$ $=\mathsf{v}J_{2}(\mathsf{v}/d)=e^{dt^{0}/\mathsf{v}}(\mathsf{v}/d)^{d}\Gamma(d+1)I_{d}(2\sqrt{q}d/\mathsf{v}).$
We get
$F_{0,1}(\mathbf{t};X)=\sum_{d>0}e^{-dt^{0}/\mathsf{v}}\frac{\mathsf{v}}{d^{2}}I_{d}(-2\sqrt{q}d/\mathsf{v})X^{-d}+\sum_{d>0}e^{dt^{0}/\mathsf{v}}\frac{\mathsf{v}}{d^{2}}I_{d}(2\sqrt{q}d/\mathsf{v})X^{d}.$
Let $q,\mathsf{v}$ be positive real numbers. By the symmetry of the modified Bessel function $I_{\alpha}(x)$ (see Appendix A), we have
(6) $F_{0,1}(\mathbf{t};X)=\sum_{d\in\mathbb{Z}_{\neq 0}}e^{dt^{0}/\mathsf{v}}\frac{\mathsf{v}}{d^{2}}I_{d}(2\sqrt{q}d/\mathsf{v})X^{d}.$
4. Mirror curve and topological recursion
4.1. Mirror curve
Let $Y$ be a coordinate on $\mathbb{C}^{*}$ and let $q=e^{t^{1}},\mathsf{v}$ be positive real numbers. The $T$-equivariant superpotential $W:\mathbb{C}^{*}\rightarrow\mathbb{C}$ is defined as
$W(Y)=t^{0}+\frac{\mathsf{v}}{2}\log q+Y+\frac{q}{Y}-\mathsf{v}\log Y.$
Consider the spectral curve $C_{q}$ defined as follows:
$C_{q}=\{(x,y)\in\mathbb{C}^{2}:x=W(e^{y})\}.$
Let $X=e^{-x/\mathsf{v}}$, $Y=e^{y}$. The curve $C_{q}$ has a parameterization:
[…]
Let $\mathbb{P}^{1}$ be the compactification of $\mathbb{C}^{*}$ with $Y\in\mathbb{C}^{*}\subset\mathbb{P}^{1}$ as its coordinate. Consider the differential $dx(Y)=(Y-q/Y-\mathsf{v})dy$. The branch points $P_{\alpha}$ of $dx$ are given by
$P_{1}=\frac{\mathsf{v}-\mathsf{v}\sqrt{1+4q/\mathsf{v}^{2}}}{2},\quad P_{2}=\frac{\mathsf{v}+\mathsf{v}\sqrt{1+4q/\mathsf{v}^{2}}}{2}.$
Near each branch point $Y=P_{\alpha}$, we choose a local coordinates $\zeta_{\alpha}$ such that
$x=\check{u}^{\alpha}-\zeta_{\alpha}^{2},\quad y=\check{v}^{\alpha}-\sum_{k=1}^{\infty}h^{\alpha}_{k}(q)\zeta^{k}_{\alpha}.$
Let $\lambda=xdy$ be the Liouville form on $\mathbb{C}^{2}$, $\Phi:=\lambda|_{C_{q}}$. Let $B(Y_{1},Y_{2})$ be a holomorphic bidifferential on $\mathbb{P}^{1}$ defined by
$B(Y_{1},Y_{2})=\frac{dY_{1}dY_{2}}{(Y_{1}-Y_{2})^{2}}.$
Around $(P_{\alpha},P_{\beta})$, we expand $B(Y_{1},Y_{2})$ as
$B(\zeta_{\alpha},\zeta_{\beta})=\Big{(}\frac{\delta_{\alpha\beta}}{(\zeta_{\alpha}-\zeta_{\beta})^{2}}+\sum_{k,l\geq 0}B^{\alpha,\beta}_{k,l}\zeta^{k}_{\alpha}\zeta^{l}_{\beta}\Big{)}d\zeta_{\alpha}d\zeta_{\beta}.$
Given $\alpha=1,2$ and $d\in\mathbb{Z}_{\geq 0}$, define
$d\xi_{\alpha,d}(p):=(2d-1)!!2^{-d}\mathrm{Res}_{p^{\prime}\rightarrow P_{\alpha}}B(p,p^{\prime})(\sqrt{-1}\zeta_{\alpha})^{-2d-1}.$
Then $d\xi_{\alpha,d}$ is a meromorphic 1-form on $\mathbb{P}^{1}$ with a single pole of order $22d+$ 2 at $P_{\alpha}$. In the local coordinate $\zeta_{\alpha}$ near $P_{\alpha}$, $d\xi_{\alpha,d}$ has the following expansion:
$d\xi_{\alpha,d}=\left(\frac{-(2d+1)!!}{2^{d}\sqrt{-1}^{2d+1}\zeta_{\alpha}^{2d+2}}+\text{analytic part in $\zeta_{\alpha}$}\right)d\zeta_{\alpha}.$
In particular, we have
$d\xi_{\alpha,0}=\sqrt{\frac{-2}{\Delta^{\alpha}(q)}}d\left(\frac{P_{\alpha}}{Y-P_{\alpha}}\right).$
4.2. The $\mathfrak{h}_{X}$-operator
The function $X=e^{-x/\mathsf{v}}$ has an essential singularity at $Y=0$:
$X=e^{-x/\mathsf{v}}=\frac{e^{-t^{0}/\mathsf{v}}}{\sqrt{q}}Ye^{-(Y+q/Y)/\mathsf{v}}.$
We also notice that $dX/X$ is a meromorphic differential on $\mathbb{P}^{1}$:
$\frac{dX}{X}=-\frac{dx}{\mathsf{v}}=\frac{Y^{2}-\mathsf{v}Y-q}{-\mathsf{v}Y^{2}}dY=\frac{(Y-P_{1})(Y-P_{2})}{-\mathsf{v}Y^{2}}dY.$
Let $D_{\epsilon}$ be a punctured disk around $Y=0$:
$D_{\epsilon}:=\{Y\in\mathbb{C}^{*}:0<|Y|<\epsilon\},$
where $\epsilon$ is a small positive real number such that $D_{\epsilon}$ does not contain any ramification points $\{P_{1},P_{2}\}$. In the punctured disk $D_{\epsilon}$, $Y$ can be expanded as a power series of $X$ by the Lagrange inversion theorem. At $Y=0$, we define the $\mathfrak{h}_{X}$-operator, which expands holomorphic functions $f(Y)$ and holomorphic differential forms $\theta(Y)$ on $D_{\epsilon}$ to Laurent series in $X$, by their local behaviors at $Y=0$:
  • For a holomorphic function $f(Y)$ on $D_{\epsilon}$, let
    $\mathfrak{h}_{X}(f):=\sum_{\mu\in\mathbb{Z}_{\neq 0}}\Big{(}\mathop{\mathrm{Res}}\limits_{Y\rightarrow 0}f(Y)X^{-\mu}\frac{dX}{X}\Big{)}X^{\mu}.$
  • For a holomorphic differential form $\theta(Y)$ on $D_{\epsilon}$, let
    $\mathfrak{h}_{X}(\theta):=\sum_{\mu\in\mathbb{Z}_{\neq 0}}\Big{(}\mathop{\mathrm{Res}}\limits_{Y\rightarrow 0}\theta(Y)X^{-\mu}\Big{)}\frac{X^{\mu}}{\mu}.$
    If $f$ is a holomorphic function on $D_{\epsilon}$ such that $\theta(Y)=df$, then $\mathfrak{h}_{X}(\theta)=\mathfrak{h}_{X}(f)$ by the integrations by parts.
For multi-holomorphic functions (resp. forms) on $(D_{\epsilon})^{\times n}$, we define $\mathfrak{h}_{X_{1},\dots,X_{n}}$ as follows:
  • Let $f(Y_{1},\dots,Y_{n})$ be a holomorphic function on $(D_{\epsilon})^{\times n}$, we define
    $\mathfrak{h}_{X_{1},\dots,X_{n}}(f):=\sum_{\mu_{1},\dots,\mu_{n}\in\mathbb{Z}_{\neq 0}}\Big{(}\mathop{\mathrm{Res}}\limits_{Y_{1}\rightarrow 0}\cdots\mathop{\mathrm{Res}}\limits_{Y_{n}\rightarrow 0}f(Y_{1},\dots,Y_{n})\prod_{i=1}^{n}X^{-\mu_{i}}_{i}\frac{dX_{i}}{X_{i}}\Big{)}\prod_{i=1}^{n}X^{\mu_{i}}_{i}.$
  • Let $\theta(Y_{1},\dots,Y_{n})$ be a holomorphic differential form from the set of sections $\Gamma((D_{\epsilon})^{\times n},\omega_{C_{q}}^{\boxtimes n})$, we define
    $\mathfrak{h}_{X_{1},\dots,X_{n}}(\theta):=\sum_{\mu_{1},\dots,\mu_{n}\in\mathbb{Z}_{\neq 0}}\Big{(}\mathop{\mathrm{Res}}\limits_{Y_{1}\rightarrow 0}\cdots\mathop{\mathrm{Res}}\limits_{Y_{n}\rightarrow 0}\theta(Y_{1},\dots,Y_{n})\prod_{i=1}^{n}X^{-\mu_{i}}_{i}\Big{)}\prod_{i=1}^{n}\frac{X^{\mu_{i}}_{i}}{\mu_{i}}.$
The $\mathfrak{h}_{X}$-operator plays an essential role in the defintion of open leaves in Section 4.3. Let
$\xi_{\alpha,0}(Y)=\sqrt{\frac{-2}{\Delta^{\alpha}(q)}}\frac{P_{\alpha}}{Y-P_{\alpha}},$
which is a meromorphic function on $\mathbb{P}^{1}$. It has a simple pole at $P_{\alpha}$ and is holomorphic elsewhere. Let
$\eta:=\xi_{2,0}-\sqrt{-1}\xi_{1,0},\quad\chi:=\xi_{2,0}+\sqrt{-1}\xi_{1,0}.$
Let $q,\mathsf{v}>0$, $\Delta^{2}(q)=\mathsf{v}\sqrt{1+4q/\mathsf{v}^{2}}>0$. Then we have
$\eta$ $=\sqrt{\frac{-2}{\Delta^{2}(q)}}\left(\frac{P_{2}}{Y-P_{2}}-\frac{P_{1}}{Y-P_{1}}\right)$
$=\sqrt{\frac{-2}{\Delta^{2}(q)}}\frac{Y\Delta^{2}(q)}{(Y-P_{1})(Y-P_{2})}$
$=\frac{1}{-\mathsf{v}}\sqrt{-2\Delta^{2}(q)}\frac{dY}{dX/X}\frac{1}{Y},$
and
$\chi$ $=\sqrt{\frac{-2}{\Delta^{2}(q)}}\left(\frac{P_{2}}{Y-P_{2}}+\frac{P_{1}}{Y-P_{1}}\right)$
$=\sqrt{\frac{-2}{\Delta^{2}(q)}}\frac{\mathsf{v}Y+2q}{(Y-P_{1})(Y-P_{2})}$
$=-\sqrt{\frac{-2}{\Delta^{2}(q)}}\left(\frac{dY}{dX/X}\frac{1}{Y}+\frac{2qdY}{\mathsf{v}dX/X}\frac{1}{Y^{2}}\right).$
Apply the operator $\mathfrak{h}_{X}$ to the meromorphic functions $\eta$ and $\chi$, we obtain two Laurent series:
$\mathfrak{h}_{X}(\eta):=\sum_{\mu\in\mathbb{Z}_{\neq 0}}A_{\mu}X^{\mu},\quad\mathfrak{h}_{X}(\chi):=\sum_{\mu\in\mathbb{Z}_{\neq 0}}B_{\mu}X^{\mu}.$
By calculation,
(7) $A_{\mu}$ $=\mathrm{Res}_{Y=0}\eta X^{-\mu}\frac{dX}{X}$
$=\frac{1}{-\mathsf{v}}\sqrt{-2\Delta^{2}(q)}e^{\mu t^{0}/\mathsf{v}}\sqrt{q}^{\mu}\mathrm{Res}_{Y=0}e^{\mu(Y+q/Y)/\mathsf{v}}\frac{dY}{Y^{\mu+1}}$
$=\frac{1}{-\mathsf{v}}\sqrt{-2\Delta^{2}(q)}e^{\mu t^{0}/\mathsf{v}}I_{\mu}\Big{(}\frac{2\sqrt{q}\mu}{\mathsf{v}}\Big{)},$
and
(8) $B_{\mu}$ $=\mathrm{Res}_{Y=0}\chi X^{-\mu}\frac{dX}{X}$
$=-\sqrt{\frac{-2}{\Delta^{2}(q)}}e^{\mu t^{0}/\mathsf{v}}\sqrt{q}^{\mu}\mathrm{Res}_{Y=0}e^{\mu(Y+q/Y)/\mathsf{v}}\left(\frac{dY}{Y^{\mu+1}}+\frac{2q}{\mathsf{v}}\frac{dY}{Y^{\mu+2}}\right)$
$=-\sqrt{\frac{-2}{\Delta^{2}(q)}}e^{\mu t^{0}/\mathsf{v}}\left(I_{\mu}\Big{(}\frac{2\sqrt{q}\mu}{\mathsf{v}}\Big{)}+\frac{2\sqrt{q}}{\mathsf{v}}I_{\mu+1}\Big{(}\frac{2\sqrt{q}\mu}{\mathsf{v}}\Big{)}\right)$
$=-\sqrt{\frac{-2}{\Delta^{2}(q)}}e^{\mu t^{0}/\mathsf{v}}\frac{2\sqrt{q}}{\mathsf{v}}I^{\prime}_{\mu}\Big{(}\frac{2\sqrt{q}\mu}{\mathsf{v}}\Big{)}.$
Remark 4.1.
More details about the residue calculation can be found in Appendix B.
4.3. Chekhov-Eynard-Orantin topological recursion
Let $\omega_{g,n}$ be the differential forms defined recursively by the Chekhov-Eynard-Orantin topological recursion [15]:
$\omega_{0,1}=0,\quad\omega_{0,2}=B(Y_{1},Y_{2}).$
For $02g-2+n>$ 0,
$\omega_{g,n}(Y_{1},\dots,Y_{n})=\sum_{\alpha\in\{1,2\}}$ $\mathrm{Res}_{Y\rightarrow P_{\alpha}}\frac{-\int_{\xi=Y}^{\hat{Y}}B(Y_{n},\xi)}{2(\log(Y)-\log(\hat{Y}))dx}\Big{(}\omega_{g-1,n+1}(Y,\hat{Y},Y_{1},\dots,Y_{n-1})$
$+\sum_{g_{1}+g_{2}=g}\sum_{I\sqcup J=\{1,\dots,n-1\}}\omega_{g_{1},|I|+1}(Y,Y_{I})\omega_{g_{2},|J|+1}(\hat{Y},Y_{J})\Big{)},$
where $Y\neq P_{\alpha}$ is in a neighbourhood of $P_{\alpha}$, and $\hat{Y}\neq Y$ is the conjugate point of $Y$ such that $x(\hat{Y})=x(Y)$.
We introduce the following notations:
  1. (1)
    Given $\alpha=1,2$ and $k\in\mathbb{Z}_{\geq 0}$, define
    $W^{\alpha}_{k}:=d((-\frac{d}{dx})^{k}(\xi_{\alpha,0})),$
    $\theta_{\alpha}(z):=\sum_{k=0}^{\infty}d\xi_{\alpha,k}z^{k},\quad\hat{\theta}_{\alpha}(z):=\sum_{k=0}^{\infty}W^{\alpha}_{k}z^{k}.$
  2. (2)
    The B-model $R$-matrix $\check{R}_{\beta}^{\ \alpha}(z)$ (which is a power series of $z$) is defined by asymptotic expansion
[…]
    Here $\gamma_{\alpha}$ is the Lefschetz thimble of the map $x$, i.e. $x(\gamma_{\alpha})-\check{u}^{\alpha}\in\mathbb{R}_{\geq 0}$.
  3. (3)
    Given $\alpha,\beta\in\{1,2\}$, $k,l\geq 0$, define
    $\check{B}^{\alpha,\beta}_{k,l}$ $:=\frac{(2k-1)!!(2l-1)!!}{2^{k+l+1}}B^{\alpha,\beta}_{2k,2l}$
    $\ =[z^{k}w^{l}]\Big{(}\frac{1}{z+w}\Big{(}\delta_{\alpha\beta}-\sum_{\gamma\in\{1,2\}}\check{R}_{\gamma}^{\ \alpha}(-z)\check{R}_{\gamma}^{\ \beta}(-w)\Big{)}\Big{)}.$
  4. (4)
    Given $\alpha=1,2$ and $k\in\mathbb{Z}_{\geq 1}$, define
    $\check{h}^{\alpha}_{k}$ $:=-\frac{2(2k-1)!!h^{\alpha}_{2k-1}}{\sqrt{-1}^{2k-1}}$
    $\ =[z^{k-1}]\Big{(}\sum_{\beta}\sqrt{-1}h^{\beta}_{1}\check{R}_{\beta}^{\ \alpha}(-z)\Big{)}.$
In similar to Section 2.8, we introduce the B-model graph sum formula for $\omega_{g,n}$. Let $\mathbf{p}=(Y_{1},\dots,Y_{n})\in(\mathbb{P}^{1})^{n}$. For a labeled graph $\vec{\Gamma}\in{\bf\Gamma}_{g,n}(\mathbb{P}^{1})$ with $L^{o}(\Gamma)=\{l_{1},\dots,l_{n}\}$, and $\bullet=\mathbf{p}$ or $O$, we assign the weight of $\vec{\Gamma}$ as
$\omega_{B}^{\bullet}(\vec{\Gamma})$ $=(-1)^{g(\vec{\Gamma})-1+n}\prod_{v\in V(\Gamma)}\Big{(}\frac{h^{\beta(v)}_{1}}{\sqrt{2}}\Big{)}^{2-2g(v)-{\mathrm{val}}(v)}\langle\prod_{h\in H(v)}\tau_{k(h)}\rangle_{g(v)}$
×$\ \ \ \ \times\prod_{e\in E(\Gamma)}\check{B}^{\beta(v_{1}(e)),\beta(v_{2}(e))}_{k(h_{1}(e)),k(h_{2}(e))}\prod_{l\in L^{1}(\Gamma)}\frac{-1}{\sqrt{-2}}\check{h}^{\beta(l)}_{k(l)}\prod_{j=1}^{n}(\check{\mathcal{L}}^{\bullet})^{\beta(l_{j})}_{k(l_{j})}(l_{j}),$
where
  • (descendant leaf)
    (ˇ$(\check{\mathcal{L}}^{\mathbf{p}})^{\beta(l_{j})}_{k(l_{j})}(l_{j})=\frac{1}{\sqrt{-2}}d\xi_{\beta(l_{j}),k(l_{j})}(Y_{j}),$
  • (open leaf)
    (ˇ$(\check{\mathcal{L}}^{O})^{\beta(l_{j})}_{k(l_{j})}(l_{j})=\frac{1}{\sqrt{-2}}\mathfrak{h}_{X_{j}}(d\xi_{\beta(l_{j}),k(l_{j})}(Y_{j})).$
We have the B-model graph sum formula from [14, Theorem 3.7]:
Theorem 4.2 (Dunin-Barkowski-Orantin-Shadrin-Spitz [14]).
For $02g-2+n>$ 0,
$\omega_{g,n}(\mathbf{p})$ $=\sum_{\vec{\Gamma}\in{\bf\Gamma}_{g,n}(\mathbb{P}^{1})}\frac{\omega_{B}^{\mathbf{p}}(\vec{\Gamma})}{|\mathrm{Aut}(\vec{\Gamma})|},$
$\mathfrak{h}_{X_{1},\dots,X_{n}}(\omega_{g,n})$ $=\sum_{\vec{\Gamma}\in{\bf\Gamma}_{g,n}(\mathbb{P}^{1})}\frac{\omega_{B}^{O}(\vec{\Gamma})}{|\mathrm{Aut}(\vec{\Gamma})|}.$
We define the B-model open potentials as follows,
  • Define the B-model disk potential $W_{0,1}(t^{0},q;X)$ as a Laurent series in $X$ with constant term zero such that
    $\Big{(}-\frac{1}{\mathsf{v}}X\frac{d}{dX}\Big{)}W_{0,1}(t^{0},q;X)=\mathfrak{h}_{X}\Big{(}\frac{\partial\Phi}{\partial t^{0}}\Big{)}.$
  • Let $\tilde{\omega}_{0,2}(Y_{1},Y_{2})$ be the meromorphic 2-form on $\mathbb{C}^{*}\times\mathbb{C}^{*}$,
    $\tilde{\omega}_{0,2}(Y_{1},Y_{2}):=\omega_{0,2}(Y_{1},Y_{2})-\frac{dX_{1}dX_{2}}{(X_{1}-X_{2})^{2}}.$
    $\tilde{\omega}_{0,2}$ is holomorphic on $(D_{\epsilon})^{\times 2}$. Define the B-model annulus invariants by
    $W_{0,2}(t^{0},q;X_{1},X_{2})=\mathfrak{h}_{X_{1},X_{2}}(\tilde{\omega}_{0,2}(Y_{1},Y_{2})).$
  • For $02g-2+n>$ 0, $\omega_{g,n}$ is holomorphic at $(D_{\epsilon})^{\times n}$, we define
    $W_{g,n}(t^{0},q;X_{1},\dots,X_{n})=\mathfrak{h}_{X_{1},\dots,X_{n}}(\omega_{g,n}(\mathbf{p})).$
5. Mirror Symmetry
5.1. Mirror symmetry of disk invariants
Let $\Phi=xdy|_{C_{q}}$ which is a 1-form on $C_{q}$. Let
$\Phi_{0}:=\frac{\partial\Phi}{\partial t^{0}}=\frac{dY}{Y}.$
Consider
$\mathfrak{h}_{X}(\Phi_{0})=\sum_{\mu\in\mathbb{Z}_{\neq 0}}R_{\mu}X^{\mu},$
where
$R_{\mu}$ $=\frac{1}{\mu}\mathrm{Res}_{Y=0}X^{-\mu}\frac{dY}{Y}$
$=\frac{1}{\mu}e^{\mu t^{0}/\mathsf{v}}\sqrt{q}^{\mu}\mathrm{Res}_{Y=0}e^{\mu(Y+q/Y)/\mathsf{v}}\frac{dY}{Y^{\mu+1}}$
$=\frac{1}{\mu}e^{\mu t^{0}/\mathsf{v}}I_{\mu}\Big{(}\frac{2\sqrt{q}\mu}{\mathsf{v}}\Big{)}.$
Compare with Equation (6), we have
$\frac{\partial}{\partial t^{0}}F_{0,1}(\mathbf{t};X)=\Big{(}\frac{1}{\mathsf{v}}X\frac{d}{dX}\Big{)}F_{0,1}=\mathfrak{h}_{X}(\Phi_{0}).$
Therefore, we obtain the following theorem:
Theorem 5.1 (Mirror symmetry of disk invariants).
$F_{0,1}(\mathbf{t};X)=-W_{0,1}(t^{0},q;X).$
5.2. Identification of open leaves
We define
$U(z)(\mathbf{t},X):=\sum_{\alpha\in\{1,2\}}\tilde{\xi}^{\alpha}(z,X)S_{z}(1,\phi_{\alpha})\Big{|}_{Q=1}.$
For $m\in\mathbb{Z}_{\geq-2}$,
$\left(\frac{1}{\mathsf{v}}X\frac{d}{dX}\right)[z^{m}](U(z)(\mathbf{t},X))=[z^{m+1}](U(z)(\mathbf{t},X)).$
In Section 3.6, we have shown that
$F_{0,1}=[z^{-2}]U(z)(\mathbf{t},X)=\sum_{d\in\mathbb{Z}_{\neq 0}}e^{dt^{0}/\mathsf{v}}\frac{\mathsf{v}}{d^{2}}I_{d}(2\sqrt{q}d/\mathsf{v})X^{d},$
$[z^{0}]U(z)(\mathbf{t},X)=\left(\frac{1}{\mathsf{v}}X\frac{d}{dX}\right)^{2}F_{0,1}=\frac{1}{\mathsf{v}}\sum_{d\in\mathbb{Z}_{\neq 0}}e^{dt^{0}/\mathsf{v}}I_{d}(2\sqrt{q}d/\mathsf{v})X^{d}.$
By quantum differential equation,
$z\frac{\partial U}{\partial t^{1}}=\sum_{\alpha\in\{1,2\}}\tilde{\xi}^{\alpha}(z,X)S_{z}(H,\phi_{\alpha})\Big{|}_{Q=1},$
and
$\frac{\partial U}{\partial t^{1}}$ $=\frac{\partial}{\partial t^{1}}\left(\frac{1}{\mathsf{v}}X\frac{d}{dX}\right)[z^{-2}](U(z)(\mathbf{t},X))$
$=\frac{\partial}{\partial t^{1}}\sum_{d\in\mathbb{Z}_{\neq 0}}e^{dt^{0}/\mathsf{v}}\frac{1}{d}I_{d}(2\sqrt{q}d/\mathsf{v})X^{d}$
$=\sum_{d\in\mathbb{Z}_{\neq 0}}e^{dt^{0}/\mathsf{v}}\frac{\sqrt{q}}{\mathsf{v}}I^{\prime}_{d}(2\sqrt{q}d/\mathsf{v})X^{d}.$
Comparing with Equation (7) and Equation (8), we have
$\mathfrak{h}_{X}(\eta)$ $=-\sqrt{-2\Delta^{2}(q)}[z^{0}]U(z)(\mathbf{t},X),$
$\mathfrak{h}_{X}(\chi)$ $=-2\sqrt{\frac{-2}{\Delta^{2}(q)}}[z^{0}]\tilde{\xi}^{\alpha}(z,X)S_{z}(H,\phi_{\alpha})\Big{|}_{Q=1}.$
Recall that
[…]
We obtain
$\sum_{\beta\in\{1,2\}}\tilde{\xi}^{\beta}(z,X)S(\hat{\phi}_{1}(q),\phi_{\beta})\Big{|}_{Q=1}$ $=[z^{0}]\frac{\sqrt{-\Delta^{2}(q)}}{2}U+[z^{-1}]\frac{1}{\sqrt{-\Delta^{2}(q)}}\frac{\partial U}{\partial t^{1}}=-\frac{\mathfrak{h}_{X}(\xi_{1,0})}{\sqrt{-2}},$
$[z^{0}]\sum_{\beta\in\{1,2\}}\tilde{\xi}^{\beta}(z,X)S(\hat{\phi}_{2}(q),\phi_{\beta})\Big{|}_{Q=1}$ $=[z^{0}]\frac{\sqrt{\Delta^{2}(q)}}{2}U+[z^{-1}]\frac{1}{\sqrt{\Delta^{2}(q)}}\frac{\partial U}{\partial t^{1}}=-\frac{\mathfrak{h}_{X}(\xi_{2,0})}{\sqrt{-2}}.$
Therefore, for $\alpha\in\{1,2\}$, $k>0$,
(9) $\sum_{\beta\in\{1,2\}}\tilde{\xi}^{\beta}(z,X)S(\hat{\phi}_{\alpha}(q),\phi_{\beta})\Big{|}_{Q=1}$ $=\mathfrak{h}_{X}\Big{(}(-\frac{d}{dx})^{k}\Big{(}\frac{-1}{\sqrt{-2}}\xi_{\alpha,0}\Big{)}\Big{)},$
$\Big{(}\sum_{\beta\in\{1,2\}}\tilde{\xi}^{\beta}(z,X)S(\hat{\phi}_{\alpha}(q),\phi_{\beta})\Big{|}_{Q=1}\Big{)}_{+}$ $=\mathfrak{h}_{X}\Big{(}\sum_{k\geq 0}\frac{-1}{\sqrt{-2}}W^{\alpha}_{k}z^{k}\Big{)}=-\mathfrak{h}_{X}\Big{(}\frac{\hat{\theta}_{\alpha}(z)}{\sqrt{-2}}\Big{)}.$
Lemma 5.2.
We have
$\theta_{\alpha}(z)=\sum_{\beta\in\{1,2\}}\check{R}_{\beta}^{\ \alpha}(-z)\hat{\theta}_{\beta}(z).$
Proof.
The lemma follows from
$d\xi_{\alpha,k}=\sum_{i=0}^{k}\sum_{\beta\in\{1,2\}}([z^{k-i}]\check{R}_{\beta}^{\ \alpha}(-z))W_{i}^{\beta},$
which is shown in the proof of [19, Theorem A]. ∎
Then we have the identification of open leaves:
Theorem 5.3.
For each ordinary leaf $l_{j}$ with $\beta(l_{j})=\beta\in\{1,2\}$ and $k(l_{j})=k\in\mathbb{Z}_{\geq 0}$, we have
(~$(\tilde{\mathcal{L}}^{O})^{\beta}_{k}(l_{j})\big{|}_{Q=1}=-(\check{\mathcal{L}}^{O})^{\beta}_{k}(l_{j}).$
Proof.
By [19, Proposition 3.10], the A- and B-model $R$-matrices are equal:
$R_{\beta}^{\ \alpha}(z)\big{|}_{Q=1}=\check{R}_{\beta}^{\ \alpha}(z).$
Then Theorem 5.3 follows from Equation (5), Equation (9), and Lemma 5.2. ∎
5.3. Mirror symmetry of annulus invariants
Define
$C(Y_{1},Y_{2})$ $:=(-\frac{\partial}{\partial x(Y_{1})}-\frac{\partial}{\partial x(Y_{2})})\Big{(}\frac{\omega_{0,2}}{dx(Y_{1})dx(Y_{2})}\Big{)}(Y_{1},Y_{2})dx(Y_{1})dx(Y_{2})$
$\ =\Big{(}-d_{1}\circ\frac{1}{dx(Y_{1})}-d_{2}\circ\frac{1}{dx(Y_{2})}\Big{)}(\tilde{\omega}_{0,2}(Y_{1},Y_{2})).$
The following proposition is proved in [21, Lemma 6.9].
Proposition 5.4.
We have
$C(Y_{1},Y_{2})=\frac{1}{2}\sum_{\alpha\in\{1,2\}}d\xi_{\alpha,0}(Y_{1})d\xi_{\alpha,0}(Y_{2}).$
Theorem 5.5 (Mirror symmetry of annulus invariants).
$F_{0,2}(\mathbf{t};X_{1},X_{2})=-W_{0,2}(t^{0},q;X_{1},X_{2}).$
Proof.
$\ \ \ \ \ \mathfrak{h}_{X_{1},X_{2}}(C)=\mathfrak{h}_{X_{1},X_{2}}(\frac{1}{2}\sum_{\alpha\in\{1,2\}}d\xi_{\alpha,0}(Y_{1})d\xi_{\alpha,0}(Y_{2}))$
$=-[z_{1}^{0}z_{2}^{0}]\sum_{\alpha,\beta,\gamma\in\{1,2\}}\tilde{\xi}^{\beta}(z_{1},X_{1})\tilde{\xi}^{\gamma}(z_{2},X_{2})S_{z_{1}}(\hat{\phi}_{\alpha}(q),\phi_{\beta})S_{z_{2}}(\hat{\phi}_{\alpha}(q),\phi_{\gamma})\big{|}_{Q=1}$
$=-[z_{1}^{0}z_{2}^{0}](z_{1}+z_{2})\sum_{\beta,\gamma\in\{1,2\}}V_{z_{1},z_{2}}(\phi_{\beta},\phi_{\gamma})\tilde{\xi}^{\beta}(z_{1},X_{1})\tilde{\xi}^{\gamma}(z_{2},X_{2})\big{|}_{Q=1}$
$=-\frac{1}{\mathsf{v}}(X_{1}\frac{\partial}{\partial X_{1}}+X_{2}\frac{\partial}{\partial X_{2}})F_{0,2}(\mathbf{t};X_{1},X_{2}).$
By the integrations by parts,
$\mathfrak{h}_{X_{1},X_{2}}(C)$ $=\mathfrak{h}_{X_{1},X_{2}}\Big{(}\Big{(}-d_{1}\circ\frac{1}{dx(Y_{1})}-d_{2}\circ\frac{1}{dx(Y_{2})}\Big{)}\tilde{\omega}_{0,2}\Big{)}$
$=\Big{(}\frac{1}{\mathsf{v}}X_{1}\frac{\partial}{\partial X_{1}}+\frac{1}{\mathsf{v}}X_{2}\frac{\partial}{\partial X_{2}}\Big{)}W_{0,2}(t^{0},q;X_{1},X_{2}).$
Since the constant terms of $F_{0,2}$ and $W_{0,2}$ are both zeros, the theorem follows immediately. ∎
5.4. All genus mirror symmetry
Theorem 5.6 (All genus mirror theorem).
For any $n>0$ and $g\geq 0$, we have
$F_{g,n}(\mathbf{t};X_{1},\dots,X_{n})=(-1)^{g-1}W_{g,n}(t^{0},q;X_{1},\dots,X_{n}).$
Proof.
For the unstable cases $(g,n)=(0,1)$ and $(0,2)$, this theorem is Theorem 5.1 and Theorem 5.5 respectively.
For stable cases $02g-2+n>$ 0, it suffices to show the A- and B-model graph sum formula in Corollary 3.9 and Theorem 4.2 are equal. By [19, Theorem A], we have
$R_{\beta}^{\ \alpha}(z)\big{|}_{Q=1}=\check{R}_{\beta}^{\ \alpha}(z),$
and
$\frac{h^{\alpha}_{1}}{\sqrt{2}}=\frac{1}{\sqrt{\Delta^{\alpha}(q)}}.$
Hence, the weights in the graph sum match except for the open leaves. Since Theorem 5.3 identifies the A- and B- open leaves, the theorem follows immediately. ∎
Appendix A Bessel functions
The special function $I_{\alpha}(x)$ in $J$-function is the modified Bessel function of the first kind. It is defined as
$I_{\alpha}(x)=\sum_{m=0}^{\infty}\frac{1}{m!\Gamma(m+\alpha+1)}(\frac{x}{2})^{2m+\alpha}.$
The derivative of $I_{\alpha}(x)$ is given by
$I^{\prime}_{\alpha}(x)=I_{\alpha+1}(x)+\frac{\alpha}{x}I_{\alpha}(x).$
Let $d\in\mathbb{Z}_{\neq 0}$ and $q,\mathsf{v}>0$, we get
$I^{\prime}_{d}\Big{(}\frac{2\sqrt{q}d}{\mathsf{v}}\Big{)}=I_{d+1}\Big{(}\frac{2\sqrt{q}d}{\mathsf{v}}\Big{)}+\frac{\mathsf{v}}{2\sqrt{q}}I_{d}\Big{(}\frac{2\sqrt{q}d}{\mathsf{v}}\Big{)}.$
For $n\in\mathbb{N}$,
$I_{n}(x)=I_{-n}(x).$
We have
$I_{0}(0)=1,\quad I_{\alpha}(0)=0\quad(\alpha>0).$
Appendix B Residue Calculation
Let $\mu\in\mathbb{Z}$, $q,\mathsf{v}>0$, consider
$C_{\mu}:=\mathrm{Res}_{Y=0}e^{\mu(Y+\frac{q}{Y})/\mathsf{v}}\frac{dY}{Y^{\mu+1}},$
$D_{\mu}:=\mathrm{Res}_{Y=0}e^{\mu(Y+\frac{q}{Y})/\mathsf{v}}\frac{dY}{Y^{\mu+2}}.$
To calculate $C_{\mu}$ and $D_{\mu}$, we expand
$e^{\mu(Y+q/Y)/\mathsf{v}}=\sum_{m,n=0}^{\infty}\frac{(\frac{\mu}{\mathsf{v}})^{n}(\frac{\mu q}{\mathsf{v}})^{m}Y^{n-m}}{n!m!}.$
For $C_{\mu},\ \mu\geq 0$, let $n=m+\mu$,
$C_{\mu}$ $=\sum_{m=0}^{\infty}\frac{(\frac{\mu}{\mathsf{v}})^{m+\mu}(\frac{\mu q}{\mathsf{v}})^{m}}{(m+\mu)!m!}$
$=\frac{1}{\sqrt{q}^{\mu}}\sum_{m=0}^{\infty}\frac{1}{m!(m+\mu)!}\Big{(}\frac{\mu\sqrt{q}}{\mathsf{v}}\Big{)}^{2m+\mu}$
$=\frac{1}{\sqrt{q}^{\mu}}I_{\mu}\Big{(}\frac{2\sqrt{q}\mu}{\mathsf{v}}\Big{)}.$
For $C_{\mu},\ \mu<0$, let $m=n-\mu$,
$C_{\mu}$ $=\sum_{n=0}^{\infty}\frac{(\frac{\mu}{\mathsf{v}})^{n}(\frac{\mu q}{\mathsf{v}})^{n-\mu}}{n!(n-\mu)!}$
$=\frac{1}{\sqrt{q}^{\mu}}\sum_{n=0}^{\infty}\frac{1}{n!(n-\mu)!}\Big{(}\frac{\mu\sqrt{q}}{\mathsf{v}}\Big{)}^{2n-\mu}$
$=\frac{1}{\sqrt{q}^{\mu}}I_{-\mu}\Big{(}\frac{2\sqrt{q}\mu}{\mathsf{v}}\Big{)}=\frac{1}{\sqrt{q}^{\mu}}I_{\mu}\Big{(}\frac{2\sqrt{q}\mu}{\mathsf{v}}\Big{)}.$
For $D_{\mu},\ \mu\geq-1$, let $n=m+\mu+1$,
$D_{\mu}$ $=\sum_{m=0}^{\infty}\frac{(\frac{\mu}{\mathsf{v}})^{m+\mu+1}(\frac{\mu q}{\mathsf{v}})^{m}}{(m+\mu+1)!m!}$
$=\frac{1}{\sqrt{q}^{\mu+1}}\sum_{m=0}^{\infty}\frac{1}{(m+\mu+1)!m!}\Big{(}\frac{\mu\sqrt{q}}{\mathsf{v}}\Big{)}^{2m+\mu+1}$
$=\frac{1}{\sqrt{q}^{\mu+1}}I_{\mu+1}\Big{(}\frac{2\sqrt{q}\mu}{\mathsf{v}}\Big{)}.$
For $D_{\mu},\ \mu<-1$, let $m=n-\mu-1$,
$D_{\mu}$ $=\sum_{n=0}^{\infty}\frac{(\frac{\mu}{\mathsf{v}})^{n}(\frac{\mu q}{\mathsf{v}})^{n-\mu-1}}{n!(n-\mu-1)!}$
$=\frac{1}{\sqrt{q}^{\mu+1}}\sum_{n=0}^{\infty}\frac{1}{n!(n-\mu-1)!}\Big{(}\frac{\mu\sqrt{q}}{\mathsf{v}}\Big{)}^{2n-\mu-1}$
$=\frac{1}{\sqrt{q}^{\mu+1}}I_{-\mu-1}\Big{(}\frac{2\sqrt{q}\mu}{\mathsf{v}}\Big{)}=\frac{1}{\sqrt{q}^{\mu+1}}I_{\mu+1}\Big{(}\frac{2\sqrt{q}\mu}{\mathsf{v}}\Big{)}.$