### v2

# All genus open mirror symmetry for the projective line

Abstract.
We prove that the Chekhov-Eynard-Orantin recursion on the mirror curve of $\mathbb{P}^{1}$ encodes all genus equivariant open Gromov-Witten invariants of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$. This result can be viewed as an all genus equivariant open mirror symmetry for $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$.
1. Introduction
1.1. Historical background and motivation
1.1.1. All genus closed mirror symmetry for the projective line
Mirror symmetry is a duality from string theory originally discovered by physicists. It says two dual string theories – type IIA and type IIB – on different Calabi-Yau 3-folds give rise to the same physics. Mathematicians became interested in this relationship around 1990 when Candelas, de la Ossa, Green, and Parkes [7] obtained a conjectural formula of the number of rational curves of arbitrary degree in the quintic 3-fold by relating it to period integrals of the quintic mirror.
By late 1990s mathematicians had established the foundation of Gromov-Witten (GW) theory as a mathematical theory of A-model topological closed strings. In this context, the genus $g$ free energy of the topological A-model is defined as a generating function of genus $g$ Gromov-Witten invariants. The mathematical aspect of genus zero closed mirror symmetry has been well-studied in many cases. Givental [22] and Lian-Liu-Yau [29] independently proved the genus zero mirror formula for the quintic Calabi-Yau 3-fold $Q$; later they extended their results to Calabi-Yau complete intersections in projective toric manifolds [23, 30, 31]. The genus-zero mirror theorem for toric Deligne-Mumford stacks is proved in [13, 11].
The higher genus mirror symmetry is much more subtle. Bershadsky-Cecotti-Ooguri-Vafa (BCOV) conjectured the genus-one and genus-two mirror formulae for the quintic 3-fold [1]. Combining the techniques of BCOV, results of Yamaguchi-Yau [35], and boundary conditions, Huang-Klemm-Quackenbush [27] proposed a mirror conjecture on $F_{g}^{Q}$ up to $g=51$. The BCOV genus-one mirror formula was first proved by A. Zinger in [40] using genus-one reduced Gromov-Witten theory, and later reproved in [28, 12] via quasimap theory and in [9] via MSP theory. The BCOV genus-two mirror formula was proved by Guo-Janda-Ruan [25] and Chang-Guo-Li [8]. The Yamaguchi–Yau’s finite generation and the holomorphic anomaly equation for quintic 3-folds were proved in [8] and [26].
The all genus closed mirror symmetry for the projective line is easier to study and there is a nice mathematical theory for the higher genus B-model. P. Dunin-Barkowski, N. Orantin, S. Shadrin and L. Spitz [14] relate the Chekhov-Eynard-Orantin topological recursion to the higher genus closed Gromov-Witten invariants of $\mathbb{P}^{1}$, proving the Norbury-Scott conjecture [34]. In [19], the above result is generalized to the equivariant setting. The main result of [19] relates the higher genus equivariant descendant Gromov-Witten potentials of $\mathbb{P}^{1}$ to the oscillatory integrals of Chekhov-Eynard-Orantin invariants of the mirror curve. This result can be viewed as an all genus equivariant closed mirror symmetry for $\mathbb{P}^{1}$. By taking the nonequivariant limit, the Norbury-Scott conjecture is recovered.
1.1.2. Open Gromov-Witten theory for the projective line
The open Gromov-Witten invariants are more difficult to study than the closed case. The all genus $S^{1}$-equivariant open Gromov-Witten theory of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ is studied via coherent boundary conditions by Buryak-Netser Zernik-Pandharipande-Tessler in [6]. A localization formula is conjectured and proved in the genus zero case. The localization formula is crucial for the computation of open Gromov-Witten theory of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$. It is natural to ask whether one can extend the result in [19] to the open string sector. This will be introduced in Section 1.1.3.
1.1.3. All genus open mirror symmetry for the projective line
In this paper, we prove the all genus open mirror symmetry for $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$, extending the result in [19] to the open string sector. On A-model side, we will study the open Gromov-Witten invariants of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ from two different but related points of view. From the first point of view, we define the higher genus open Gromov-Witten invariants of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ via integration over the fixed locus of the $S^{1}$-action and consider the corresponding generating functions in Section 3. On B-model side, instead of taking the oscillatory integrals, we take the expansion of the Chekhov-Eynard-Orantin invariants under certain local coordinate at the puncture of the mirror curve. The first main theorem (Theorem 1.1) of this paper states that this expansion coincides with the generating function of higher genus open Gromov-Witten invariants of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ under the mirror map. The second main theorem (Theorem 1.2) generalizes the above result to the case when descendant insertions appear.
From the second point of view, we define the higher genus open Gromov-Witten invariants of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ as a graph sum in Section 6.2 and Section 6.3. The vertex factor in this graph sum formula is given by the open Gromov-Witten invariants via integration over the fixed locus defined in Section 3. In other words, the open Gromov-Witten invariants via integration over the fixed locus are the building blocks in the second point of view of the open Gromov-Witten invariants of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$. In [6], this graph sum formula is conjectured to have a geometric definition via coherent boundary conditions. So we call the second point of view the open Gromov-Witten invariants via coherent boundary conditions. The third main theorem (Theorem 1.3) of this paper is the all genus mirror symmetry for open Gromov-Witten invariants of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ via coherent boundary conditions.
We would like to remark that similar open mirror symmetry phenomenon appears in the case of toric Calabi-Yau 3-folds/3-orbifolds, where the open Gromov-Witten invariants are defined via integration over the fixed locus. Based on the work of Eynard-Orantin [15] and Mariño [33], Bouchard-Klemm-Mariño-Pasquetti (BKMP) [4, 5] proposed a new formalism of the topological B-model in terms of the Chekhov-Eynard-Orantin invariants of the mirror curve. BKMP conjectured a precise correspondence, known as the BKMP Remodeling Conjecture, between the local expansion of the Chekhov-Eynard-Orantin invariants at the puncture of the mirror curve and the generating function of open Gromov-Witten invariants of toric Calabi-Yau 3-folds/3-orbifolds. The mathematical study of the Remodeling Conjecture can be found, for example, in [3, 10, 16, 17, 18, 20, 21, 36, 37, 38, 39]. The difference in our case is that our target space $\mathbb{P}^{1}$ is compact and non-Calabi-Yau.
We emphasize the following feature of our main results. Since $\mathbb{P}^{1}$ and $\mathbb{R}\mathbb{P}^{1}$ are compact, the non-equivariant limit of the $S^{1}$-equivariant open Gromov-Witten invariants of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ defined in Section 6 is conjectured [6, Conjecture 1] to be equal to the non-equivariant open Gromov-Witten invariants of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$, defined geometrically via coherent boundary conditions. In general, the computation of higher genus open Gromov-Witten invariants is quite difficult. The main results of our paper provide an effective algorithm of computing higher genus open Gromov-Witten invariants of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ recursively.
1.2. Statement of the main result
In this paper, we denote the complex projective line by $\mathbb{P}^{1}$. Let $t\in T=S^{1}$ act on $\mathbb{P}^{1}$ by
$t\cdot[z_{1},z_{2}]=[tz_{1},t^{-1}z_{2}].$
Let $\mathbb{C}[\mathsf{v}]=H_{T}^{*}(\mathrm{point};\mathbb{C})$ be the $T$-equivariant cohomology of a point. Let $p_{1}=[1,0]$ and $p_{2}=[0,1]$ be the $T$ fixed points. The $T$-equivariant cohomology of $\mathbb{P}^{1}$ is given by
$H^{*}_{T}(\mathbb{P}^{1};\mathbb{C})=\mathbb{C}[H,\mathsf{v}]/\langle(H+\mathsf{v}/2)(H-\mathsf{v}/2)\rangle$
where $\deg H=\deg\mathsf{v}=2$. Let
$L:=\{[e^{{\rm i}\varphi},e^{-{\rm i}\varphi}]\in\mathbb{P}^{1}:\varphi\in\mathbb{R}\}$
be the Lagrangian submanifold of $\mathbb{P}^{1}$, which is preserved by the $T$-action. By taking a Möbius transform, we can identify the pair $(\mathbb{P}^{1},L)$ with $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$.
In Section 3.5, we will define the generating function $F_{g,n}(\mathbf{t},Q;X_{1},\dots,X_{n})$ of genus $g$, $n$ boundary circles open Gromov-Witten invariants of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ via integration over the fixed locus of the $S^{1}$-action. Here $\mathbf{t}=t^{0}1+t^{1}H$, $Q$ is the Novikov variable encoding the degree of the stable maps to $\mathbb{P}^{1}$, and $X_{1},\cdots,X_{n}$ are variables encoding the winding numbers (viewed as the open string coordinates). We will also define the generating function $\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\cdots\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{g,m,n}^{\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1},T}(\mathbf{t},X_{1},\dots,X_{n})$ of genus $g$, $m$ descendant insertion classes, $n$ boundary circles open Gromov-Witten invariants of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$.
In Section 4.1, we will study the mirror curve of $\mathbb{P}^{1}$. Consider the $T$-equivariant superpotential $W:\mathbb{C}^{*}\rightarrow\mathbb{C}$ defined as
$W(Y)=t^{0}+\frac{\mathsf{v}}{2}\log q+Y+\frac{q}{Y}-\mathsf{v}\log Y,$
where $q=e^{t^{1}}$. Let $X=e^{-x/\mathsf{v}}$, $Y=e^{y}$. Consider the mirror curve $C_{q}$ defined as follows:
$C_{q}=\{(x,y)\in\mathbb{C}^{2}:x=W(e^{y})\}.$
In Section 4.3, we will study the Chekhov-Eynard-Orantin invariants $\omega_{g,n}$ of the mirror curve $C_{q}$. Then we use $\omega_{g,n}$ to define the B-model open primary potential $W_{g,n}(t^{0},q;X_{1},\dots,X_{n})$. In Section 4.4, we study the oscillatory integrals of $\omega_{g,n}$ and define the B-model open descendant potential $W_{g,m,n}(t^{0},q;\mathcal{E}_{1},\dots,\mathcal{E}_{m};X_{1},\dots,X_{n})$, where $\mathcal{E}_{1},\dots,\mathcal{E}_{m}\in K_{T}(\mathbb{P}^{1})$.
The following three theorems are the main results of this paper:
Theorem 1.1 (= Theorem 5.7).
For any $n>0$ and $g\geq 0$, we have
$F_{g,n}(\mathbf{t},1;X_{1},\dots,X_{n})=(-1)^{g-1}W_{g,n}(t^{0},q;X_{1},\dots,X_{n}).$
Theorem 1.2 (= Theorem 5.8).
Suppose $g,m,n\geq 0$, for any $\mathcal{E}_{1},\dots,\mathcal{E}_{m}\in K_{T}(\mathbb{P}^{1})$, we have
$W_{g,m,n}(t^{0},q;\mathcal{E}_{1},\dots,\mathcal{E}_{m};X_{1},\dots,X_{n})$
$=(-1)^{g-1}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\frac{\kappa_{z_{1}}(\mathcal{E}_{1})}{z_{1}-\psi_{1}},\dots,\frac{\kappa_{z_{m}}(\mathcal{E}_{m})}{z_{m}-\psi_{m}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{g,m,n}^{\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1},T}(\mathbf{t},X_{1},\dots,X_{n}).$
Here the characteristic class $\kappa_{z}$ is defined in Section 4.4 and $z_{1},\cdots,z_{m}$ are formal variables. The dependence of $W_{g,m,n}$ on $z_{1},\cdots,z_{m}$ is via oscillatory integrals defined in Section $4.$
In Section 6.1 and 6.2, we will study the moduli specification $S$ and the combinatorial graph sum formula of open Gromov-Witten invariants $\text{OGW}(S,\vec{a},\vec{\gamma})$, where the vertex factor is given by the open Gromov-Witten invariants via integration over the fixed locus defined in Section 3. In [6], $\text{OGW}(S,\vec{a},\vec{\gamma})$ is conjectured to have a geometric definition via coherent boundary conditions. We will define the decorated graph $\Gamma_{S}$ and the generating function $\text{OGW}(\Gamma_{S},\vec{a},\vec{\gamma},\mathbf{t})$ in Section 6.3. In Section 6.4, we define $\text{OGW}_{g,m,n}(\Gamma_{S},\frac{\kappa_{z_{i}}(\mathcal{E}_{i})}{z_{i}-\psi_{i}},\mathbf{t})$ as a function of $z_{i}$ composed of $\text{OGW}(\Gamma_{S},\vec{a},\vec{\gamma},\mathbf{t})$. The B-model potential $G_{g,m,n}(\Gamma_{S},\mathbf{t})$ is defined in Section 6.4 via graph sum formula, where the building blocks are Chekhov-Eynard-Orantin invariants. The third main theorem is the mirror symmetry for $\text{OGW}_{g,m,n}(\Gamma_{S},\frac{\kappa_{z_{i}}(\mathcal{E}_{i})}{z_{i}-\psi_{i}},\mathbf{t})$.
Theorem 1.3 (= Theorem 6.9).
Let $S$ be a pure connected moduli specification with one black vertex, $n$ white vertices and $m$ marked labels $\sf I$}. Then we have
$\text{OGW}_{g,m,n}(\Gamma_{S},\frac{\kappa_{z_{i}}(\mathcal{E}_{i})}{z_{i}-\psi_{i}},\mathbf{t})=G_{g,m,n}(\Gamma_{S},\mathbf{t}).$
1.3. Overview of the paper
In Section 2, we study the equivariant closed Gromov-Witten theory of $\mathbb{P}^{1}$ and give the graph sum formula for the descendant Gromov-Witten potential. In Section 3, we define the equivariant open Gromov-Witten theory of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ via integration over the fixed locus of the $S^{1}$-action and define the A-model potential $F_{g,n}$ and $\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\frac{\kappa_{z_{1}}(\mathcal{E}_{1})}{z_{1}-\psi_{1}},\dots,\frac{\kappa_{z_{m}}(\mathcal{E}_{m})}{z_{m}-\psi_{m}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{g,m,n}^{\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1},T}$. In Section 4, we study the mirror curve of $\mathbb{P}^{1}$ and the Chekhov-Eynard-Orantin topological recursion. Then we use the Chekhov-Eynard-Orantin invariants $\omega_{g,n}$ to define the B-model potential $W_{g,m,n}$. In Section 5, we prove the all genus open mirror symmetry for $F_{g,n}$ and $\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\frac{\kappa_{z_{1}}(\mathcal{E}_{1})}{z_{1}-\psi_{1}},\dots,\frac{\kappa_{z_{m}}(\mathcal{E}_{m})}{z_{m}-\psi_{m}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{g,m,n}^{\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1},T}$. In Section 6, we consider the open Gromov-Witten invariants $\text{OGW}(S,\vec{a},\vec{\gamma})$ and prove the corresponding all genus open mirror symmetry.
Acknowledgements
The authors would like to thank Bohan Fang, Chiu-Chu Melissa Liu, and Song Yu for useful discussions. The second author is partially supported by the Natural Science Foundation of Beijing, China (grant No. 1252008) and NSFC (grant No. 12571067).
2. Equivariant closed Gromov-Witten theory of $\mathbb{P}^{1}$
2.1. Equivariant cohomology of $\mathbb{P}^{1}$
Let $t\in T=S^{1}$ act on $\mathbb{P}^{1}$ by
$t\cdot[z_{1},z_{2}]=[tz_{1},t^{-1}z_{2}].$
The $T$-equivariant cohomology of $\mathbb{P}^{1}$ is given by
$H^{*}_{T}(\mathbb{P}^{1};\mathbb{C})=\mathbb{C}[H,\mathsf{v}]/\langle(H+\mathsf{v}/2)(H-\mathsf{v}/2)\rangle.$
Let $p_{1}=[1,0]$ and $p_{2}=[0,1]$ be the $T$-fixed points. Then $H|_{p_{1}}=-\mathsf{v}/2$, $H|_{p_{2}}=\mathsf{v}/2$. The $T$-equivariant Poincaré dual of $p_{1}$ and $p_{2}$ are $H-\mathsf{v}/2$ and $H+\mathsf{v}/2$, respectively.
Let
$\phi_{1}:=-\frac{H-\mathsf{v}/2}{\mathsf{v}},\phi_{2}:=\frac{H+\mathsf{v}/2}{\mathsf{v}}\in H^{*}_{T}(\mathbb{P}^{1};\mathbb{C})\otimes_{\mathbb{C}[\mathsf{v}]}\mathbb{C}[\mathsf{v},\frac{1}{\mathsf{v}}].$
We have
$\phi_{1}+\phi_{2}=1,\quad\phi_{\alpha}\cup\phi_{\beta}=\delta_{\alpha\beta}\phi_{\alpha},$
$(\phi_{\alpha},\phi_{\beta})_{\mathbb{P}^{1},T}:=\int_{\mathbb{P}^{1}}\phi_{\alpha}\cup\phi_{\beta}=\delta_{\alpha\beta}\int_{\mathbb{P}^{1}}\phi_{\alpha}=\frac{\delta_{\alpha\beta}}{\Delta^{\alpha}},\quad\alpha,\beta\in\{1,2\},$
where
$\Delta^{1}=-\mathsf{v},\quad\Delta^{2}=\mathsf{v}.$
For convenience, we introduce the notations $R_{T}:=\mathbb{C}[\mathsf{v}]$ and ${S}_{T}:=\mathbb{C}[\mathsf{v},\frac{1}{\mathsf{v}}]$, and we let $\bar{S}_{T}$ be the minimal field extension of $S_{T}$ containing $\{\sqrt{\Delta^{\alpha}}:\alpha=1,2\}$. Then $\{\phi_{\alpha}:\alpha=1,2\}$ is a canonical basis of the semisimple Frobenius algebra
$(H^{*}_{T}(\mathbb{P}^{1};\mathbb{C})\otimes_{R_{T}}\bar{S}_{T},\cup,(\cdot,\cdot)_{\mathbb{P}^{1},T})$
over $\bar{S}_{T}$. Let
$\hat{\phi}_{\alpha}:=\sqrt{\Delta^{\alpha}}\phi_{\alpha}.$
Then $\{\hat{\phi}_{\alpha}:\alpha=1,2\}$ is the normalized canonical basis of the semisimple Frobenius algebra $H^{*}_{T}(\mathbb{P}^{1};\mathbb{C})\otimes_{R_{T}}\bar{S}_{T}$:
$\hat{\phi}_{\alpha}\cup\hat{\phi}_{\beta}=\delta_{\alpha\beta}\sqrt{\Delta^{\alpha}}\hat{\phi}_{\alpha},\quad(\hat{\phi}_{\alpha},\hat{\phi}_{\beta})_{\mathbb{P}^{1},T}=\delta_{\alpha\beta}.$
2.2. Equivariant Gromov-Witten invariants of $\mathbb{P}^{1}$
Let $E(\mathbb{P}^{1})$ denote the effective curve classes in $H_{2}(\mathbb{P}^{1};\mathbb{Z})$. Given nonnegative integers $g,n$ and an effective curve class $\beta\in E(\mathbb{P}^{1})$, let $\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{1},\beta)$ be the moduli stack of genus $g$, $n$-pointed, degree $\beta$ stable maps to $\mathbb{P}^{1}$. Let $\mathrm{ev}_{i}:\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{1},\beta)\rightarrow\mathbb{P}^{1}$ be the evaluation map at the $i$-th marked point. The $T$-action on $\mathbb{P}^{1}$ induces an $T$-action on $\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{1},\beta)$ and the evaluation map $\mathrm{ev}_{i}$ is $T$-equivariant.
For $i=1,\dots,n$, let $\mathbb{L}_{i}$ be the $i$-th tautological line bundle over $\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{1},\beta)$ formed by the cotangent line at the $i$-th marked point. Define the $i$-th descendant class $\psi_{i}$ as
$\psi_{i}:=c_{1}(\mathbb{L}_{i})\in H^{2}(\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{1},\beta);\mathbb{Q}).$
Given $\gamma_{1},\dots,\gamma_{n}\in H^{*}_{T}(\mathbb{P}^{1};\mathbb{C})$ and nonnegative integers $a_{1},\dots,a_{n}$, we define genus-$g$, degree-$\beta$, $T$-equivariant descendant Gromov-Witten invariants of $\mathbb{P}^{1}$:
$\langle\tau_{a_{1}}(\gamma_{1})\dots\tau_{a_{n}}(\gamma_{n})\rangle_{g,n,\beta}^{\mathbb{P}^{1},T}:=\int_{[\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{1},\beta)]^{\mathrm{vir}}}\prod_{i=1}^{n}\psi_{i}^{a_{i}}\mathrm{ev}_{i}^{*}(\gamma_{i})\in\mathbb{C}[\mathsf{v}].$
The genus-$g$, degree-$\beta$, $T$-equivariant primary Gromov-Witten invariants of $\mathbb{P}^{1}$ is defined as
$\langle\gamma_{1}\dots\gamma_{n}\rangle_{g,n,\beta}^{\mathbb{P}^{1},T}:=\langle\tau_{0}(\gamma_{1})\dots\tau_{0}(\gamma_{n})\rangle_{g,n,\beta}^{\mathbb{P}^{1},T}.$
Define the Novikov ring
$\Lambda_{\mathrm{nov}}:=\mathbb{C}[\widehat{E(\mathbb{P}^{1})}]=\Bigg\{\sum_{\beta\in E(\mathbb{P}^{1})}c_{\beta}Q^{\beta}:c_{\beta}\in\mathbb{C}\Bigg\},$
where $Q$ is the Novikov variable. Let $\mathbf{t}=t^{0}1+t^{1}H$, we define the following double correlator:
$\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\tau_{a_{1}}(\gamma_{1}),\dots,\tau_{a_{n}}(\gamma_{n})\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{g,n}^{\mathbb{P}^{1},T}:=\sum_{\beta\in E(\mathbb{P}^{1})}\sum_{m=0}^{\infty}\frac{Q^{\beta}}{m!}\langle\tau_{a_{1}}(\gamma_{1}),\dots,\tau_{a_{n}}(\gamma_{n}),\mathbf{t}^{m}\rangle_{g,n+m,\beta}^{\mathbb{P}^{1},T}.$
For $j=1,\dots,n$, introduce formal variables
${\bf{u}}_{j}={\bf{u}}_{j}(z)=\sum_{a\geq 0}(u_{j})_{a}z^{a}$
where $(u_{j})_{a}\in H^{*}_{T}(\mathbb{P}^{1})\otimes_{R_{T}}\bar{S}_{T}$. Define
$\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}{\bf u}_{1},\dots,{\bf u}_{n}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{g,n}^{\mathbb{P}^{1},T}=\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}{\bf u}_{1}(\psi),\dots,{\bf u}_{n}(\psi)\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{g,n}^{\mathbb{P}^{1},T}=\sum_{a_{1},\dots,a_{n}\geq 0}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}(u_{1})_{a_{1}}\psi^{a_{1}},\dots,(u_{n})_{a_{n}}\psi^{a_{n}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{g,n}^{\mathbb{P}^{1},T}.$
Let $z_{1},\dots,z_{n}$ be formal variables and $\gamma_{1},\dots,\gamma_{n}\in H^{*}_{T}(\mathbb{P}^{1})\otimes_{R_{T}}\bar{S}_{T}$. Define
$\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\frac{\gamma_{1}}{z_{1}-\psi},\dots,\frac{\gamma_{n}}{z_{n}-\psi}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{g,n}^{\mathbb{P}^{1},T}=\sum_{a_{1},\dots,a_{n}\in\mathbb{Z}_{\geq 0}}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\gamma_{1}\psi^{a_{1}},\dots,\gamma_{n}\psi^{a_{n}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{g,n}^{\mathbb{P}^{1},T}\prod_{i=1}^{n}z_{i}^{-a_{i}-1}.$
We use the conventions that
$\langle\frac{\gamma}{z-\psi}\rangle_{0,1,0}^{\mathbb{P}^{1},T}:=z\int_{\mathbb{P}^{1}}\gamma,$
$\langle\frac{\gamma_{1}}{z-\psi},\gamma_{2}\rangle_{0,2,0}^{\mathbb{P}^{1},T}:=\int_{\mathbb{P}^{1}}\gamma_{1}\cup\gamma_{2},$
$\langle\frac{\gamma_{1}}{z_{1}-\psi_{1}},\frac{\gamma_{2}}{z_{2}-\psi_{2}}\rangle_{0,2,0}^{\mathbb{P}^{1},T}:=\frac{1}{z_{1}+z_{2}}\int_{\mathbb{P}^{1}}\gamma_{1}\cup\gamma_{2}.$
2.3. Equivariant quantum cohomology and Frobenius structure
The $T$-equivariant quantum cohomology ring $QH_{T}^{*}(\mathbb{P}^{1})$ of $\mathbb{P}^{1}$ is defined by its genus-zero primary Gromov-Witten invariants. As a $\mathbb{C}[\mathsf{v}]$-module, $QH_{T}^{*}(\mathbb{P}^{1})=H^{*}_{T}(\mathbb{P}^{1})$. The ring structure is given by the quantum product $\star$:
$(\gamma_{1}\star\gamma_{2},\gamma_{3})_{\mathbb{P}^{1},T}=\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\gamma_{1},\gamma_{2},\gamma_{3}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{0,3}^{\mathbb{P}^{1},T}.$
In other words,
$\gamma_{1}\star\gamma_{2}=\gamma_{1}\cup\gamma_{2}+q\left(\int_{\mathbb{P}^{1}}\gamma_{1}\right)\left(\int_{\mathbb{P}^{1}}\gamma_{2}\right),$
where $\cup$ is the cup product in $H^{*}_{T}(\mathbb{P}^{1})$ and $q=Qe^{t^{1}}$. In particular,
$H\star H=\mathsf{v}^{2}/4+q.$
The $T$-equivariant quantum cohomology ring of $\mathbb{P}^{1}$ is given by
$QH^{*}_{T}(\mathbb{P}^{1};\mathbb{C})=\mathbb{C}[H,\mathsf{v},q]/\langle(H+\mathsf{v}/2)\star(H-\mathsf{v}/2)-q\rangle,$
where $\deg H=\deg\mathsf{v}=2$ and $\deg q=4$.
Let
$\phi_{1}(q)=\frac{1}{2}-\frac{H}{\mathsf{v}\sqrt{1+4q/\mathsf{v}^{2}}},\quad\phi_{2}(q)=\frac{1}{2}+\frac{H}{\mathsf{v}\sqrt{1+4q/\mathsf{v}^{2}}}.$
Then
$\phi_{\alpha}(q)\star\phi_{\beta}(q)=\delta_{\alpha\beta}\phi_{\alpha}(q),\quad(\phi_{\alpha}(q),\phi_{\beta}(q))_{\mathbb{P}^{1},T}=\frac{\delta_{\alpha\beta}}{\Delta^{\alpha}(q)},$
where
$\Delta^{1}(q)=-\mathsf{v}\sqrt{1+4q/\mathsf{v}^{2}},\quad\Delta^{2}(q)=-\Delta^{1}(q).$
Let $S:=\bar{S}_{T}[[q]]$. Then $\{\phi_{1}(q),\phi_{2}(q)\}$ is a canonical basis of the semisimple Frobenius algebra
$(H^{*}_{T}(\mathbb{P}^{1};\mathbb{C})\otimes_{R_{T}}S,\star,(\cdot,\cdot)_{\mathbb{P}^{1},T}).$
The normalized canonical basis is defined by
$\{\hat{\phi}_{\alpha}(q):=\sqrt{\Delta^{\alpha}(q)}\phi_{\alpha}(q):\alpha=1,2\}.$
They satisfy
$\hat{\phi}_{\alpha}(q)\star\hat{\phi}_{\beta}(q)=\delta_{\alpha\beta}\sqrt{\Delta^{\alpha}(q)}\hat{\phi}_{\alpha}(q),\quad(\hat{\phi}_{\alpha}(q),\hat{\phi}_{\beta}(q))_{\mathbb{P}^{1},T}=\delta_{\alpha\beta}.$
2.4. Canonical coordinates and the transition matrix
Let $\{t^{0},t^{1}\}$ be the flat coordinates with respect to the basis $\{T_{0}=1,T_{1}=H\}$, and let $\{u^{1},u^{2}\}$ be the canonical coordinates with respect to the basis $\{\phi_{1}(q),\phi_{2}(q)\}$. Then
$\frac{\partial}{\partial u^{1}}$ $=\frac{1}{2}\frac{\partial}{\partial t^{0}}+\frac{1}{\Delta^{1}(q)}\frac{\partial}{\partial t^{1}},$
$\frac{\partial}{\partial u^{2}}$ $=\frac{1}{2}\frac{\partial}{\partial t^{0}}+\frac{1}{\Delta^{2}(q)}\frac{\partial}{\partial t^{1}}.$
The canonical coordinates $u^{1}$ and $u^{2}$ are characterized by the above equations up to a constant in $\mathbb{C}[\mathsf{v},\frac{1}{\mathsf{v}}]$. We choose the canonical coordinates $(u^{1},u^{2})$ such that
$\lim_{q\rightarrow 0}(u^{1}-t^{0}+\mathsf{v}t^{1}/2)=0,\quad\lim_{q\rightarrow 0}(u^{2}-t^{0}-\mathsf{v}t^{1}/2)=0.$
We use Greek alphabet $\alpha\in\{1,2\}$ to represent canonical coordinates and the Latin alphabet $i\in\{0,1\}$ to represent flat coordinates. Let $\Psi=(\Psi_{i}^{\ \alpha})$ be the transition matrix between the flat and quantum normalized canonical bases:
$T_{i}=\sum_{\alpha\in\{1,2\}}\Psi_{i}^{\ \alpha}\hat{\phi}_{\alpha}(q).$
It can be rewritten as
$\frac{\partial}{\partial t^{i}}=\sum_{\alpha\in\{1,2\}}\Psi_{i}^{\ \alpha}\sqrt{\Delta^{\alpha}(q)}\frac{\partial}{\partial u^{\alpha}},$
which is equivalent to
$\frac{du^{\alpha}}{\sqrt{\Delta^{\alpha}(q)}}=\sum_{i\in\{0,1\}}dt^{i}\Psi_{i}^{\ \alpha}.$
Then
$\Psi_{0}^{\ \alpha}=\frac{1}{\sqrt{\Delta^{\alpha}(q)}},\quad\Psi_{1}^{\ \alpha}=\frac{\sqrt{\Delta^{\alpha}(q)}}{2}.$
Let $\Psi^{-1}=(\Psi^{-1})_{\alpha}^{\ i}$, so that
$\sum_{i\in\{0,1\}}(\Psi^{-1})_{\alpha}^{\ i}\Psi_{i}^{\ \beta}=\delta_{\alpha}^{\ \beta}.$
Then
$(\Psi^{-1})_{\alpha}^{\ 0}=\frac{\sqrt{\Delta^{\alpha}(q)}}{2},\quad(\Psi^{-1})_{\alpha}^{\ 1}=\frac{1}{\sqrt{\Delta^{\alpha}(q)}}.$
2.5. The $\mathcal{S}$-operator
For any $a,b\in H^{*}_{T}(\mathbb{P}^{1};\bar{S}_{T})$, define the $\mathcal{S}$-operator by
$(a,\mathcal{S}(b))_{\mathbb{P}^{1},T}=(a,b)_{\mathbb{P}^{1},T}+\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}a,\frac{b}{z-\psi}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{0,2}^{\mathbb{P}^{1},T}.$
We consider several different (flat) bases for $H^{*}_{T}(\mathbb{P}^{1};\bar{S}_{T})$:
  1. (1)
    The natural basis: $T_{0}=1$ and $T_{1}=H$.
  2. (2)
    The basis dual to the natural basis: $T^{0}=H$ and $T^{1}=1$.
  3. (3)
    The classicial canonical basis $\{\phi_{1},\phi_{2}\}$.
  4. (4)
    The basis dual to the classicial canonical basis with respect to the $T$-equivariant Poincaré pairing: $\phi^{1}=\Delta^{1}\phi_{1}$, $\phi^{2}=\Delta^{2}\phi_{2}$.
  5. (5)
    The normalized classical basis $\hat{\phi}_{1}=\sqrt{\Delta^{1}}\phi_{1}$ and $\hat{\phi}_{2}=\sqrt{\Delta^{2}}\phi_{2}$, which is self-dual: $\hat{\phi}^{1}=\hat{\phi}_{1}$ and $\hat{\phi}^{2}=\hat{\phi}_{2}$.
We also consider several different bases for $H^{*}_{T}(\mathbb{P}^{1};\bar{S}_{T})\otimes_{\bar{S}_{T}}S$:
  1. (1)
    The quantum canonical basis $\{\phi_{1}(q),\phi_{2}(q)\}$.
  2. (2)
    The basis dual to the quantun canonical basis with respect to the $T$-equivariant Poincaré pairing: $\phi^{1}(q)=\Delta^{1}(q)\phi_{1}(q)$, $\phi^{2}(q)=\Delta^{2}(q)\phi_{2}(q)$.
  3. (3)
    The normalized quantum canonical basis $\{\hat{\phi}_{1}(q),\hat{\phi}_{2}(q)\}$, which is self-dual.
We introduce some additional notations. For $\alpha,\beta\in\{1,2\}$, define
$S^{\alpha}_{\ \beta}=(\phi^{\alpha},\mathcal{S}(\phi_{\beta}))_{\mathbb{P}^{1},T},\quad S^{\hat{\underline{\alpha}}}_{\ \beta}=(\hat{\phi}_{\alpha}(q),\mathcal{S}(\phi_{\beta}))_{\mathbb{P}^{1},T},\quad S^{\hat{\underline{\alpha}}}_{\ \hat{\underline{\beta}}}=(\hat{\phi}_{\alpha}(q),\mathcal{S}(\hat{\phi}_{\beta}(q)))_{\mathbb{P}^{1},T}.$
For $\alpha,\beta\in\{1,2\}$ and $i\in\{0,1\}$, define
$S_{i}^{\ \hat{\alpha}}=(T_{i},\mathcal{S}(\hat{\phi}^{\alpha}))_{\mathbb{P}^{1},T}.$
For $a,b\in H^{*}_{T}(\mathbb{P}^{1};\bar{S}_{T})$, define
$S_{z}(a,b)$ $=(a,\mathcal{S}(b))_{\mathbb{P}^{1},T}$
$V_{z_{1},z_{2}}(a,b)$ $=\frac{(a,b)_{\mathbb{P}^{1},T}}{z_{1}+z_{2}}+\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\frac{a}{z_{1}-\psi_{1}},\frac{b}{z_{2}-\psi_{2}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{0,2}^{\mathbb{P}^{1},T}.$
We have the following identity:
$V_{z_{1},z_{2}}(a,b)=\frac{1}{z_{1}+z_{2}}\sum_{\alpha\in\{1,2\}}S_{z_{1}}(\hat{\phi}_{\alpha}(q),a)S_{z_{2}}(\hat{\phi}_{\alpha}(q),b).$
Remark 2.1.
Let $t^{\prime}=t^{1}H$ and $t^{\prime\prime}=t^{0}1$. By the divisor equation, we have
$\ (a,b)_{\mathbb{P}^{1},T}+\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}a,\frac{b}{z-\psi}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{0,2}^{\mathbb{P}^{1},T}$
$=$ $\ (a,be^{t^{\prime}/z})_{\mathbb{P}^{1},T}+\sum_{m=0}^{\infty}\sum_{\beta\in E(\mathbb{P}^{1})\atop{(\beta,m)\neq(0,0)}}\frac{Q^{\beta}e^{\int_{\beta}t^{\prime}}}{m!}\langle a,\frac{be^{t^{\prime}/z}}{z-\psi},(t^{\prime\prime})^{m}\rangle^{\mathbb{P}^{1},T}_{0,2+m,\beta}.$
The factor $e^{\int_{\beta}t^{\prime}}$ plays the role of $Q^{\beta}$, so the operator $\mathcal{S}(b)\big|_{Q=1}$ is well-defined.
2.6. Equivariant $J$-function
The $T$-equivariant $J$-function $J(z)$ is characterized by
$(J(z),a)_{\mathbb{P}^{1},T}=(1,\mathcal{S}(a)\big|_{Q=1})_{\mathbb{P}^{1},T},$
for any $a\in H^{*}_{T}(\mathbb{P}^{1};\bar{S}_{T})$. Equivalently,
$J(z)=1+\sum_{\alpha\in\{1,2\}}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}1,\frac{\hat{\phi}_{\alpha}}{z-\psi}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{0,2}^{\mathbb{P}^{1},T}\Big|_{Q=1}\hat{\phi}_{\alpha}.$
By the genus zero mirror theorem [22, 29],
$J(z)=e^{(t^{0}+t^{1}H)/z}\Bigg(1+\sum_{d=1}^{\infty}\frac{q^{d}}{\prod_{m=1}^{d}(H+\mathsf{v}/2+mz)\prod_{m=1}^{d}(H-\mathsf{v}/2+mz)}\Bigg),$
where $q=e^{t^{1}}$.
Let $J(z)=J^{1}\phi_{1}+J^{2}\phi_{2}$. Then for $\alpha=1,2$, we have
(1) $J^{\alpha}$ $=e^{(t^{0}+t^{1}\Delta^{\alpha}/2)/z}\sum_{d=0}^{\infty}\frac{q^{d}}{d!z^{d}}\frac{1}{\prod_{m=1}^{d}(\Delta^{\alpha}+mz)}$
$=e^{(t^{0}+t^{1}\Delta^{\alpha}/2)/z}\sum_{m=0}^{\infty}\Big(\frac{\sqrt{q}}{z}\Big)^{2m}\frac{\Gamma(\Delta^{\alpha}/z+1)}{m!\Gamma(\Delta^{\alpha}/z+m+1)}$
$=e^{t^{0}/z}z^{\Delta^{\alpha}/z}\Gamma(\Delta^{\alpha}/z+1)I_{\Delta^{\alpha}/z}\Big(\frac{2\sqrt{q}}{z}\Big),$
where the function $I_{\alpha}(x)$ is the modified Bessel function of first kind in Appendix A.
2.7. The $R$-matrix
Let $U$ denote the diagnoal matrix $\text{diag}(u^{1},u^{2})$. The results in [24] imply the following statement.
Theorem 2.2.
There exists a unique matrix power series $R(z)=\mathbf{1}+R_{1}z+R_{2}z^{2}+\dots$ satisfying the following properties:
  1. (1)
    The entries of each $R_{k}$ lie in $\bar{S}_{T}[[q,t^{0}]]$.
  2. (2)
    $\tilde{S}=\Psi R(z)e^{U/z}$ is a fundamental solution to the $T$-equivariant big quantum differential equation.
  3. (3)
    $R$ satisfies the unitary condition $R^{T}(-z)R(z)=\mathbf{1}$.
  4. (4)
    For $\alpha,\beta\in\{1,2\}$, we have
    $\lim_{q,t^{0}\rightarrow 0}R_{\alpha}^{\ \beta}(z)=\delta_{\alpha\beta}\exp\left(-\sum_{n=1}^{\infty}\frac{B_{2n}}{2n(2n-1)}\left(\frac{z}{\Delta^{\beta}}\right)^{2n-1}\right).$
2.8. The graph sum formula for descendant Gromov-Witten potential
In this subsection, we introduce the graph sum formula for the generating functions of descendant Gromov-Witten invariants. Given a connected graph $\Gamma$, we introduce the following notation.
  1. (1)
    $V(\Gamma)$ is the set of vertices in $\Gamma$.
  2. (2)
    $E(\Gamma)$ is the set of edges in $\Gamma$.
  3. (3)
    $H(\Gamma)$ is the set of half-edges in $\Gamma$.
  4. (4)
    $L^{o}(\Gamma)$ is the set of ordinary leaves in $\Gamma$. The ordinary leaves are ordered: $L^{o}(\Gamma)=\{l_{1},\dots,l_{n}\}$ where $n$ is the number of ordinary leaves.
  5. (5)
    $L^{1}(\Gamma)$ is the set of dilaton leaves in $\Gamma$. The dilaton leaves are unordered.
Withe the above notation, we introduce the following labels:
  1. (1)
    (genus) $g:V(\Gamma)\rightarrow\mathbb{Z}_{\geq 0}$.
  2. (2)
    (marking) $\beta:V(\Gamma)\rightarrow\{1,2\}$. This induces $\beta:L(\Gamma)=L^{o}(\Gamma)\cup L^{1}(\Gamma)\rightarrow\{1,2\}$, as follows: if $l\in L(\Gamma)$ is a leaf attached to a vertex $v\in V(\Gamma)$, define $\beta(l)=\beta(v)$.
  3. (3)
    (height) $k:H(\Gamma)\rightarrow\mathbb{Z}_{\geq 0}$.
Given an edge $e$, let $h_{1}(e)$ and $h_{2}(e)$ be the two half-edges associated to $e$. The order of the two half-edges does not affect the graph sum formula in this paper. Given a vertex $v\in V(\Gamma)$, let $H(v)$ denote the set of half-edges emanating from $v$. The valency of the vertex $v$ is equal to the cardinality of the set $H(v)$: ${\mathrm{val}}(v)=|H(v)|$. A labeled graph $\vec{\Gamma}=(\Gamma,g,\beta,k)$ is stable if
$2g(v)-2+{\mathrm{val}}(v)>0$
for all $v\in V(\Gamma)$.
Let ${\bf\Gamma}(\mathbb{P}^{1})$ denote the set of all stable labeled graphs $\vec{\Gamma}=(\Gamma,g,\beta,k)$. The genus of a stable labeled graph $\vec{\Gamma}$ is defined to be
$g(\vec{\Gamma}):=\sum_{v\in V(\Gamma)}g(v)+|E(\Gamma)|-|V(\Gamma)|+1=\sum_{v\in V(\Gamma)}(g(v)-1)+(\sum_{e\in E(\Gamma)}1)+1.$
Define
${\bf\Gamma}_{g,n}(\mathbb{P}^{1})=\{\vec{\Gamma}=(\Gamma,g,\beta,k)\in{\bf\Gamma}(\mathbb{P}^{1}):g(\vec{\Gamma})=g,|L^{o}(\Gamma)|=n\}.$
We assign weights to leaves, edges, and vertices of a labeled graph $\vec{\Gamma}\in{\bf\Gamma}(\mathbb{P}^{1})$ as follows.
  1. (1)
    Ordinary leaves. To each ordinary leaf $l_{j}\in L^{o}(\Gamma)$ with $\beta(l_{j})=\beta\in\{1,2\}$ and $k(l_{j})=k\in\mathbb{Z}_{\geq 0}$, we assign the following descendant weight:
    $(\mathcal{L}^{\bf u})^{\beta}_{k}(l_{j})=[z^{k}](\sum_{\alpha,\gamma\in\{1,2\}}\left(\frac{{\bf u}^{\alpha}_{j}(z)}{\sqrt{\Delta^{\alpha}(q)}}S^{\hat{\underline{\gamma}}}_{\ \hat{\underline{\alpha}}}(z)\right)_{+}R(-z)_{\gamma}^{\ \beta}).$
    where $(\cdot)_{+}$ means taking the nonnegative powers of $z$.
  2. (2)
    Dilaton leaves. To each dilaton leaf $l\in L^{1}(\Gamma)$ with $\beta(l)=\beta\in\{1,2\}$ and $2\leq k(l)=k\in\mathbb{Z}_{\geq 0}$, we assign
    $(\mathcal{L}^{1})^{\beta}_{k}(l)=[z^{k-1}](-\sum_{\alpha\in\{1,2\}}\frac{1}{\sqrt{\Delta^{\alpha}(q)}}R_{\alpha}^{\ \beta}(-z)).$
  3. (3)
    Edges. To an edge connecting a vertex marked by $\alpha\in\{1,2\}$ and a vertex marked by $\beta\in\{1,2\}$, and with heights $k$ and $l$ at the corresponding half-edges, we assign
    $\mathcal{E}^{\alpha,\beta}_{k,l}=[z^{k}w^{l}]\Big(\frac{1}{z+w}(\delta_{\alpha,\beta}-\sum_{\gamma\in\{1,2\}}R_{\gamma}^{\ \alpha}(-z)R_{\gamma}^{\ \beta}(-w))\Big).$
  4. (4)
    Vertices. To a vertex $v$ with genus $g(v)=g\in\mathbb{Z}_{\geq 0}$ and with marking $\beta(v)=\beta$, with $n$ ordinary leaves and half-edges attached to it with heights $k_{1},\dots,k_{n}\in\mathbb{Z}_{\geq 0}$ and $m$ more dilaton leaves with heights $k_{n+1},\dots,k_{n+m}\in\mathbb{Z}_{\geq 0}$, we assign
    $\Big(\sqrt{\Delta^{\beta}(q)}\Big)^{2g(v)-2+{\mathrm{val}}(v)}\langle\tau_{k_{1}}\dots\tau_{k_{n+m}}\rangle_{g},$
    where $\langle\tau_{k_{1}}\dots\tau_{k_{n+m}}\rangle_{g}=\int_{\overline{\mathcal{M}}_{g,n+m}}\psi_{1}^{k_{1}}\dots\psi_{n+m}^{k_{n+m}}$.
We define the weight of a labeled graph $\vec{\Gamma}\in{\bf\Gamma}_{g,n}(\mathbb{P}^{1})$ to be
$\omega^{\bf u}_{A}(\vec{\Gamma})=$ $\prod_{v\in V(\Gamma)}\Big(\sqrt{\Delta^{\beta(v)}(q)}\Big)^{2g(v)-2+{\mathrm{val}}(v)}\langle\prod_{h\in H(v)}\tau_{k(h)}\rangle_{g(v)}\prod_{e\in E(\Gamma)}\mathcal{E}^{\beta(v_{1}(e)),\beta(v_{2}(e))}_{k(h_{1}(e)),k(h_{2}(e))}$
$\cdot\prod_{l\in L^{1}(\Gamma)}(\mathcal{L}^{1})^{\beta(l)}_{k(l)}(l)\prod_{j=1}^{n}(\mathcal{L}^{\bf u})^{\beta(l_{j})}_{k(l_{j})}(l_{j}).$
With the about definition of the weight of a labeled graph, we have the following theorem which expresses the $T$-equivariant descendant Gromov-Witten potential of $\mathbb{P}^{1}$ in terms of graph sum.
Theorem 2.3 ([24]).
Suppose that 2g2+n>02g-2+n>0. Then
$\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}{\bf u}_{1},\dots,\mathbf{u}_{n}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{g,n}^{\mathbb{P}^{1},T}=\sum_{\vec{\Gamma}\in{\bf\Gamma}_{g,n}(\mathbb{P}^{1})}\frac{\omega^{\mathbf{u}}_{A}(\vec{\Gamma})}{|\mathrm{Aut}(\vec{\Gamma})|}.$
3. Equivariant open Gromov-Witten theory of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ via integration over the fixed locus
3.1. Moduli spaces of stable maps to ($\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}$)
Let
$L:=\{[e^{{\rm i}\varphi},e^{-{\rm i}\varphi}]\in\mathbb{P}^{1}:\varphi\in\mathbb{R}\}$
be the Lagrangian submanifold of $\mathbb{P}^{1}$, which is preserved by the $T$-action. By taking a Möbius transform, we can identify the pair $(\mathbb{P}^{1},L)$ with $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$. Let $D_{1}$ and $D_{2}$ be the two hemispheres centered at $p_{1}$ and $p_{2}$ respectively. Then we have
$H_{2}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})=\mathbb{Z}[D_{1}]\oplus\mathbb{Z}[D_{2}].$
Sometimes, we identify the relative homology group $H_{2}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ to $\mathbb{Z}^{2}$ so that the generators $(1,0)$ and $(0,1)$ represent $D_{1}$ and $D_{2}$ respectively. The relative homology group $H_{2}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ could also be viewed as the extension of the homology groups $H_{2}(\mathbb{P}^{1})$ and $H_{1}(\mathbb{R}\mathbb{P}^{1})$. Consider the short exact sequence
[…]
For $d[\mathbb{P}^{1}]\in H_{2}(\mathbb{P}^{1})$, the map $\delta:H^{2}(\mathbb{P}^{1})\rightarrow H_{2}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ is defined by $\delta(d[\mathbb{P}^{1}])=(d,d)$. The connecting map $\partial:H_{2}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})\rightarrow H_{1}(\mathbb{R}\mathbb{P}^{1})$ is given by $\partial(a,b)=(b-a)[\mathbb{R}\mathbb{P}^{1}]$.
Let $(\Sigma,\partial\Sigma)$ be a prestable bordered Riemann surface, and let $\partial\Sigma=R_{1}\cup\dots\cup R_{h}$. The (small) genus $g(\Sigma)$ of $\Sigma$ is the genus of the closed Riemann surface obtained by capping each boundary component of $\Sigma$ with a disk. A genus-$g$ bordered Riemann surface with $h$ boundary components is called a Riemann surface of topological type $(g,h)$.
The double $\Sigma_{\mathbb{C}}$ of the prestable bordered Riemann surface $(\Sigma,\partial\Sigma)$ is obtained by gluing the surface $\Sigma$ and the conjugate surface $\bar{\Sigma}$ along the boundary $\partial\Sigma$ by the Schwartz reflection principle. The Riemann surface $\Sigma_{\mathbb{C}}$ is endowed with a canonical involution $b:\Sigma_{\mathbb{C}}\rightarrow\Sigma_{\mathbb{C}}$.
We concern about the maps $u:(\Sigma,\partial\Sigma)\rightarrow(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$. The degree of the map $u$ is the element
$\vec{d}=(d_{-},d_{+})=u_{*}[\Sigma]\in H_{2}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}).$
Let
$\mu_{i}[\mathbb{R}\mathbb{P}^{1}]=(u\mid_{R_{i}})_{*}[R_{i}]\in H_{1}(\mathbb{R}\mathbb{P}^{1})=\mathbb{Z}[\mathbb{R}\mathbb{P}^{1}],\quad i=1,\cdots,h.$
The number $\mu_{i}$ is called the $i$-th winding number. Let $\beta^{\prime}=u_{*}[\Sigma]\in H_{2}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ and let $\vec{\mu}=(\mu_{1},\dots,\mu_{h})\in\mathbb{Z}_{\neq 0}^{h}$. Then there exists $d\in\mathbb{Z}_{\geq 0}$ such that
$\beta^{\prime}=d[\mathbb{P}^{1}]+\sum_{i=1}^{h}\mu_{i}[\mathbb{R}\mathbb{P}^{1}].$
By the short exact sequence (2), we have
$d_{+}-d_{-}=\sum_{j=1}^{h}\mu_{j},$
$d=d_{+}-\sum_{\{j:\mu_{j}>0\}}\mu_{j}=d_{-}+\sum_{\{j:\mu_{j}<0\}}\mu_{j}.$
Let $(\Sigma,\partial\Sigma,x_{1},\dots,x_{n})$ be a prestable bordered Riemann surface with $n$ interior marked points, and consider stable maps:
$u:(\Sigma,\partial\Sigma,x_{1},\dots,x_{n})\rightarrow(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}).$
Let $\overline{\mathcal{M}}_{(g,h),n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})$ be the moduli space of degree $\beta^{\prime}$ stable maps to $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ from type $(g,h)$ bordered Riemann surfaces with $n$ interior marked points, such that the winding numbers are given by $\mu_{i}\in\mathbb{Z}$.
3.2. Equivariant open Gromov-Witten invariants
Let $\gamma_{1},\dots,\gamma_{n}\in H^{*}_{T}(\mathbb{P}^{1};\mathbb{C})$, $\beta^{\prime}\in H_{2}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$. The $T$-action on $\mathbb{P}^{1}$ induces a $T$-action on the moduli space $\overline{\mathcal{M}}_{(g,h),n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})$. Let $F:=\overline{\mathcal{M}}_{(g,h),n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})^{T}$ be the $T$-fixed locus. The evaluation map $\mathrm{ev}_{i}:\overline{\mathcal{M}}_{(g,h),n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})\to\mathbb{P}^{1}$ is $T$-equivariant.
For $i=1,\dots,n$, let $\mathbb{L}_{i}$ be the $i$-th tautological line bundle over $\overline{\mathcal{M}}_{(g,h),n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})$ formed by the cotangent line at the $i$-th marked point. Define the $i$-th descendant class $\psi_{i}$ as
$\psi_{i}:=c_{1}(\mathbb{L}_{i})\in H^{2}(\overline{\mathcal{M}}_{(g,h),n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu});\mathbb{Q}).$
We choose a $T$-equivariant lift $\psi_{i}^{T}\in H^{2}_{T}(\overline{\mathcal{M}}_{(g,h),n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu});\mathbb{Q})$ of $\psi_{i}$. We define the $T$-equivariant open Gromov-Witten invariants of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ via integration over the fixed locus:
$\langle\tau_{a_{1}}(\gamma_{1})\dots\tau_{a_{n}}(\gamma_{n})\rangle^{\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1},T}_{g,\beta^{\prime},\vec{\mu}}:=\int_{[F]^{\mathrm{vir}}}\frac{\prod_{i=1}^{n}(\psi_{i}^{T})^{a_{i}}\mathrm{ev}_{i}^{*}(\gamma_{i})|_{F}}{e_{T}(N^{\mathrm{vir}})}\in\mathbb{C}[\mathsf{v},\frac{1}{\mathsf{v}}],$
where $F]^{\mathrm{vir}}$ is the virtual fundamental class of $F$, and $N^{\mathrm{vir}}$ is the virtual normal bundle of $F$. Since $F$ is a compact orbifold without boundary, the above integral is well-defined.
3.3. Virtual localization formula
Definition 3.1 (Decorated graphs (a)).
Let $n\in\mathbb{Z}_{\geq 0}$ and $\beta^{\prime}=d[\mathbb{P}^{1}]+\sum_{i=1}^{h}\mu_{i}[\mathbb{R}\mathbb{P}^{1}]\in H_{2}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$. A genus $g$, $n$-pointed, degree $\beta^{\prime}$ decorated graph for $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ is a tuple $\vec{\Gamma}=(\Gamma,\vec{f},\vec{d},\vec{g},\vec{s})$ consisting of the following data.
  1. (1)
    $\Gamma$ is a compact, connected 1-dimensional CW complex. Let $V(\Gamma)\sqcup V_{\circ}(\Gamma)$ denote the set of vertices in $\Gamma$, denoted by $\bullet$ and $\circ$, respectively. The $\circ$ vertex is univalent. Let $E(\Gamma)$ denote the set of edges, where an edge $e$ is a line connecting two $\bullet$ vertices. Let $H(\Gamma)=\{h_{j}:j=1,\dots,h\}$ denote the set of half-edges, where a half-edge $h_{j}$ is a line connecting one $\bullet$ vertex and one $\circ$ vertex. We require the half-edges to be ordered. Let $F(\Gamma)$ be the set of flags:
    $\{(e,v)\in E(\Gamma)\times V(\Gamma):v\in e\}\cup\{(h_{j},v)\in H(\Gamma)\times V(\Gamma):v\in h_{j}\}.$
    For each $v\in V(\Gamma)$, let $F_{v}$ (resp. $E_{v}$, $H_{v}$) denote the flags (resp. edges, half-edges) attached to $v$, and let ${\mathrm{val}}(v)=|F_{v}|$ denote the number of flags incident to $v$.
  2. (2)
    The label map $\vec{f}:V(\Gamma)\rightarrow\{1,2\}$ labels each $\bullet$ with a number. If $v_{1},v_{2}\in V(\Gamma)$ are connected by an edge, we require $\vec{f}(v_{1})\neq\vec{f}(v_{2})$.
  3. (3)
    The degree map $\vec{d}:E(\Gamma)\cup H(\Gamma)\rightarrow\mathbb{Z}_{>0}$ sends an edge $e$ (resp. a half-edge $h_{j}$) to a positive integer $\vec{d}(e)=d_{e}$ (resp. $\vec{d}(h_{j})=d_{h_{j}}$).
  4. (4)
    The genus map $\vec{g}:V(\Gamma)\rightarrow\mathbb{Z}_{\geq 0}$ sends a vertex $v\in V(\Gamma)$ to a nonnegative integer $g_{v}$.
  5. (5)
    The marking map $\vec{s}:\{1,2,\dots,n\}\rightarrow V(\Gamma)$. For each $v\in V(\Gamma)$, define $S_{v}:=\vec{s}^{-1}(v)$, and $n_{v}=|S_{v}|$.
The data is required to satisfy the following conditions:
  • (i)
    (genus) $g=\sum_{v\in V(\Gamma)}g_{v}+|E(\Gamma)|-|V(\Gamma)|+1$.
  • (ii)
    (degree) $d=\sum_{e\in E(\Gamma)}d_{e}$.
  • (iii)
    (winding numbers) Let $h_{j}\in H(\Gamma)$ be the $j$-th half-edge, and let $v_{j}\in V(\Gamma)$ be its incident $\bullet$ vertex, then $\mu_{j}=(-1)^{\vec{f}(v_{j})}\vec{d}(h_{j})$.
Let $G_{g,n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})$ be the set of all decorated graphs $\vec{\Gamma}=(\Gamma,\vec{f},\vec{d},\vec{g},\vec{s})$ satisfying the above constraints.
Given a decorated graph $\vec{\Gamma}\in G_{g,n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})$. Let $V^{S}(\Gamma)$ be the set of stable vertices:
$V^{S}(\Gamma):=\{v\in V(\Gamma):2g_{v}-2+{\mathrm{val}}(v)+n_{v}>0\}.$
Let $\mathrm{Aut}(\vec{\Gamma})$ denote the group of automorphisms of $\vec{\Gamma}$, and let $A^{0}_{\vec{\Gamma}}$ be the group of covering automorphisms
$A^{0}_{\vec{\Gamma}}=\prod_{e\in E(\Gamma)}\mathbb{Z}_{\vec{d}(e)}\times\prod_{h_{j}\in H(\Gamma)}\mathbb{Z}_{\vec{d}(h_{j})}.$
The automorphism group $A_{\vec{\Gamma}}$ fits in the short exact sequence:
$1\rightarrow A^{0}_{\vec{\Gamma}}\rightarrow A_{\vec{\Gamma}}\rightarrow\mathrm{Aut}(\vec{\Gamma})\rightarrow 1.$
We set
$\overline{\mathcal{M}}_{\vec{\Gamma}}:=\prod_{v\in V^{S}(\Gamma)}\overline{\mathcal{M}}_{g_{v},F_{v}\cup S_{v}},\quad F_{\vec{\Gamma}}:=[\overline{\mathcal{M}}_{\vec{\Gamma}}/A_{\vec{\Gamma}}].$
Every decorated graph $\vec{\Gamma}\in G_{g,n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})$ represents a topological type of the $T$-invariant stable open map $u:(\Sigma,\partial\Sigma,x_{1},\dots,x_{n})\rightarrow(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ in the following way:
  1. (1)
    Every $v\in V(\Gamma)$ represents a stable curve or a single point $\Sigma_{v}$ in the domain curve $\Sigma$. If $\Sigma_{v}$ is a stable curve, it is of genus-$g_{v}$ with marked points $S_{v}$. The map $u$ contracts $\Sigma_{v}$ to the fixed-point $p_{\vec{f}(v)}$.
  2. (2)
    Every half-edge $h_{j}\in H(\Gamma)$ represents a disk $D_{j}$ in the domain curve $\Sigma$. The map $u|_{D_{j}}$ is a standard degree $\vec{d}(h_{j})=|\mu_{j}|$ cover of disk branched at $p_{\vec{f}(v_{j})}$, where $v_{j}$ is the $\bullet$ vertex attached at $h_{j}$.
  3. (3)
    Every edge $e\in E(\Gamma)$ represents a sphere $\mathbb{P}^{1}\cong C\subset\Sigma$, so that $u|_{C}$ is a degree $\vec{d}(e)=d_{e}$ cover of $\mathbb{P}^{1}$ branched at $p_{1},p_{2}$.
Vice versa, given a $T$-invariant stable open map $u$, one can read off its decorated graph $\vec{\Gamma}$. Therefore, we can use the decorated graphs to describe the connected components of the $T$-fixed locus of the moduli space $\overline{\mathcal{M}}_{(g,h),n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})$:
(3) $\overline{\mathcal{M}}_{(g,h),n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})^{T}=\bigcup_{\vec{\Gamma}\in G_{g,n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})}F_{\vec{\Gamma}}.$
[…]
Figure 1. The left figure is an example of the graph $\vec{\Gamma}$ $\in$ $G_{g,n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})$. The right figure is an example of stable map corresponding to $\vec{\Gamma}$.
Given $\vec{\Gamma}\in G_{g,n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})$, we introduce the following notations:
  • Let $J_{\pm}=\{j\in\{1,\dots,h\}:\pm\mu_{j}>0\}$.
  • (weight) We define
    ${\bf w}(p_{1})=-\mathsf{v},\quad{\bf w}(p_{2})=\mathsf{v},$
    For a flag $f\in F_{v}$, we define
    ${\bf w}_{f}:=\left\{\begin{aligned} &\frac{{\bf w}(p_{\vec{f}(v)})}{d_{e}},\quad f=(e,v),\\ &\frac{{\bf w}(p_{\vec{f}(v)})}{d_{h_{j}}},\quad f=(h_{j},v).\end{aligned}\right$.
  • (vertex contribution) Let $g\in\mathbb{Z}_{\geq 0}$,
    ${\bf h}(p_{i},g)=\frac{\Lambda^{\vee}_{g}({\bf w}(p_{i}))}{{\bf w}(p_{i})},$
    where $\Lambda^{\vee}_{g}(\mathsf{u}):=\sum_{i=0}^{g}(-1)^{i}\lambda_{i}\mathsf{u}^{g-i}$ and $\lambda_{i}$ is the $i$-th Hodge class on $\overline{\mathcal{M}}_{g,n}$.
  • (edge contribution) For $d\in\mathbb{Z}_{>0}$, we define
    ${\bf h}(e,d)=\frac{(-1)^{d}d^{2d}}{(d!)^{2}\mathsf{v}^{2d}}.$
  • (disk factor) For $\mu\in\mathbb{Z}_{>0}$, we define
    $D^{1}(\mu)=(-1)^{\mu+1}\frac{\mu^{\mu-2}}{\mu!\mathsf{v}^{\mu-2}},\quad D^{2}(\mu)=\frac{\mu^{\mu-2}}{\mu!\mathsf{v}^{\mu-2}}.$
By the virtual localization formula in [6], we get the following proposition.
Proposition 3.2.
Let $\beta^{\prime}=d[\mathbb{P}^{1}]+\sum_{i=1}^{h}\mu_{i}[\mathbb{R}\mathbb{P}^{1}]\in H_{2}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$. Then for $\gamma_{1},\dots,\gamma_{n}\in H^{*}_{T}(\mathbb{P}^{1})$ and $g,a_{1},\dots,a_{n}\in\mathbb{Z}_{\geq 0}$, we have
$\langle\tau_{a_{1}}(\gamma_{1}),\dots,\tau_{a_{n}}(\gamma_{n})\rangle^{\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1},T}_{g,\beta^{\prime},\vec{\mu}}$
$=$ $\sum_{\vec{\Gamma}\in G_{g,n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})}\frac{1}{|\mathrm{Aut}(\vec{\Gamma})|}\prod_{e\in E(\Gamma)}\frac{{\bf h}(e,d_{e})}{d_{e}}\prod_{v\in V(\Gamma)}\Big({\bf w}(p_{\vec{f}(v)})^{|E_{v}|}\prod_{i\in S_{v}}i^{*}_{p_{\vec{f}(v)}}\gamma_{i}\Big)$
$\cdot\prod_{j\in J_{-}}D^{1}(-\mu_{j})\prod_{j\in J_{+}}D^{2}(\mu_{j})\prod_{v\in V(\Gamma)}\int_{\overline{\mathcal{M}}_{g_{v},F_{v}\cup S_{v}}}\frac{{\bf h}(p_{\vec{f}(v)},g_{v})\prod_{i\in S_{v}}\psi_{i}^{a_{i}}}{\prod_{h_{j}\in H_{v}}{\bf w}_{(h_{j},v)}\prod_{f\in F_{v}}({\bf w}_{f}-\psi_{f})}.$
We use the following convention for the unstable integrals:
$\int_{\overline{\mathcal{M}}_{0,1}}\frac{1}{{\bf w}-\psi}={\bf w},\quad\int_{\overline{\mathcal{M}}_{0,2}}\frac{\psi_{2}^{a}}{{\bf w}-\psi_{1}}=(-{\bf w})^{a}$ $,\quad a\in\mathbb{Z}_{\geq 0},$
$\int_{\overline{\mathcal{M}}_{0,2}}\frac{1}{({\bf w}_{1}-\psi_{1})({\bf w}_{2}-\psi_{2})}=\frac{1}{{\bf w}_{1}+{\bf w}_{2}}$ .
3.4. Open/descendant correspondence
Given $u:(\Sigma,\partial\Sigma,{\bf x})\rightarrow(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ that represents a point $\xi\in\overline{\mathcal{M}}_{(g,h),n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})^{T}$, we have
$\Sigma=C\cup\bigcup_{j=1}^{h}D_{j},$
where $C$ is a closed nodal curve of genus $g$ with marked points $x_{1},\dots,x_{n}$, and $D_{j}$’s are holomorphic disks. $C$ and $D_{j}$ intersect at $y_{j}$. Let $\hat{u}=u|_{C}$ and $u_{j}=u|_{D_{j}}$. Then $\hat{u}:(C,{\bf x},{\bf y})\rightarrow\mathbb{P}^{1}$ represents a point $\hat{\xi}\in\overline{\mathcal{M}}_{g,n+h}(\mathbb{P}^{1},\beta)^{T}$, where the $(n+j)$-th marked point $x_{n+j}$ is $y_{j}$.
The following definition is introduced in [32, Definition 52].
Definition 3.3 (Decorated graphs (b)).
Define $G_{g,n+h}(\mathbb{P}^{1},\beta)$ to be the set of all decorated graphs $\hat{\Gamma}=(\Gamma,\vec{f},\vec{d},\vec{g},\vec{s})$ defined as follows. Let $n\in\mathbb{Z}_{\geq 0}$, $h\in\mathbb{Z}_{>0}$ and $\beta=d[\mathbb{P}^{1}]\in E(\mathbb{P}^{1})$. A genus $g$, $(n+h)$-pointed, degree $\beta$ decorated graph for $\mathbb{P}^{1}$ is a tuple $\hat{\Gamma}=(\Gamma,\vec{f},\vec{d},\vec{g},\vec{s})$ consisting of the following data.
  1. (1)
    $\Gamma$ is a compact, connected 1-dimensional CW complex. Let $V(\Gamma)$ denote the set of vertices in $\Gamma$, denoted by $\bullet$. Let $E(\Gamma)$ denote the set of edges, where an edge $e$ is a line connecting two $\bullet$ vertices. Let $F(\Gamma)$ be the set of flags:
    $\{(e,v)\in E(\Gamma)\times V(\Gamma):v\in e\}.$
    For each $v\in V(\Gamma)$, let $F_{v}$ (resp. $E_{v}$) denote the flags (resp. edges) attached to $v$, and let ${\mathrm{val}}(v)=|F_{v}|=|E_{v}|$ denote the number of flags (resp. edges) incident to $v$.
  2. (2)
    The label map $\vec{f}:V(\Gamma)\rightarrow\{1,2\}$ labels each $\bullet$ with a number. If $v_{1},v_{2}\in V(\Gamma)$ are connected by an edge, we require $\vec{f}(v_{1})\neq\vec{f}(v_{2})$.
  3. (3)
    The degree map $\vec{d}:E(\Gamma)\rightarrow\mathbb{Z}_{>0}$ sends an edge $e$ to a positive integer $\vec{d}(e)=d_{e}$.
  4. (4)
    The genus map $\vec{g}:V(\Gamma)\rightarrow\mathbb{Z}_{\geq 0}$ sends a vertex $v\in V(\Gamma)$ to a nonnegative integer $g_{v}$.
  5. (5)
    The marking map $\vec{s}:\{1,2,\dots,n+h\}\rightarrow V(\Gamma)$. For each $v\in V(\Gamma)$, define $S_{v}:=\vec{s}^{-1}(v)$, and $n_{v}=|S_{v}|$.
The data is required to satisfy the following conditions:
  • (i)
    (genus) $g=\sum_{v\in V(\Gamma)}g_{v}+|E(\Gamma)|-|V(\Gamma)|+1$.
  • (ii)
    (degree) $d=\sum_{e\in E(\Gamma)}d_{e}$.
Let $F_{\hat{\Gamma}}$ be the connected component corresponding to $\hat{\Gamma}$. We have
$\overline{\mathcal{M}}_{g,n+h}(\mathbb{P}^{1},\beta)^{T}=\bigcup_{\hat{\Gamma}\in G_{g,n+h}(\mathbb{P}^{1},\beta)}F_{\hat{\Gamma}}$.
By [32, Theorem 73], we get
Proposition 3.4.
Let $\beta=d[\mathbb{P}^{1}]\in E(\mathbb{P}^{1})$. Then for $\gamma_{1},\dots,\gamma_{n+h}\in H^{*}_{T}(\mathbb{P}^{1})$ and $g,a_{1},\dots,a_{n+h}\in\mathbb{Z}_{\geq 0}$, we have
$\langle\tau_{a_{1}}(\gamma_{1}),\dots,\tau_{a_{n+h}}(\gamma_{n+h})\rangle^{\mathbb{P}^{1},T}_{g,n+h,\beta}$
$=$ $\sum_{\hat{\Gamma}\in G_{g,n+h}(\mathbb{P}^{1},\beta)}\frac{1}{|\mathrm{Aut}(\hat{\Gamma})|}\prod_{e\in E(\Gamma)}\frac{{\bf h}(e,d_{e})}{d_{e}}\prod_{v\in V(\Gamma)}\Big({\bf w}(p_{\vec{f}(v)})^{|E_{v}|}\prod_{i\in S_{v}}i^{*}_{p_{\vec{f}(v)}}\gamma_{i}\Big)$
$\cdot\prod_{v\in V(\Gamma)}\int_{\overline{\mathcal{M}}_{g_{v},E_{v}\cup S_{v}}}\frac{{\bf h}(p_{\vec{f}(v)},g_{v})\prod_{i\in S_{v}}\psi_{i}^{a_{i}}}{\prod_{e\in E_{v}}({\bf w}_{(e,v)}-\psi_{(e,v)})}.$
Let $\vec{\Gamma}$ and $\hat{\Gamma}$ be the decorated graphs corresponding to the map $u$ and $\hat{u}$ respectively. There is a map
(4) $G_{g,n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})\rightarrow G_{g,n+h}(\mathbb{P}^{1},\beta):\vec{\Gamma}\mapsto\hat{\Gamma}$.
[…]
By Proposition 3.2, Equation (4) and Proposition 3.4, we give a open/descendant correspondence theorem:
Theorem 3.5.
Let $\beta^{\prime}=d[\mathbb{P}^{1}]+\sum_{i=1}^{h}\mu_{i}[\mathbb{R}\mathbb{P}^{1}]\in H_{2}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$. Then for $\gamma_{1},\dots,\gamma_{n}\in H^{*}_{T}(\mathbb{P}^{1})$ and $g,a_{1},\dots,a_{n}\in\mathbb{Z}_{\geq 0}$, we have
$\langle\tau_{a_{1}}(\gamma_{1})\dots\tau_{a_{n}}(\gamma_{n})\rangle^{\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1},T}_{g,\beta^{\prime},\vec{\mu}}=\prod_{j\in J_{-}}D^{1}(-\mu_{j})\prod_{j\in J_{+}}D^{2}(\mu_{j})\ \ \ \ \ \ \ \ \ \ \ \ \$
$\cdot\int_{[\overline{\mathcal{M}}_{g,n+h}(\mathbb{P}^{1},\beta)]^{{\mathrm{vir}},T}}\frac{\prod_{i=1}^{n}\psi_{i}^{a_{i}}\mathrm{ev}_{i}^{*}(\gamma_{i})\prod_{j\in J_{-}}\mathrm{ev}_{n+j}^{*}\phi_{1}\prod_{j\in J_{+}}\mathrm{ev}_{n+j}^{*}\phi_{2}}{\prod_{j=1}^{h}\frac{\mathsf{v}}{\mu_{j}}(\frac{\mathsf{v}}{\mu_{j}}-\psi_{n+j})}.$
3.5. Generating function of open Gromov-Witten invariants
Consider the generating function of genus $g$, $h$ boundary circles open Gromov-Witten invariants of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$: let $\mathbf{t}=t^{0}1+t^{1}H$,
$F_{g,h}(\mathbf{t},Q;X_{1},\dots,X_{h})=\sum_{\beta\in E(\mathbb{P}^{1})\atop{l\geq 0}}\sum_{\vec{\mu}=(\mu_{1},\dots,\mu_{h})\atop{\mu_{i}\in\mathbb{Z}_{\neq 0}}}\frac{Q^{\beta}}{l!}\langle\mathbf{t}^{l}\rangle^{\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1},T}_{g,\beta^{\prime},\vec{\mu}}\prod_{i=1}^{h}X_{i}^{\mu_{i}}.$
Let
$\tilde{\xi}^{1}_{k}(X)=\sum_{d<0}(\frac{\mathsf{v}}{d})^{-(k+2)}D^{1}(-d)X^{d},$
$\tilde{\xi}^{2}_{k}(X)=\sum_{d>0}(\frac{\mathsf{v}}{d})^{-(k+2)}D^{2}(d)X^{d}.$
For $\alpha\in\{1,2\}$,
$\Big(\frac{1}{\mathsf{v}}X\frac{d}{dX}\Big)\tilde{\xi}^{\alpha}_{k}(X)=\tilde{\xi}^{\alpha}_{k+1}(X),$
$\tilde{\xi}^{\alpha}(z,X):=\sum_{k\in\mathbb{Z}_{\geq-2}}z^{k}\tilde{\xi}^{\alpha}_{k}(X).$
With the above notations and the Theorem 3.5, we have
Proposition 3.6.
  • (1)
    (disk invariants)
    $\ \ \ \ \ F_{0,1}(\mathbf{t},Q;X)$
    $=\sum_{d<0}S_{z}(1,\phi_{1})\Big|_{z=\frac{\mathsf{v}}{d}}D^{1}(-d)X^{d}+\sum_{d>0}S_{z}(1,\phi_{2})\Big|_{z=\frac{\mathsf{v}}{d}}D^{2}(d)X^{d}$
    $=\sum_{\alpha\in\{1,2\}}\Big(\frac{1}{\Delta^{\alpha}}\tilde{\xi}^{\alpha}_{-2}(X)+(\frac{t^{0}}{\Delta^{\alpha}}+\frac{t^{1}}{2})\tilde{\xi}^{\alpha}_{-1}(X)+\sum_{k\geq 0}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\phi_{\alpha}\psi^{k}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{0,1}^{\mathbb{P}^{1},T}\tilde{\xi}^{\alpha}_{k}(X)\Big)$
    $=[z^{-2}]\sum_{\alpha\in\{1,2\}}S_{z}(1,\phi_{\alpha})\tilde{\xi}^{\alpha}(z,X).$
  • (2)
    (annulus invariants)
    $\ \ \ \ \ F_{0,2}(\mathbf{t},Q;X_{1},X_{2})-F_{0,2}(0;X_{1},X_{2})$
    $=\sum_{k_{1},k_{2}\geq 0}\sum_{\alpha_{1},\alpha_{2}\in\{1,2\}}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\phi_{\alpha_{1}}\psi^{k_{1}},\phi_{\alpha_{2}}\psi^{k_{2}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{0,2}^{\mathbb{P}^{1},T}\tilde{\xi}_{k_{1}}^{\alpha_{1}}(X_{1})\tilde{\xi}_{k_{2}}^{\alpha_{2}}(X_{2})$
    $=[z_{1}^{-1}z_{2}^{-1}]\sum_{\alpha_{1},\alpha_{2}\in\{1,2\}}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\frac{\phi_{\alpha_{1}}}{z_{1}-\psi_{1}},\frac{\phi_{\alpha_{2}}}{z_{2}-\psi_{2}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{0,2}^{\mathbb{P}^{1},T}\tilde{\xi}^{\alpha_{1}}(z_{1},X_{1})\tilde{\xi}^{\alpha_{2}}(z_{2},X_{2}),$
    where
    $\frac{1}{\mathsf{v}}\Big(X_{1}\frac{\partial}{\partial X_{1}}+X_{2}\frac{\partial}{\partial X_{2}}\Big)F_{0,2}(0;X_{1},X_{2})=\sum_{\alpha\in\{1,2\}}\frac{1}{\Delta^{\alpha}}\tilde{\xi}^{\alpha}_{0}(X_{1})\tilde{\xi}^{\alpha}_{0}(X_{2}).$
    So we have
    $\ \ \ \ \ \frac{1}{\mathsf{v}}\Big(X_{1}\frac{\partial}{\partial X_{1}}+X_{2}\frac{\partial}{\partial X_{2}}\Big)F_{0,2}(\mathbf{t},Q;X_{1},X_{2})$
    $=[z_{1}^{0}z_{2}^{0}](z_{1}+z_{2})\sum_{\alpha_{1},\alpha_{2}\in\{1,2\}}V_{z_{1},z_{2}}(\phi_{\alpha_{1}},\phi_{\alpha_{2}})\tilde{\xi}^{\alpha_{1}}(z_{1},X_{1})\tilde{\xi}^{\alpha_{2}}(z_{2},X_{2}).$
  • (3)
    For 2g2+n>02g-2+n>0,
    $\ \ \ \ \ F_{g,n}(\mathbf{t},Q;X_{1},\dots,X_{n})$
    $=\sum_{k_{1},\dots,k_{n}\geq 0}\sum_{\alpha_{1},\dots,\alpha_{n}\in\{1,2\}}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\phi_{\alpha_{1}}\psi^{k_{1}},\dots,\phi_{\alpha_{n}}\psi^{k_{n}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{g,n}^{\mathbb{P}^{1},T}\prod_{j=1}^{n}\tilde{\xi}_{k_{j}}^{\alpha_{j}}(X_{j})$
    $=[z_{1}^{-1}\dots z_{n}^{-1}]\sum_{\alpha_{1},\dots,\alpha_{n}\in\{1,2\}}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\frac{\phi_{\alpha_{1}}}{z_{1}-\psi_{1}},\dots,\frac{\phi_{\alpha_{n}}}{z_{n}-\psi_{n}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{g,n}^{\mathbb{P}^{1},T}\prod_{j=1}^{n}\tilde{\xi}^{\alpha_{j}}(z_{j},X_{j}).$
Let $\vec{\Gamma}\in{\bf\Gamma}_{g,n}(\mathbb{P}^{1})$ be a labeled graph defined in Section 2.8. To each ordinary leaf $l_{j}\in L^{o}(\Gamma)$ with $\beta(l_{j})=\beta\in\{1,2\}$ and $k(l_{j})=k\in\mathbb{Z}_{\geq 0}$, we assign the following weight (open leaf)
(5) $(\tilde{\mathcal{L}}^{O})^{\beta}_{k}(l_{j})=[z^{k}](\sum_{\alpha,\gamma\in\{1,2\}}(\tilde{\xi}^{\alpha}(z,X_{j})S^{\hat{\underline{\gamma}}}_{\ \alpha})_{+}R(-z)_{\gamma}^{\ \beta}).$
We define a new weight $\widetilde{\omega}^{O}_{A}(\vec{\Gamma})$ of a labeled graph $\vec{\Gamma}\in{\bf\Gamma}_{g,n}(\mathbb{P}^{1})$ as
$\widetilde{\omega}^{O}_{A}(\vec{\Gamma})=$ $\prod_{v\in V(\Gamma)}\Big(\sqrt{\Delta^{\beta(v)}(q)}\Big)^{2g(v)-2+{\mathrm{val}}(v)}\langle\prod_{h\in H(v)}\tau_{k(h)}\rangle_{g(v)}\prod_{e\in E(\Gamma)}\mathcal{E}^{\beta(v_{1}(e)),\beta(v_{2}(e))}_{k(h_{1}(e)),k(h_{2}(e))}$
$\cdot\prod_{l\in L^{1}(\Gamma)}(\mathcal{L}^{1})^{\beta(l)}_{k(l)}(l)\prod_{j=1}^{n}(\tilde{\mathcal{L}}^{O})^{\beta(l_{j})}_{k(l_{j})}(l_{j}).$
Then by Theorem 2.3 and Proposition 3.6, we have the following graph sum formula for $F_{g,n}$.
Theorem 3.7.
For 2g2+n>02g-2+n>0, we have
$F_{g,n}(\mathbf{t},Q;X_{1},\dots,X_{n})=\sum_{\vec{\Gamma}\in{\bf\Gamma}_{g,n}(\mathbb{P}^{1})}\frac{\widetilde{\omega}_{A}^{O}(\vec{\Gamma})}{|\mathrm{Aut}(\vec{\Gamma})|}.$
Definition 3.8.
We define
$F_{g,n}(\mathbf{t};X_{1},\dots,X_{n}):=F_{g,n}(\mathbf{t},1;X_{1},\dots,X_{n}).$
By Remark 2.1 and Proposition 3.6, the restriction of $F_{g,n}$ to $Q=1$ is well-defined.
Theorem 3.7 implies
Corollary 3.9.
Let $\omega_{A}^{O}(\vec{\Gamma})=\widetilde{\omega}_{A}^{O}(\vec{\Gamma})\big|_{Q=1}$. For 2g2+n>02g-2+n>0, we have
$F_{g,n}(\mathbf{t};X_{1},\dots,X_{n})=\sum_{\vec{\Gamma}\in{\bf\Gamma}_{g,n}(\mathbb{P}^{1})}\frac{\omega_{A}^{O}(\vec{\Gamma})}{|\mathrm{Aut}(\vec{\Gamma})|}.$
Convention 3.10.
From this point onwards, we use the specialization $Q=1$, $q=e^{t^{1}}$ in the calculation.
For $g,n,m\geq 0$, $a_{i}\in\mathbb{Z}_{\geq 0}$ and $\gamma_{i}\in H^{*}_{T}(\mathbb{P}^{1})$, we also define the following double correlators as the generating function of open Gromov-Witten invariants with $m$ descendant insertions and $n$ boundary conpoments:
$\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\tau_{a_{1}}(\gamma_{1})\dots\tau_{a_{m}}(\gamma_{m})\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{g,m,n}^{\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1},T}(\mathbf{t},X_{1},\dots,X_{n})$
$:=\sum_{\beta\in E(\mathbb{P}^{1})\atop{l\geq 0}}\sum_{\vec{\mu}=(\mu_{1},\dots,\mu_{n})\atop{\mu_{i}\in\mathbb{Z}_{\neq 0}}}\frac{1}{l!}\langle\tau_{a_{1}}(\gamma_{1})\dots\tau_{a_{m}}(\gamma_{m}),\mathbf{t}^{l}\rangle_{g,\beta^{\prime},\vec{\mu}}^{\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1},T}\prod_{i=1}^{n}X_{i}^{\mu_{i}}.$
We also have
$\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\tau_{a_{1}}(\gamma_{1})\dots\tau_{a_{m}}(\gamma_{m})\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{g,m,n}^{\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1},T}(\mathbf{t},X_{1},\dots,X_{n})$
$=\sum_{k_{1},\dots,k_{n}\geq 0\atop{\alpha_{1},\dots,\alpha_{n}\in\{1,2\}}}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\tau_{a_{1}}(\gamma_{1}),\dots,\tau_{a_{m}}(\gamma_{m}),\phi_{\alpha_{1}}\psi^{k_{1}},\dots,\phi_{\alpha_{n}}\psi^{k_{n}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{g,n}^{\mathbb{P}^{1},T}\prod_{j=1}^{n}\tilde{\xi}_{k_{j}}^{\alpha_{j}}(X_{j}).$
For a labeled graph $\vec{\Gamma}\in{\bf\Gamma}_{g,m+n}(\mathbb{P}^{1})$ with $L^{o}(\Gamma)=\{l_{1},\dots,l_{m+n}\}$, we introduce a new weight $\mathsf{w}_{A}(\vec{\Gamma})$ for the labeled graph $\vec{\Gamma}\in{\bf\Gamma}_{g,m+n}(\mathbb{P}^{1})$ as
$\mathsf{w}_{A}(\vec{\Gamma})=$ $\prod_{v\in V(\Gamma)}\Big(\sqrt{\Delta^{\beta(v)}(q)}\Big)^{2g(v)-2+{\mathrm{val}}(v)}\langle\prod_{h\in H(v)}\tau_{k(h)}\rangle_{g(v)}\prod_{e\in E(\Gamma)}\mathcal{E}^{\beta(v_{1}(e)),\beta(v_{2}(e))}_{k(h_{1}(e)),k(h_{2}(e))}$
$\cdot\prod_{l\in L^{1}(\Gamma)}(\mathcal{L}^{1})^{\beta(l)}_{k(l)}(l)\prod_{j=1}^{m}(\mathcal{L}^{\bf u})_{k(l_{j})}^{\beta(l_{j})}(l_{j})\prod_{j=1}^{n}(\tilde{\mathcal{L}}^{O})^{\beta(l_{m+j})}_{k(l_{m+j})}(l_{m+j})\Big|_{Q=1}.$
Theorem 3.11.
Suppose $g,m\geq 0$, $n\geq 1$ and 2g2+m+n>02g-2+m+n>0. It holds that
$\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}{\bf u}_{1},\dots,{\bf u}_{m}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{g,m,n}^{\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1},T}(\mathbf{t},X_{1},\dots,X_{n})=\sum_{\vec{\Gamma}\in{\bf\Gamma}_{g,m+n}(\mathbb{P}^{1})}\frac{\mathsf{w}_{A}(\vec{\Gamma})}{|\mathrm{Aut}(\vec{\Gamma})|}.$
3.6. Explicit formula of $F_{0,1}$
Following from Proposition 3.6, the generating function $F_{0,1}$ is given by
$F_{0,1}(\mathbf{t};X)=\sum_{d>0}\Big(J_{1}(-\mathsf{v}/d)D^{1}(d)X^{-d}+J_{2}(\mathsf{v}/d)D^{2}(d)X^{d}\Big),$
where $J_{\alpha}(z):=(J(z),\phi_{\alpha})_{\mathbb{P}^{1},T}$ are the components of the $J$-function in Section 2.6.
By Equation (1), for $d>0$
$J^{1}(-\mathsf{v}/d)$ $=-\mathsf{v}J_{1}(-\mathsf{v}/d)=e^{-dt^{0}/\mathsf{v}}(-\mathsf{v}/d)^{d}\Gamma(d+1)I_{d}(-2\sqrt{q}d/\mathsf{v}),$
$J^{2}(\mathsf{v}/d)$ $=\mathsf{v}J_{2}(\mathsf{v}/d)=e^{dt^{0}/\mathsf{v}}(\mathsf{v}/d)^{d}\Gamma(d+1)I_{d}(2\sqrt{q}d/\mathsf{v}).$
We get
$F_{0,1}(\mathbf{t};X)=\sum_{d>0}e^{-dt^{0}/\mathsf{v}}\frac{\mathsf{v}}{d^{2}}I_{d}(-2\sqrt{q}d/\mathsf{v})X^{-d}+\sum_{d>0}e^{dt^{0}/\mathsf{v}}\frac{\mathsf{v}}{d^{2}}I_{d}(2\sqrt{q}d/\mathsf{v})X^{d}.$
Let $q,\mathsf{v}$ be positive real numbers. By the symmetry of the modified Bessel function $I_{\alpha}(x)$ (see Appendix A), we have
(6) $F_{0,1}(\mathbf{t};X)=\sum_{d\in\mathbb{Z}_{\neq 0}}e^{dt^{0}/\mathsf{v}}\frac{\mathsf{v}}{d^{2}}I_{d}(2\sqrt{q}d/\mathsf{v})X^{d}.$
4. Mirror curve and topological recursion
4.1. Mirror curve
Let $Y$ be a coordinate on $\mathbb{C}^{*}$ and let $q=e^{t^{1}},\mathsf{v}$ be positive real numbers. The $T$-equivariant superpotential $W:\mathbb{C}^{*}\rightarrow\mathbb{C}$ is defined as
$W(Y)=t^{0}+\frac{\mathsf{v}}{2}\log q+Y+\frac{q}{Y}-\mathsf{v}\log Y.$
Consider the spectral curve $C_{q}$ defined as follows:
$C_{q}=\{(x,y)\in\mathbb{C}^{2}:x=W(e^{y})\}.$
Let $X=e^{-x/\mathsf{v}}$, $Y=e^{y}$. The curve $C_{q}$ has a parameterization:
$\left\{\begin{aligned} x(Y)&=t^{0}+\frac{\mathsf{v}}{2}\log q+Y+\frac{q}{Y}-\mathsf{v}\log Y,\\ y(Y)&=\log Y.\end{aligned}\right$.
Let $\mathbb{P}^{1}$ be the compactification of $\mathbb{C}^{*}$ with $Y\in\mathbb{C}^{*}\subset\mathbb{P}^{1}$ as its coordinate. Consider the differential $dx(Y)=(Y-q/Y-\mathsf{v})dy$. The branch points $P_{\alpha}$ of $dx$ are given by
$P_{1}=\frac{\mathsf{v}-\mathsf{v}\sqrt{1+4q/\mathsf{v}^{2}}}{2},\quad P_{2}=\frac{\mathsf{v}+\mathsf{v}\sqrt{1+4q/\mathsf{v}^{2}}}{2}.$
Near each branch point $Y=P_{\alpha}$, we choose a local coordinates $\zeta_{\alpha}$ such that
$x=\check{u}^{\alpha}-\zeta_{\alpha}^{2},\quad y=\check{v}^{\alpha}-\sum_{k=1}^{\infty}h^{\alpha}_{k}(q)\zeta^{k}_{\alpha}.$
Let $\lambda=xdy$ be the Liouville form on $\mathbb{C}^{2}$, $\Phi:=\lambda|_{C_{q}}$. Let $B(Y_{1},Y_{2})$ be a holomorphic bidifferential on $\mathbb{P}^{1}$ defined by
$B(Y_{1},Y_{2})=\frac{dY_{1}dY_{2}}{(Y_{1}-Y_{2})^{2}}.$
Around $(P_{\alpha},P_{\beta})$, we expand $B(Y_{1},Y_{2})$ as
$B(\zeta_{\alpha},\zeta_{\beta})=\Big(\frac{\delta_{\alpha\beta}}{(\zeta_{\alpha}-\zeta_{\beta})^{2}}+\sum_{k,l\geq 0}B^{\alpha,\beta}_{k,l}\zeta^{k}_{\alpha}\zeta^{l}_{\beta}\Big)d\zeta_{\alpha}d\zeta_{\beta}.$
Given $\alpha=1,2$ and $d\in\mathbb{Z}_{\geq 0}$, define
$d\xi_{\alpha,d}(p):=(2d-1)!!2^{-d}\mathrm{Res}_{p^{\prime}\rightarrow P_{\alpha}}B(p,p^{\prime})(\sqrt{-1}\zeta_{\alpha})^{-2d-1}.$
Then $d\xi_{\alpha,d}$ is a meromorphic 1-form on $\mathbb{P}^{1}$ with a single pole of order $2d+2$ at $P_{\alpha}$. In the local coordinate $\zeta_{\alpha}$ near $P_{\alpha}$, $d\xi_{\alpha,d}$ has the following expansion:
$d\xi_{\alpha,d}=\left(\frac{-(2d+1)!!}{2^{d}\sqrt{-1}^{2d+1}\zeta_{\alpha}^{2d+2}}+\text{analytic part in $\zeta_{\alpha}$}\right)d\zeta_{\alpha}.$
In particular, we have
$d\xi_{\alpha,0}=\sqrt{\frac{-2}{\Delta^{\alpha}(q)}}d\left(\frac{P_{\alpha}}{Y-P_{\alpha}}\right).$
4.2. The $\mathfrak{h}_{X}$-operator
The function $X=e^{-x/\mathsf{v}}$ has an essential singularity at $Y=0$:
$X=e^{-x/\mathsf{v}}=\frac{e^{-t^{0}/\mathsf{v}}}{\sqrt{q}}Ye^{-(Y+q/Y)/\mathsf{v}}.$
We also notice that $dX/X$ is a meromorphic differential on $\mathbb{P}^{1}$:
$\frac{dX}{X}=-\frac{dx}{\mathsf{v}}=\frac{Y^{2}-\mathsf{v}Y-q}{-\mathsf{v}Y^{2}}dY=\frac{(Y-P_{1})(Y-P_{2})}{-\mathsf{v}Y^{2}}dY.$
Let $D_{\epsilon}$ be a punctured disk around $Y=0$:
$D_{\epsilon}:=\{Y\in\mathbb{C}^{*}:0<|Y|<\epsilon\},$
where $\epsilon$ is a small positive real number such that $D_{\epsilon}$ does not contain any ramification points $\{P_{1},P_{2}\}$. In the punctured disk $D_{\epsilon}$, $Y$ can be expanded as a power series of $X$ by the Lagrange inversion theorem. At $Y=0$, we define the $\mathfrak{h}_{X}$-operator, which expands holomorphic functions $f(Y)$ and holomorphic differential forms $\theta(Y)$ on $D_{\epsilon}$ to Laurent series in $X$, by their local behaviors at $Y=0$:
  • For a holomorphic function $f(Y)$ on $D_{\epsilon}$, let
    $\mathfrak{h}_{X}(f):=\sum_{\mu\in\mathbb{Z}_{\neq 0}}\Big(\mathop{\mathrm{Res}}\limits_{Y\rightarrow 0}f(Y)X^{-\mu}\frac{dX}{X}\Big)X^{\mu}.$
  • For a holomorphic differential form $\theta(Y)$ on $D_{\epsilon}$, let
    $\mathfrak{h}_{X}(\theta):=\sum_{\mu\in\mathbb{Z}_{\neq 0}}\Big(\mathop{\mathrm{Res}}\limits_{Y\rightarrow 0}\theta(Y)X^{-\mu}\Big)\frac{X^{\mu}}{\mu}.$
    If $f$ is a holomorphic function on $D_{\epsilon}$ such that $\theta(Y)=df$, then $\mathfrak{h}_{X}(\theta)=\mathfrak{h}_{X}(f)$ by the integrations by parts.
For multi-holomorphic functions (resp. forms) on $(D_{\epsilon})^{\times n}$, we define $\mathfrak{h}_{X_{1},\dots,X_{n}}$ as follows:
  • Let $f(Y_{1},\dots,Y_{n})$ be a holomorphic function on $(D_{\epsilon})^{\times n}$, we define
    $\mathfrak{h}_{X_{1},\dots,X_{n}}(f):=\sum_{\mu_{1},\dots,\mu_{n}\in\mathbb{Z}_{\neq 0}}\Big(\mathop{\mathrm{Res}}\limits_{Y_{1}\rightarrow 0}\cdots\mathop{\mathrm{Res}}\limits_{Y_{n}\rightarrow 0}f(Y_{1},\dots,Y_{n})\prod_{i=1}^{n}X^{-\mu_{i}}_{i}\frac{dX_{i}}{X_{i}}\Big)\prod_{i=1}^{n}X^{\mu_{i}}_{i}.$
  • Let $\theta(Y_{1},\dots,Y_{n})$ be a holomorphic differential form from the set of sections $\Gamma((D_{\epsilon})^{\times n},\omega_{C_{q}}^{\boxtimes n})$, we define
    $\mathfrak{h}_{X_{1},\dots,X_{n}}(\theta):=\sum_{\mu_{1},\dots,\mu_{n}\in\mathbb{Z}_{\neq 0}}\Big(\mathop{\mathrm{Res}}\limits_{Y_{1}\rightarrow 0}\cdots\mathop{\mathrm{Res}}\limits_{Y_{n}\rightarrow 0}\theta(Y_{1},\dots,Y_{n})\prod_{i=1}^{n}X^{-\mu_{i}}_{i}\Big)\prod_{i=1}^{n}\frac{X^{\mu_{i}}_{i}}{\mu_{i}}.$
The $\mathfrak{h}_{X}$-operator plays an essential role in the definition of open leaves in Section 4.3. Let
$\xi_{\alpha,0}(Y)=\sqrt{\frac{-2}{\Delta^{\alpha}(q)}}\frac{P_{\alpha}}{Y-P_{\alpha}},$
which is a meromorphic function on $\mathbb{P}^{1}$. It has a simple pole at $P_{\alpha}$ and is holomorphic elsewhere. Let
$\eta:=\xi_{2,0}-\sqrt{-1}\xi_{1,0},\quad\chi:=\xi_{2,0}+\sqrt{-1}\xi_{1,0}.$
Let $q,\mathsf{v}>0$, $\Delta^{2}(q)=\mathsf{v}\sqrt{1+4q/\mathsf{v}^{2}}>0$. Then we have
$\eta$ $=\sqrt{\frac{-2}{\Delta^{2}(q)}}\left(\frac{P_{2}}{Y-P_{2}}-\frac{P_{1}}{Y-P_{1}}\right)$
$=\sqrt{\frac{-2}{\Delta^{2}(q)}}\frac{Y\Delta^{2}(q)}{(Y-P_{1})(Y-P_{2})}$
$=\frac{1}{-\mathsf{v}}\sqrt{-2\Delta^{2}(q)}\frac{dY}{dX/X}\frac{1}{Y},$
and
$\chi$ $=\sqrt{\frac{-2}{\Delta^{2}(q)}}\left(\frac{P_{2}}{Y-P_{2}}+\frac{P_{1}}{Y-P_{1}}\right)$
$=\sqrt{\frac{-2}{\Delta^{2}(q)}}\frac{\mathsf{v}Y+2q}{(Y-P_{1})(Y-P_{2})}$
$=-\sqrt{\frac{-2}{\Delta^{2}(q)}}\left(\frac{dY}{dX/X}\frac{1}{Y}+\frac{2qdY}{\mathsf{v}dX/X}\frac{1}{Y^{2}}\right).$
Apply the operator $\mathfrak{h}_{X}$ to the meromorphic functions $\eta$ and $\chi$, we obtain two Laurent series:
$\mathfrak{h}_{X}(\eta):=\sum_{\mu\in\mathbb{Z}_{\neq 0}}A_{\mu}X^{\mu},\quad\mathfrak{h}_{X}(\chi):=\sum_{\mu\in\mathbb{Z}_{\neq 0}}B_{\mu}X^{\mu}.$
By calculation,
(7) $A_{\mu}$ $=\mathrm{Res}_{Y=0}\eta X^{-\mu}\frac{dX}{X}$
$=\frac{1}{-\mathsf{v}}\sqrt{-2\Delta^{2}(q)}e^{\mu t^{0}/\mathsf{v}}\sqrt{q}^{\mu}\mathrm{Res}_{Y=0}e^{\mu(Y+q/Y)/\mathsf{v}}\frac{dY}{Y^{\mu+1}}$
$=\frac{1}{-\mathsf{v}}\sqrt{-2\Delta^{2}(q)}e^{\mu t^{0}/\mathsf{v}}I_{\mu}\Big(\frac{2\sqrt{q}\mu}{\mathsf{v}}\Big),$
and
(8) $B_{\mu}$ $=\mathrm{Res}_{Y=0}\chi X^{-\mu}\frac{dX}{X}$
$=-\sqrt{\frac{-2}{\Delta^{2}(q)}}e^{\mu t^{0}/\mathsf{v}}\sqrt{q}^{\mu}\mathrm{Res}_{Y=0}e^{\mu(Y+q/Y)/\mathsf{v}}\left(\frac{dY}{Y^{\mu+1}}+\frac{2q}{\mathsf{v}}\frac{dY}{Y^{\mu+2}}\right)$
$=-\sqrt{\frac{-2}{\Delta^{2}(q)}}e^{\mu t^{0}/\mathsf{v}}\left(I_{\mu}\Big(\frac{2\sqrt{q}\mu}{\mathsf{v}}\Big)+\frac{2\sqrt{q}}{\mathsf{v}}I_{\mu+1}\Big(\frac{2\sqrt{q}\mu}{\mathsf{v}}\Big)\right)$
$=-\sqrt{\frac{-2}{\Delta^{2}(q)}}e^{\mu t^{0}/\mathsf{v}}\frac{2\sqrt{q}}{\mathsf{v}}I^{\prime}_{\mu}\Big(\frac{2\sqrt{q}\mu}{\mathsf{v}}\Big).$
Remark 4.1.
More details about the residue calculation can be found in Appendix B.
4.3. Chekhov-Eynard-Orantin topological recursion
Let $\omega_{g,n}$ be the differential forms defined recursively by the Chekhov-Eynard-Orantin topological recursion [15]:
$\omega_{0,1}=0,\quad\omega_{0,2}=B(Y_{1},Y_{2}).$
For 2g2+n>02g-2+n>0,
$\omega_{g,n}(Y_{1},\dots,Y_{n})=\sum_{\alpha\in\{1,2\}}$ $\mathrm{Res}_{Y\rightarrow P_{\alpha}}\frac{-\int_{\xi=Y}^{\hat{Y}}B(Y_{n},\xi)}{2(\log(Y)-\log(\hat{Y}))dx}\Big(\omega_{g-1,n+1}(Y,\hat{Y},Y_{1},\dots,Y_{n-1})$
$+\sum_{g_{1}+g_{2}=g}\sum_{I\sqcup J=\{1,\dots,n-1\}}\omega_{g_{1},|I|+1}(Y,Y_{I})\omega_{g_{2},|J|+1}(\hat{Y},Y_{J})\Big),$
where $Y\neq P_{\alpha}$ is in a neighbourhood of $P_{\alpha}$, and $\hat{Y}\neq Y$ is the conjugate point of $Y$ such that $x(\hat{Y})=x(Y)$.
We introduce the following notations:
  1. (1)
    Given $\alpha=1,2$ and $k\in\mathbb{Z}_{\geq 0}$, define
    $W^{\alpha}_{k}:=d((-\frac{d}{dx})^{k}(\xi_{\alpha,0})),$
    $\theta_{\alpha}(z):=\sum_{k=0}^{\infty}d\xi_{\alpha,k}z^{k},\quad\hat{\theta}_{\alpha}(z):=\sum_{k=0}^{\infty}W^{\alpha}_{k}z^{k}.$
  2. (2)
    The B-model $R$-matrix $\check{R}_{\beta}^{\ \alpha}(z)$ (which is a power series of $z$) is defined by asymptotic expansion
    $\check{R}_{\beta}^{\ \alpha}(z)\sim\frac{\sqrt{z}e^{-\check{u}^{\alpha}/z}}{2\sqrt{\pi}}\int_{\gamma_{\alpha}}e^{x/z}d\xi_{\beta,0}.$
    Here $\gamma_{\alpha}$ is the Lefschetz thimble of the map $x$, i.e. $x(\gamma_{\alpha})-\check{u}^{\alpha}\in\mathbb{R}_{\geq 0}$.
  3. (3)
    Given $\alpha,\beta\in\{1,2\}$, $k,l\geq 0$, define
    $\check{B}^{\alpha,\beta}_{k,l}$ $=\frac{(2k-1)!!(2l-1)!!}{2^{k+l+1}}B^{\alpha,\beta}_{2k,2l}$
    $\ =[z^{k}w^{l}]\Big(\frac{1}{z+w}\Big(\delta_{\alpha\beta}-\sum_{\gamma\in\{1,2\}}\check{R}_{\gamma}^{\ \alpha}(-z)\check{R}_{\gamma}^{\ \beta}(-w)\Big)\Big).$
  4. (4)
    Given $\alpha=1,2$ and $k\in\mathbb{Z}_{\geq 1}$, define
    $\check{h}^{\alpha}_{k}$ $=-\frac{\sqrt{-1}^{2k-1}(2k-1)!!}{2^{k-1}}h^{\alpha}_{2k-1}$
    $\ =[z^{k-1}]\Big(\sum_{\beta}\sqrt{-1}h^{\beta}_{1}\check{R}_{\beta}^{\ \alpha}(-z)\Big).$
In similar to Section 2.8, we introduce the B-model graph sum formula for $\omega_{g,n}$. Let $\mathbf{p}=(Y_{1},\dots,Y_{n})\in(\mathbb{P}^{1})^{n}$. For a labeled graph $\vec{\Gamma}\in{\bf\Gamma}_{g,n}(\mathbb{P}^{1})$ with $L^{o}(\Gamma)=\{l_{1},\dots,l_{n}\}$, and $\bullet=\mathbf{p}$ or $O$, we assign the weight of $\vec{\Gamma}$ as
$\omega_{B}^{\bullet}(\vec{\Gamma})$ $=(-1)^{g(\vec{\Gamma})-1+n}\prod_{v\in V(\Gamma)}\Big(\frac{h^{\beta(v)}_{1}}{\sqrt{2}}\Big)^{2-2g(v)-{\mathrm{val}}(v)}\langle\prod_{h\in H(v)}\tau_{k(h)}\rangle_{g(v)}$
$\ \ \ \ \times\prod_{e\in E(\Gamma)}\check{B}^{\beta(v_{1}(e)),\beta(v_{2}(e))}_{k(h_{1}(e)),k(h_{2}(e))}\prod_{l\in L^{1}(\Gamma)}\frac{-1}{\sqrt{-2}}\check{h}^{\beta(l)}_{k(l)}\prod_{j=1}^{n}(\check{\mathcal{L}}^{\bullet})^{\beta(l_{j})}_{k(l_{j})}(l_{j}),$
where
  • (descendant leaf)
    $(\check{\mathcal{L}}^{\mathbf{p}})^{\beta(l_{j})}_{k(l_{j})}(l_{j})=\frac{1}{\sqrt{-2}}d\xi_{\beta(l_{j}),k(l_{j})}(Y_{j}),$
  • (open leaf)
    $(\check{\mathcal{L}}^{O})^{\beta(l_{j})}_{k(l_{j})}(l_{j})=\frac{1}{\sqrt{-2}}\mathfrak{h}_{X_{j}}(d\xi_{\beta(l_{j}),k(l_{j})}(Y_{j})).$
We have the B-model graph sum formula from [14, Theorem 3.7]:
Theorem 4.2 (Dunin-Barkowski-Orantin-Shadrin-Spitz [14]).
For 2g2+n>02g-2+n>0,
$\omega_{g,n}(\mathbf{p})$ $=\sum_{\vec{\Gamma}\in{\bf\Gamma}_{g,n}(\mathbb{P}^{1})}\frac{\omega_{B}^{\mathbf{p}}(\vec{\Gamma})}{|\mathrm{Aut}(\vec{\Gamma})|},$
$\mathfrak{h}_{X_{1},\dots,X_{n}}(\omega_{g,n})$ $=\sum_{\vec{\Gamma}\in{\bf\Gamma}_{g,n}(\mathbb{P}^{1})}\frac{\omega_{B}^{O}(\vec{\Gamma})}{|\mathrm{Aut}(\vec{\Gamma})|}.$
We define the B-model open primary potentials as follows,
  • Define the B-model disk potential $W_{0,1}(t^{0},q;X)$ as a Laurent series in $X$ with constant term zero such that
    $\Big(-\frac{1}{\mathsf{v}}X\frac{d}{dX}\Big)W_{0,1}(t^{0},q;X)=\mathfrak{h}_{X}\Big(\frac{\partial\Phi}{\partial t^{0}}\Big).$
  • Let $\tilde{\omega}_{0,2}(Y_{1},Y_{2})$ be the meromorphic 2-form on $\mathbb{C}^{*}\times\mathbb{C}^{*}$,
    $\tilde{\omega}_{0,2}(Y_{1},Y_{2}):=\omega_{0,2}(Y_{1},Y_{2})-\frac{dX_{1}dX_{2}}{(X_{1}-X_{2})^{2}}.$
    $\tilde{\omega}_{0,2}$ is holomorphic on $(D_{\epsilon})^{\times 2}$. Define the B-model annulus invariants by
    $W_{0,2}(t^{0},q;X_{1},X_{2})=\mathfrak{h}_{X_{1},X_{2}}(\tilde{\omega}_{0,2}(Y_{1},Y_{2})).$
  • For 2g2+n>02g-2+n>0, $\omega_{g,n}$ is holomorphic at $(D_{\epsilon})^{\times n}$, we define
    $W_{g,n}(t^{0},q;X_{1},\dots,X_{n})=\mathfrak{h}_{X_{1},\dots,X_{n}}(\omega_{g,n}(\mathbf{p})).$
4.4. SYZ T-dual
Definition 4.3 (equivariant K-theoretic framing).
We define
$\widetilde{\mathrm{ch}}_{z}:K_{T}(\mathbb{P}^{1})\rightarrow H^{*}_{T}(\mathbb{P}^{1};\mathbb{Q})\left[\!\left[\frac{\mathsf{v}}{z}\right]\!\right]$
by the following two properties:
  • (a)
    $\widetilde{\mathrm{ch}}_{z}$ is a homomorphism of additive groups:
    $\widetilde{\mathrm{ch}}_{z}(\mathcal{E}_{1}\oplus\mathcal{E}_{2})=\widetilde{\mathrm{ch}}_{z}(\mathcal{E}_{1})+\widetilde{\mathrm{ch}}_{z}(\mathcal{E}_{2}).$
  • (b)
    If $\mathcal{E}$ is a $T$-equivariant line bundle on $\mathbb{P}^{1}$, then
    $\widetilde{\mathrm{ch}}_{z}(\mathcal{E})=\exp\left(-\frac{2\pi\sqrt{-1}(c_{1})_{T}(\mathcal{E})}{z}\right).$
For any $\mathcal{E}\in K_{T}(\mathbb{P}^{1})$, we define the $K$-theoretic framing of $\mathcal{E}$ by
$\kappa_{z}(\mathcal{E}):=(-z)^{1-(c_{1})_{T}(T\mathbb{P}^{1})/z}\Gamma\left(1-\frac{(c_{1})_{T}(T\mathbb{P}^{1})}{z}\right)\widetilde{\mathrm{ch}}_{z}(\mathcal{E}),$
where $(c_{1})_{T}(T\mathbb{P}^{1})=2H$.
Definition 4.4 (equivariant SYZ T-dual).
Let $\mathcal{E}=\mathcal{O}_{\mathbb{P}^{1}}(l_{1}p_{1}+l_{2}p_{2})$ be an equivariant ample line bundle on $\mathbb{P}^{1}$, where $l_{1},l_{2}$ are integers such that $l_{1}+l_{2}>0$. We define the equivariant SYZ T-dual $\mathrm{SYZ}(\mathcal{E})$ of $\mathcal{E}$ to be the oriented path in $\mathbb{C}$ (see Figure 2). We extend the definition additively to the equivariant K-theory group $K_{T}(\mathbb{P}^{1})$.
$\infty+(-2l_{1}-1)\pi i$$(-2l_{1}-1)\pi i$$(2l_{2}-1)\pi i$$\infty+(2l_{2}-1)\pi i$
Figure 2. SYZ $T$-dual of $\mathcal{E}$ in $\mathbb{C}$
Suppose $g,m,n\geq 0$ and 2g2+m+n>02g-2+m+n>0, for any $\mathcal{E}_{1},\dots,\mathcal{E}_{m}\in K_{T}(\mathbb{P}^{1})$, we define the B-model open descendant potential as
(9) $W_{g,m,n}(t^{0},q;\mathcal{E}_{1},\dots,\mathcal{E}_{m};X_{1},\dots,X_{n})$
$=\int_{y_{1}\in\mathrm{SYZ}(\mathcal{E}_{1})}\cdots\int_{y_{m}\in\mathrm{SYZ}(\mathcal{E}_{m})}e^{\sum_{i=1}^{m}x(y_{i})/z_{i}}\mathfrak{h}_{X_{1},\dots,X_{n}}(\omega_{g,m+n}),$
where $\int_{y_{i}\in\mathrm{SYZ}(\mathcal{E}_{i})}$ acts on the $i$-th variables of $\omega_{g,m+n}$ and $\mathfrak{h}_{X_{1},\dots,X_{n}}$ acts on the last $n$ variables.
For a labeled graph $\vec{\Gamma}\in{\bf\Gamma}_{g,m+n}(\mathbb{P}^{1})$ with $L^{o}(\Gamma)=\{l_{1},\dots,l_{m+n}\}$. For the ordinary leaf $l_{j}\in L^{o}(\Gamma)$ with $1\leq j\leq m$, we assign the SYZ weight (SYZ leaf factor)
$(\check{\mathcal{L}}^{\mathrm{SYZ}})^{\beta(l_{j})}_{k(l_{j})}(l_{j})=\frac{1}{\sqrt{-2}}\int_{y_{j}\in\mathrm{SYZ}(\mathcal{E}_{j})}e^{x(y_{j})/z_{j}}d\xi_{\beta(l_{j}),k(l_{j})}(Y_{j}),$
and for the ordinary leaf $l_{m+j}\in L^{o}(\Gamma)$ with $1\leq j\leq n$, we assign the open leaf factor
$(\check{\mathcal{L}}^{O})^{\beta(l_{m+j})}_{k(l_{m+j})}(l_{m+j})=\frac{1}{\sqrt{-2}}\mathfrak{h}_{X_{j}}(d\xi_{\beta(l_{m+j}),k(l_{m+j})}(Y_{m+j})).$
We have the following B-model graph sum formula of $W_{g,m,n}$ with 2g2+m+n>02g-2+m+n>0:
(10) $W_{g,m,n}(t^{0},q;\mathcal{E}_{1},\dots,\mathcal{E}_{m};X_{1},\dots,X_{n})=\sum_{\vec{\Gamma}\in{\bf\Gamma}_{g,m+n}(\mathbb{P}^{1})}\frac{{\mathsf{w}}_{B}(\vec{\Gamma})}{|\mathrm{Aut}(\vec{\Gamma})|},$
where
${\mathsf{w}}_{B}(\vec{\Gamma})$ $=(-1)^{g(\vec{\Gamma})-1+m+n}\prod_{v\in V(\Gamma)}\Big(\frac{h^{\beta(v)}_{1}}{\sqrt{2}}\Big)^{2-2g(v)-{\mathrm{val}}(v)}\langle\prod_{h\in H(v)}\tau_{k(h)}\rangle_{g(v)}$
$\ \ \ \ \times\prod_{e\in E(\Gamma)}\check{B}^{\beta(v_{1}(e)),\beta(v_{2}(e))}_{k(h_{1}(e)),k(h_{2}(e))}\prod_{l\in L^{1}(\Gamma)}\frac{-1}{\sqrt{-2}}\check{h}^{\beta(l)}_{k(l)}$
$\ \ \ \ \times\ \ \prod_{j=1}^{m}(\check{\mathcal{L}}^{\mathrm{SYZ}})^{\beta(l_{j})}_{k(l_{j})}(l_{j})\prod_{j=1}^{n}(\check{\mathcal{L}}^{O})^{\beta(l_{m+j})}_{k(l_{m+j})}(l_{m+j}).$
5. Mirror Symmetry
5.1. Mirror symmetry of disk invariants
Let $\Phi=xdy|_{C_{q}}$ which is a 1-form on $C_{q}$. Let
$\Phi_{0}:=\frac{\partial\Phi}{\partial t^{0}}=\frac{dY}{Y}.$
Consider
$\mathfrak{h}_{X}(\Phi_{0})=\sum_{\mu\in\mathbb{Z}_{\neq 0}}R_{\mu}X^{\mu},$
where
$R_{\mu}$ $=\frac{1}{\mu}\mathrm{Res}_{Y=0}X^{-\mu}\frac{dY}{Y}$
$=\frac{1}{\mu}e^{\mu t^{0}/\mathsf{v}}\sqrt{q}^{\mu}\mathrm{Res}_{Y=0}e^{\mu(Y+q/Y)/\mathsf{v}}\frac{dY}{Y^{\mu+1}}$
$=\frac{1}{\mu}e^{\mu t^{0}/\mathsf{v}}I_{\mu}\Big(\frac{2\sqrt{q}\mu}{\mathsf{v}}\Big).$
Compare with Equation (6), we have
$\frac{\partial}{\partial t^{0}}F_{0,1}(\mathbf{t};X)=\Big(\frac{1}{\mathsf{v}}X\frac{d}{dX}\Big)F_{0,1}=\mathfrak{h}_{X}(\Phi_{0}).$
Therefore, we obtain the following theorem:
Theorem 5.1 (Mirror symmetry of disk invariants).
$F_{0,1}(\mathbf{t};X)=-W_{0,1}(t^{0},q;X).$
5.2. Identification of open leaves
We define
$U(z)(\mathbf{t},X):=\sum_{\alpha\in\{1,2\}}\tilde{\xi}^{\alpha}(z,X)S_{z}(1,\phi_{\alpha})\Big|_{Q=1}.$
For $m\in\mathbb{Z}_{\geq-2}$,
$\left(\frac{1}{\mathsf{v}}X\frac{d}{dX}\right)[z^{m}](U(z)(\mathbf{t},X))=[z^{m+1}](U(z)(\mathbf{t},X)).$
In Section 3.6, we have shown that
$F_{0,1}=[z^{-2}]U(z)(\mathbf{t},X)=\sum_{d\in\mathbb{Z}_{\neq 0}}e^{dt^{0}/\mathsf{v}}\frac{\mathsf{v}}{d^{2}}I_{d}(2\sqrt{q}d/\mathsf{v})X^{d},$
$z^{0}]U(z)(\mathbf{t},X)=\left(\frac{1}{\mathsf{v}}X\frac{d}{dX}\right)^{2}F_{0,1}=\frac{1}{\mathsf{v}}\sum_{d\in\mathbb{Z}_{\neq 0}}e^{dt^{0}/\mathsf{v}}I_{d}(2\sqrt{q}d/\mathsf{v})X^{d}.$
By quantum differential equation,
$z\frac{\partial U}{\partial t^{1}}=\sum_{\alpha\in\{1,2\}}\tilde{\xi}^{\alpha}(z,X)S_{z}(H,\phi_{\alpha})\Big|_{Q=1},$
and
$\frac{\partial U}{\partial t^{1}}$ $=\frac{\partial}{\partial t^{1}}\left(\frac{1}{\mathsf{v}}X\frac{d}{dX}\right)[z^{-2}](U(z)(\mathbf{t},X))$
$=\frac{\partial}{\partial t^{1}}\sum_{d\in\mathbb{Z}_{\neq 0}}e^{dt^{0}/\mathsf{v}}\frac{1}{d}I_{d}(2\sqrt{q}d/\mathsf{v})X^{d}$
$=\sum_{d\in\mathbb{Z}_{\neq 0}}e^{dt^{0}/\mathsf{v}}\frac{\sqrt{q}}{\mathsf{v}}I^{\prime}_{d}(2\sqrt{q}d/\mathsf{v})X^{d}.$
Comparing with Equation (7) and Equation (8), we have
$\mathfrak{h}_{X}(\eta)$ $=-\sqrt{-2\Delta^{2}(q)}[z^{0}]U(z)(\mathbf{t},X),$
$\mathfrak{h}_{X}(\chi)$ $=-2\sqrt{\frac{-2}{\Delta^{2}(q)}}[z^{0}]\tilde{\xi}^{\alpha}(z,X)S_{z}(H,\phi_{\alpha})\Big|_{Q=1}.$
Recall that
$\hat{\phi}_{\alpha}(q)=(\Psi^{-1})_{\alpha}^{\ 0}1+(\Psi^{-1})_{\alpha}^{\ 1}H,$
$\sum_{\beta\in\{1,2\}}\tilde{\xi}^{\beta}(z,X)S(\hat{\phi}_{\alpha}(q),\phi_{\beta})\Big|_{Q=1}=(\Psi^{-1})_{\alpha}^{\ 0}U+z(\Psi^{-1})_{\alpha}^{\ 1}\frac{\partial U}{\partial t^{1}}.$
We obtain
$\sum_{\beta\in\{1,2\}}\tilde{\xi}^{\beta}(z,X)S(\hat{\phi}_{1}(q),\phi_{\beta})\Big|_{Q=1}$ $=[z^{0}]\frac{\sqrt{-\Delta^{2}(q)}}{2}U+[z^{-1}]\frac{1}{\sqrt{-\Delta^{2}(q)}}\frac{\partial U}{\partial t^{1}}=-\frac{\mathfrak{h}_{X}(\xi_{1,0})}{\sqrt{-2}},$
$[z^{0}]\sum_{\beta\in\{1,2\}}\tilde{\xi}^{\beta}(z,X)S(\hat{\phi}_{2}(q),\phi_{\beta})\Big|_{Q=1}$ $=[z^{0}]\frac{\sqrt{\Delta^{2}(q)}}{2}U+[z^{-1}]\frac{1}{\sqrt{\Delta^{2}(q)}}\frac{\partial U}{\partial t^{1}}=-\frac{\mathfrak{h}_{X}(\xi_{2,0})}{\sqrt{-2}}.$
Therefore, for $\alpha\in\{1,2\}$, $k>0$,
(11) $\sum_{\beta\in\{1,2\}}\tilde{\xi}^{\beta}(z,X)S(\hat{\phi}_{\alpha}(q),\phi_{\beta})\Big|_{Q=1}$ $=\mathfrak{h}_{X}\Big((-\frac{d}{dx})^{k}\Big(\frac{-1}{\sqrt{-2}}\xi_{\alpha,0}\Big)\Big),$
$\Big(\sum_{\beta\in\{1,2\}}\tilde{\xi}^{\beta}(z,X)S(\hat{\phi}_{\alpha}(q),\phi_{\beta})\Big|_{Q=1}\Big)_{+}$ $=\mathfrak{h}_{X}\Big(\sum_{k\geq 0}\frac{-1}{\sqrt{-2}}W^{\alpha}_{k}z^{k}\Big)=-\mathfrak{h}_{X}\Big(\frac{\hat{\theta}_{\alpha}(z)}{\sqrt{-2}}\Big).$
Lemma 5.2.
We have
$\theta_{\alpha}(z)=\sum_{\beta\in\{1,2\}}\check{R}_{\beta}^{\ \alpha}(-z)\hat{\theta}_{\beta}(z).$
Proof.
The lemma follows from
$d\xi_{\alpha,k}=\sum_{i=0}^{k}\sum_{\beta\in\{1,2\}}([z^{k-i}]\check{R}_{\beta}^{\ \alpha}(-z))W_{i}^{\beta},$
which is shown in the proof of [19, Theorem A]. ∎
Then we have the identification of open leaves:
Theorem 5.3.
For each ordinary leaf $l_{j}$ with $\beta(l_{j})=\beta\in\{1,2\}$ and $k(l_{j})=k\in\mathbb{Z}_{\geq 0}$, we have
$(\tilde{\mathcal{L}}^{O})^{\beta}_{k}(l_{j})\big|_{Q=1}=-(\check{\mathcal{L}}^{O})^{\beta}_{k}(l_{j}).$
Proof.
By [19, Proposition 3.10], the A- and B-model $R$-matrices are equal:
$R_{\beta}^{\ \alpha}(z)\big|_{Q=1}=\check{R}_{\beta}^{\ \alpha}(z).$
Then Theorem 5.3 follows from Equation (5), Equation (11), and Lemma 5.2. ∎
5.3. Mirror symmetry of annulus invariants
Define
$C(Y_{1},Y_{2})$ $=(-\frac{\partial}{\partial x(Y_{1})}-\frac{\partial}{\partial x(Y_{2})})\Big(\frac{\omega_{0,2}}{dx(Y_{1})dx(Y_{2})}\Big)(Y_{1},Y_{2})dx(Y_{1})dx(Y_{2})$
$\ =\Big(-d_{1}\circ\frac{1}{dx(Y_{1})}-d_{2}\circ\frac{1}{dx(Y_{2})}\Big)(\tilde{\omega}_{0,2}(Y_{1},Y_{2})).$
The following proposition is proved in [21, Lemma 6.9].
Proposition 5.4.
We have
$C(Y_{1},Y_{2})=\frac{1}{2}\sum_{\alpha\in\{1,2\}}d\xi_{\alpha,0}(Y_{1})d\xi_{\alpha,0}(Y_{2}).$
Theorem 5.5 (Mirror symmetry of annulus invariants).
$F_{0,2}(\mathbf{t};X_{1},X_{2})=-W_{0,2}(t^{0},q;X_{1},X_{2}).$
Proof.
$\ \ \ \ \ \mathfrak{h}_{X_{1},X_{2}}(C)=\mathfrak{h}_{X_{1},X_{2}}(\frac{1}{2}\sum_{\alpha\in\{1,2\}}d\xi_{\alpha,0}(Y_{1})d\xi_{\alpha,0}(Y_{2}))$
$=-[z_{1}^{0}z_{2}^{0}]\sum_{\alpha,\beta,\gamma\in\{1,2\}}\tilde{\xi}^{\beta}(z_{1},X_{1})\tilde{\xi}^{\gamma}(z_{2},X_{2})S_{z_{1}}(\hat{\phi}_{\alpha}(q),\phi_{\beta})S_{z_{2}}(\hat{\phi}_{\alpha}(q),\phi_{\gamma})\big|_{Q=1}$
$=-[z_{1}^{0}z_{2}^{0}](z_{1}+z_{2})\sum_{\beta,\gamma\in\{1,2\}}V_{z_{1},z_{2}}(\phi_{\beta},\phi_{\gamma})\tilde{\xi}^{\beta}(z_{1},X_{1})\tilde{\xi}^{\gamma}(z_{2},X_{2})\big|_{Q=1}$
$=-\frac{1}{\mathsf{v}}(X_{1}\frac{\partial}{\partial X_{1}}+X_{2}\frac{\partial}{\partial X_{2}})F_{0,2}(\mathbf{t};X_{1},X_{2}).$
By the integrations by parts,
$\mathfrak{h}_{X_{1},X_{2}}(C)$ $=\mathfrak{h}_{X_{1},X_{2}}\Big(\Big(-d_{1}\circ\frac{1}{dx(Y_{1})}-d_{2}\circ\frac{1}{dx(Y_{2})}\Big)\tilde{\omega}_{0,2}\Big)$
$=\Big(\frac{1}{\mathsf{v}}X_{1}\frac{\partial}{\partial X_{1}}+\frac{1}{\mathsf{v}}X_{2}\frac{\partial}{\partial X_{2}}\Big)W_{0,2}(t^{0},q;X_{1},X_{2}).$
Since the constant terms of $F_{0,2}$ and $W_{0,2}$ are both zeros, the theorem follows immediately. ∎
For $(g,m,n)=(0,1,1)$, we define $W_{0,1,1}(t^{0},q;\mathcal{E}_{1};X_{2})$ as the unique Laurent series of $X_{2}$ determined by
$\Big(\frac{1}{\mathsf{v}}X_{2}\frac{d}{dX_{2}}+z_{1}^{-1}\Big)W_{0,1,1}(t^{0},q;\mathcal{E}_{1};X_{2})=\int_{y_{1}\in\mathrm{SYZ}(\mathcal{E}_{1})}e^{x(y_{1})/z_{1}}\mathfrak{h}_{X_{2}}(C).$
Theorem 5.6.
$W_{0,1,1}(t^{0},q;\mathcal{E}_{1};X_{2})=-\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\frac{\kappa_{z_{1}}(\mathcal{E}_{1})}{z_{1}-\psi_{1}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{0,1,1}^{\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1},T}(\mathbf{t},X_{2}).$
Proof.
By [19, Equation (15)], we know
$z\int_{y\in\mathrm{SYZ}(\mathcal{E})}e^{x/z}\frac{d\xi_{\beta,0}}{\sqrt{-2}}=\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\hat{\phi}_{\beta}(q),\frac{\kappa_{z}(\mathcal{E})}{z-\psi}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{0,2}^{\mathbb{P}^{1},T}\big|_{Q=1}.$
Then
$-z_{1}\int_{y_{1}\in\mathrm{SYZ}(\mathcal{E}_{1})}e^{x(y_{1})/z_{1}}\mathfrak{h}_{X_{2}}(C)$
$=\sum_{\alpha\in\{1,2\}}z_{1}\int_{y_{1}\in\mathrm{SYZ}(\mathcal{E}_{1})}\frac{d\xi_{\alpha,0}(Y_{1})}{\sqrt{-2}}\cdot\mathfrak{h}_{X_{2}}(\frac{d\xi_{\alpha,0}(Y_{2})}{\sqrt{-2}})$
$=[z_{2}^{0}]\sum_{\alpha,\gamma\in\{1,2\}}\tilde{\xi}^{\gamma}(z_{2},X_{2})S_{z_{1}}(\hat{\phi}_{\alpha}(q),\kappa_{z_{1}}(\mathcal{E}_{1}))S_{z_{2}}(\hat{\phi}_{\alpha}(q),\phi_{\gamma})\big|_{Q=1}$
$=[z_{2}^{0}]\sum_{\gamma=1,2}\tilde{\xi}^{\gamma}(z_{2},X_{2})\cdot(z_{1}+z_{2})\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\frac{\kappa_{z_{1}}(\mathcal{E}_{1})}{z_{1}-\psi_{1}},\frac{\phi_{\gamma}}{z_{2}-\psi_{2}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{0,2}^{\mathbb{P}^{1},T}\big|_{Q=1}$
$=\Big(\frac{1}{\mathsf{v}}X_{2}\frac{d}{dX_{2}}+z_{1}^{-1}\Big)z_{1}\sum_{\gamma\in\{1,2\},k\geq 0}\tilde{\xi}^{\gamma}_{k}(X_{2})\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\frac{\kappa_{z_{1}}(\mathcal{E}_{1})}{z_{1}-\psi_{1}},\phi_{\gamma}\psi_{2}^{k}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{0,2}^{\mathbb{P}^{1},T}\big|_{Q=1}$
$=\Big(\frac{1}{\mathsf{v}}X_{2}\frac{d}{dX_{2}}+z_{1}^{-1}\Big)z_{1}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\frac{\kappa_{z_{1}}(\mathcal{E}_{1})}{z_{1}-\psi_{1}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}^{\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1},T}_{0,1,1}.$
5.4. All genus mirror symmetry
Theorem 5.7 (All genus mirror theorem (a)).
For any $n>0$ and $g\geq 0$, we have
$F_{g,n}(\mathbf{t};X_{1},\dots,X_{n})=(-1)^{g-1}W_{g,n}(t^{0},q;X_{1},\dots,X_{n}).$
Proof.
For the unstable cases $(g,n)=(0,1)$ and $(0,2)$, this theorem is Theorem 5.1 and Theorem 5.5 respectively.
For stable cases 2g2+n>02g-2+n>0, it suffices to show the A- and B-model graph sum formula in Corollary 3.9 and Theorem 4.2 are equal. By [19, Theorem A], we have
$R_{\beta}^{\ \alpha}(z)\big|_{Q=1}=\check{R}_{\beta}^{\ \alpha}(z),$
and
$\frac{h^{\alpha}_{1}}{\sqrt{2}}=\frac{1}{\sqrt{\Delta^{\alpha}(q)}}.$
Hence, the weights in the graph sum match except for the open leaves. Since Theorem 5.3 identifies the A- and B- open leaves, the theorem follows immediately. ∎
Theorem 5.8 (All genus mirror theorem (b)).
Suppose $g,m,n\geq 0$, for any $\mathcal{E}_{1},\dots,\mathcal{E}_{m}\in K_{T}(\mathbb{P}^{1})$, we have
$W_{g,m,n}(t^{0},q;\mathcal{E}_{1},\dots,\mathcal{E}_{m};X_{1},\dots,X_{n})$
$=(-1)^{g-1}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\frac{\kappa_{z_{1}}(\mathcal{E}_{1})}{z_{1}-\psi_{1}},\dots,\frac{\kappa_{z_{m}}(\mathcal{E}_{m})}{z_{m}-\psi_{m}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{g,m,n}^{\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1},T}(\mathbf{t},X_{1},\dots,X_{n}).$
Proof.
For the case $n=0$, this theorem is Theorem [19, Theorem B]. For the case $m=0$, this theorem is Theorem 5.7. For the unstable case $(g,m,n)=(0,1,1)$, this theorem is Theorem 5.6.
For the stable case, it suffices to show the A- and B-model graph sum formula in Theorem 3.11 and (10) are equal. Following the proof of Theorem 5.7, it remains to show the $\mathrm{SYZ}$ leaf matches. By [19, Equation (15)], we know
$\int_{y\in\mathrm{SYZ}(\mathcal{E})}e^{x(y)/z}\frac{W^{\beta}_{i}}{\sqrt{-2}}=-z^{-i-1}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\hat{\phi}_{\beta}(q),\frac{\kappa_{z}(\mathcal{E})}{z-\psi}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{0,2}^{\mathbb{P}^{1},T}\Big|_{Q=1}.$
By Lemma 5.2, we have
$(\check{\mathcal{L}}^{\mathrm{SYZ}})^{\beta(l_{j})}_{k(l_{j})}(l_{j})=-\sum_{i=0}^{k}\sum_{\gamma\in\{1,2\}}([z^{k-i}]\check{R}_{\gamma}^{\ \beta}(-z))z_{j}^{-i-1}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\hat{\phi}_{\gamma}(q),\frac{\kappa_{z_{j}}(\mathcal{E}_{j})}{z_{j}-\psi}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{0,2}^{\mathbb{P}^{1},T}\Big|_{Q=1}.$
In A-model, ${\bf u}_{j}(z)=\sum_{\alpha\in\{1,2\}}{\bf u}_{j}^{\alpha}(z)\phi_{\alpha}(q)$ is replaced by
(12) ${\bf u}_{j}^{\alpha}(z)=\sum_{a\geq 0}\big(\phi^{\alpha}(q),\kappa_{z_{j}}(\mathcal{E}_{j})\big)_{\mathbb{P}^{1},T}z_{j}^{-a-1}z^{a},$
so
$[z^{i}]\left(\frac{{\bf u}_{j}^{\alpha}(z)}{\sqrt{\Delta^{\alpha}(q)}}S^{\hat{\underline{\gamma}}}_{\ \hat{\underline{\alpha}}}(z)\right)_{+}$
$=\sum_{b\geq 0}z_{j}^{-i-b-2}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\hat{\phi}_{\gamma}(q),\tau_{b}(\hat{\phi}_{\alpha}(q))\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{0,2}^{\mathbb{P}^{1},T}\big(\hat{\phi}_{\alpha}(q),\kappa_{z_{j}}(\mathcal{E}_{j})\big)_{\mathbb{P}^{1},T}$
$=z_{j}^{-i-1}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\hat{\phi}_{\gamma}(q),\frac{\kappa_{z_{j}}(\mathcal{E}_{j})}{z_{j}-\psi}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{0,2}^{\mathbb{P}^{1},T}.$
In other words, for the ${\bf u}_{j}(z)$ corresponding to $\frac{\kappa_{z_{j}}(\mathcal{E}_{z_{j}})}{z_{j}-\psi}$, we have
$(\mathcal{L}^{\bf u})_{k(l_{j})}^{\beta(l_{j})}(l_{j})\Big|_{Q=1}=-(\check{\mathcal{L}}^{\mathrm{SYZ}})^{\beta(l_{j})}_{k(l_{j})}(l_{j}).$
The theorem follows immediately. ∎
6. Geometric open Gromov-Witten theory of $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$
6.1. Geometric counting via coherent boundary conditions
In previous sections, we established the counting of stable maps from bordered Riemann surfaces to $(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1})$ by means of integration over the fixed locus of the moduli space $\overline{\mathcal{M}}_{(g,h),n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})$. We further demonstrated that these virtual invariants are systematically encoded by the topological recursion, via the local expansion of the multidifferential forms $\omega_{g,n}$ on the boundary of mirror curve $C_{q}$.
In symplectic geometry, the enumeration of stable maps by means of integration over the global moduli space $\overline{\mathcal{M}}_{(0,h),n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})$ has also garnered significant interest. We refer to this global enumeration as geometric counting. One approach defines it via multisections intersection, while another employs the coherent boundary conditions and the localization formula for the orbifold with corners. In the non-equivariant limit, the construction using the coherent boundary conditions reduces to the non-equivariant counting defined by sections of bundles.
The geometric counting contains two contribution: one arises from the integration over the fixed locus $F$, which is the virtual counting defined in Section 3.2; the other one is the corner contribution.
The work [6, Theorem 2.18, Definition 5.9] proposes a combinatorial formula for the geometric counting, which is expressed solely in terms of fixed-point graphs . In the genus 0 case, by carefully adjusting the coefficients preceding each fixed-point graph, they incorporate the corner contribution into the formula that comprises only fixed-point loci. Meanwhile, the authors conjecture that there exists a geometric construction via coherent boundary conditions whereby this combinatorial formula holds for geometric counting in all genera. If this conjecture is proven, the counting of fixed-point graphs, (i.e., the virtual counting we define in Section 3.2), will serve as the fundamental building block for calculating geometric intersection numbers. Our Mirror Theorem 5.7 and 5.8 provide a recursive algorithm for geometric counting.
6.2. Combinatorial formula
Given the moduli space $\overline{\mathcal{M}}_{(g,h),n}(\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1}|\beta^{\prime},\vec{\mu})$, one can associate a decorated graph $S=(V_{b}\sqcup V_{w},E,g,{\sf{I}},\vec{d},\mu)$, where
  • (i)
    $V_{b}$ is the set of black vertex, $V_{w}$ is the set of white vertex, and $E$ is the set of edges. There is exactly one black vertex and $l(\vec{\mu})$ white vertices. There are exactly a single edge connecting the black vertex and each white vertex.
  • (ii)
    The black vertex is decorated by $g,{\sf I},\vec{d}$, corresponding the genus, marked points and the degree of the moduli space.
  • (iii)
    Each white vertex is decorated by a unique entry $\mu_{i}$ in $\vec{\mu}$. The white vertices are unordered, so the decoration is unique.
A moduli specification $S$ is the disjoint union of such decorated graphs. An isomorphism between two moduli specifications is a graph isomorphism which respects the decorations.
Given $\vec{a}=(a_{i})_{i\in{\sf I}}$ and $\vec{\gamma}=(\gamma_{i})_{i\in{\sf I}}$, where $a_{i}\in\mathbb{Z}_{\geq 0}$ and $\gamma_{i}\in H^{*}_{T}(\mathbb{P}^{1})$. If the moduli specification $S$ is connected, we define
$I(S,\vec{a},\vec{\gamma})=\langle\tau_{a_{1}}(\gamma_{1})\dots\tau_{a_{n}}(\gamma_{n})\rangle^{\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1},T}_{g,\beta^{\prime},\vec{\mu}}.$
For disconnected moduli specification $S$ with finitely many connected components $S_{i}$, we define
$I(S,\vec{a},\vec{\gamma})=\prod_{i}I(S_{i},\vec{a}|_{S_{i}},\vec{\gamma}|_{S_{i}}).$
A moduli specification is pure if $\mu(w)\neq 0$ for each white vertex $w$.
Definition 6.1.
A morphism from a moduli specification $S^{new}$ to a moduli specification $S^{old}$ is a decorated graph $M=(V_{b}\sqcup V_{w},E,H^{CB},\widetilde{E},g,{\sf I},\vec{d},\mu,\sigma,\mu^{old})$ such that
  • (a)
    $(V_{b}\sqcup V_{w},E,g,{\sf I},\vec{d},\mu)$ is the moduli specification $S^{new}$.
  • (b)
    The wavy edges $\widetilde{E}$ connect white vertices.
  • (c)
    The contracted boundary half-edges $H^{CB}$ are half-edges which emanate from black vertices.
  • (d)
    A wavy flag is a pair of a wavy edge and an incident white vertex:
    $\{(\tilde{e},v)\in\widetilde{E}\times V_{w}:v\in\tilde{e}\}$.
    For any white vertex $v$, let $\sigma_{v}$ be a cyclic order of the wavy flags emanate from $v$.
  • (e)
    An involution $\tau$ on wavy flags: if $v,v^{\prime}$ are two incident vertices of $\tilde{e}\in\tilde{E}$, then $\tau(\tilde{e},v):=(\tilde{e},v^{\prime})$.
  • (f)
    The cyclic order $\sigma$ and the involution $\tau$ form the set of cycles of $\sigma^{-1}\circ\tau$:
    $\quad\{(v_{1},\tilde{e}_{1},v_{2},\dots,\tilde{e}_{m-1},v_{m})~|~v_{i}\in V_{w},\tilde{e}_{i}\in\widetilde{E},v_{i},v_{i+1}\in\tilde{e}_{i},v_{1}=v_{m}\}/\sim,$
    where the equivalence relation $\sim$ identifies tuples that differ by a cyclic permutation. Here for the white vertex without incident wavy edge, we assign the cycle to be its own. The function $\mu^{old}$ from the set of cycles of $\sigma^{-1}\circ\tau$ to $\mathbb{Z}$.
  • (g)
    $S^{old}$ is the desingularization of $M$, where the desingularization is the moduli specification as follows.
    • $\bullet$
[…]
    • $\bullet$
      For such a black vertex $v^{\prime}$, we define
      $\quad\quad\vec{d}(v^{\prime})=\sum_{v}\vec{d}(v),\quad{\sf I}(v^{\prime})=\bigsqcup_{v}{\sf I}(v),\quad g(v^{\prime})=h^{1}(G)+\sum_{v}g(v),$
      where we sum or union over the black vertices $v$ in $G$ associated to the connected component represented by $v^{\prime}$.
    • $\bullet$
      For a cycle $c$ from the set of $\sigma^{-1}\circ\tau$ in the connected component corresponding to $v^{\prime}$, we assign a white vertex $w$ connecting $v^{\prime}$, with $\mu(w)=\mu^{old}(c)$.
    • $\bullet$
      For each contracted boundary half-edge $h$ in the connected component corresponding to $v^{\prime}$, we assign a white vertex $w$ connecting $v^{\prime}$, with $\mu(w)=0$.
Let $\mathrm{Hom}(S^{new},S^{old})$ be the set of isomorphism types of morphisms from $S^{new}$ to $S^{old}$.
For a contracted boundary half-edge $h$ of $M$, which is attached to a black vertex $v$, we can assign a degree $d$ to it:
$d(h):=d_{+}(v)-\sum_{w\in V_{w}(v):\mu(w)>0}\mu(w),$
where $V_{w}(v)$ is the set of white vertices attached to $v$.
In order to state the combinatorial formula proposed in [6], we also introduce the boundary contribution graphs.
Definition 6.2.
A boundary contribution (BC) graph $G=(V,E,L,\mu)$ consists of the following data:
  • (a)
    The set of vertices $V$. The set of edges $E=\vec{E}\sqcup\widetilde{E}$, where $\vec{E}$ is the set of directed edges and $\widetilde{E}$ is the set of wavy edges. The set of the oriented loops $L$. It requires that every vertex has exactly one incoming edge and one outgoing edge. Every wavy edges connects two different vertices.
  • (b)
[…]
  • (c)
    The perimeter function $\mu:F^{new}\cup F^{old}\rightarrow\mathbb{Z}$.
Definition 6.3.
A metric on a boundary contribution graph $G$ is an assignment of length $x_{e}\in\mathbb{R}$ for every $e\in\vec{E}$ such that
  • (a)
    The sum of lengths of the oriented edges of any circuit of the graph is an integer.
  • (b)
    The sum of lengths of the oriented edges of a new or old face is the perimeter of that face.
  • (c)
    If $e\in\vec{E}$ is an edge of a new face $f$, then $x_{e}\mu(f)>0$.
Let $W_{G}\subset\mathbb{R}^{\vec{E}}$ be the space of metrics on $G$ and $\overline{W}_{G}$ be its closure in $\mathbb{R}^{\vec{E}}$. Let $\mathcal{E}\subset\vec{E}$ be a set of directed edges such that $\mathcal{E}\sqcup\widetilde{E}$ forms a spanning forest of $G$. We define
$\Omega_{G}=\bigwedge_{e\in\mathcal{E}}dx_{e},$
and define the volume of a BC graph $G$ as
$\text{Vol}_{G}=(-1)^{alt}\left|\Bigg(\prod_{f\in F^{new}\backslash L}\mu(f)\Bigg)\int_{\overline{W}_{G}}\Omega_{G}\right|,$
where altalt is the number of edges of $\widetilde{E}$ which touch two new faces $f_{1},f_{2}$ with $\mu(f_{1})\mu(f_{2})<0$.
We can assign a boundary contribution graph for morphism $M\in\mathrm{Hom}(S^{new},S^{old})$ in the following way (See Figure 3).
Definition 6.4.
Given a morphism
$M=(V_{b}\sqcup V_{w},E,H^{CB},\widetilde{E},g,{\sf I},\vec{d},\mu,\sigma,\mu^{old})\in\mathrm{Hom}(S^{new},S^{old}),$
one can assign a BC graph, denoted by $BC(M)=(V,E,L,\mu)$ as follows:
  • For any wavy flag $(\tilde{e},v)$, assign a vertex.
  • For the wavy flags $(\tilde{e},v),(\tilde{e},v^{\prime})$, where $\tilde{e}$ connects two vertices $v,v^{\prime}$, one assign a wavy edge connecting the corresponding vertices.
  • Draw the directed edges $\vec{e}$ with respect to the cycles $\sigma_{v}$.
  • Assign a loop for the isolated white vertex.
  • So far, we get a graph $G$. The new faces $f^{new}$ of $G$ corresponds to the white vertices $w$ of $M$, and we assign $\mu(f^{new})=\mu(w)$. The old faces $f^{old}$ corresponds to the cycles $c$ of type $\sigma^{-1}\circ\tau$, and we assign $\mu(f^{old})=\mu^{old}(c)$.
[…]
Figure 3. boundary contribution graph
The left hand side is an example of morphism between moduli specification (just show white vertices and wavy edges); the right hand side is the associated boundary contribution graph. Here $\tilde{e}_{ij}$ is the assigned vertex of the wavy flag from $i$ to $j$.
Now we state the combinatorial formula in [6].
Definition 6.5.
Let $S$ be a moduli specification, $\vec{a},\vec{\gamma}$ be vectors of nonnegative integers and cohomology classes in $H^{*}_{T}(\mathbb{P}^{1})$. Define the combinatorial open Gromov-Witten invariants of $(S,\vec{a},\vec{\gamma})$ as:
$\text{OGW}(S,\vec{a},\vec{\gamma})=\sum_{S^{\prime}\text{ is pure}}\sum_{M\in\mathrm{Hom}(S^{\prime},S)}\frac{1}{|\mathrm{Aut}(M)|}\frac{\text{Vol}_{BC(M)}}{(-\mathsf{v})^{|\widetilde{E}|}}\left(\prod_{h\in H^{CB}(M)}\frac{d(h)}{\mathsf{v}}\right)I(S^{\prime},\vec{a},\vec{\gamma}).$
The main conjecture in [6] states that
Conjecture 6.6.
There exists a geometric definition for $\text{OGW}(S,\vec{a},\vec{\gamma})$ for all moduli specifications $S$.
Remark 6.7.
It is proved in [6] that the coherent boundary condition serves as the geometric definition of $\text{OGW}(S,\vec{a},\vec{\gamma})$ in genus 0.
6.3. Generating function of geometric counting
Let $S=(V_{b}\sqcup V_{w},E,g,{\sf{I}},\vec{d},\mu)$ be a moduli specification, $\vec{a}=(a_{i})_{i\in{\sf I}}$ and $\vec{\gamma}=(\gamma_{i})_{i\in{\sf I}}$. We could change the decoration of $S$ in the following two ways:
  • (1)
    (add marked points) For every $l\in\mathbb{Z}_{\geq 1}$, we define $S^{[l]}$ by extending the set of labels on the black vertices to ${\sf I}^{[l]}={\sf I}\sqcup\{1,\dots,l\}$, while keeping all other decorations unchanged.
    The vector $\vec{a}^{[l]}$ is extended accordingly by setting $\vec{a}^{[l]}_{i}=a_{i}$ for $i\in{\sf I}$ and $\vec{a}^{[l]}_{i}=0$ for $i\in\{1,\dots,l\}$. The vector $\vec{\gamma}^{[l]}$ is extended by setting $\vec{\gamma}^{[l]}=\gamma_{i}$ for $i\in{\sf I}$ and $\vec{\gamma}^{[l]}_{i}=\mathbf{t}$ for $i\in\{1,\dots,l\}$, where $\mathbf{t}\in H^{*}_{T}(\mathbb{P}^{1})$.
  • (2)
    (forget degree) For a black vertex $v\in V_{b}$, its decoration $d_{v}=d(v)$ corresponding to the curve degree $\beta\in H_{2}(\mathbb{P}^{1})$ is given by
    $d(v)=d_{+}(v)-\sum_{w\in V_{w}(v):\mu(w)>0}\mu(w),$
    where $V_{w}(v)$ is the set of white vertices attached to $v$.
    By forgetting the decoration $\vec{d}$, we obtain from $S$ a decorated graph
    $\Gamma_{S}=(V_{b}\sqcup V_{w},E,g,{\sf{I}},\mu).$
    We say $\Gamma_{S}$ is pure if $\mu(w)\neq 0$ for each white vertex $w$.
    Conversely, given a vector $(d_{v})_{v\in V_{b}}$, we define $\Gamma_{S}[d_{v}]$ as the moduli specification by assigning $\vec{d}$ as follows
    (13) $\qquad d_{+}(v)=d_{v}+\sum_{w\in V_{w}(v):\mu(w)>0}\mu(w),\quad d_{-}(v)=d_{v}-\sum_{w\in V_{w}(v):\mu(w)<0}\mu(w).$
    Different vectors $(d_{v})$ may yield the same moduli specification (i.e. indistinguishable at the level of the decorated graph $S$). We therefore define $d_{v}]$ as the equivalence class of all such $(d_{v})$ that determine the same moduli specification.
Assume that $S$ is a connected moduli specification, we would like to define the generating function of geometric counting as:
$\text{OGW}(\Gamma_{S},\vec{a},\vec{\gamma},\mathbf{t}):=\sum_{d\geq 0}\sum_{l\geq 0}\frac{1}{l!}\text{OGW}(\Gamma_{S}[d]^{[l]},\vec{a}^{[l]},\vec{\gamma}^{[l]}),$
where $\Gamma_{S}[d]^{[l]}$ is the moduli specification by adding the degree $d$ and extending $l$ marked labels on the unique black vertex. The dependence of $\text{OGW}(\Gamma_{S},\vec{a},\vec{\gamma},\mathbf{t})$ on $\mathbf{t}$ is reflected in $\vec{\gamma}^{[l]}$, where $\mathbf{t}$ is assigned to $\{1,\dots,l\}$.
Furthermore, assume $S$ is pure ($\mu$ is non-zero on white vertices), we would like to use the combinatorial formula in Definition 6.5 to express $\text{OGW}(\Gamma_{S},\vec{a},\vec{\gamma},\mathbf{t})$.
We introduce the following notations:
  • For two moduli specifications $S^{\prime},S$, we can construct a morphism $\Gamma_{M}\in\mathrm{Hom}(\Gamma_{S^{\prime}},\Gamma_{S})$ by dropping the condition on the equality of $d_{v}$, following the pattern of Definition 6.1.
  • For a morphism $M\in\mathrm{Hom}(S^{\prime},S)$, the association of the boundary contribution graph $BC(M)$ is independent of the decoration $d_{v}$ on the black vertices of $M$, so we can also define the boundary contribution graph for $\Gamma_{M}\in\mathrm{Hom}(\Gamma_{S^{\prime}},\Gamma_{S})$, denoted by $BC(\Gamma_{M})$.
  • Let $v$ be a black vertex of $\Gamma_{M}$, and let $g_{v},{\sf I}_{v},V_{w}(v)$ be the genus, marked labels, the set of white vertices attached to $v$, respectively. We define
    $I(\vec{a}|_{v},\vec{\gamma}|_{v},\mathbf{t})=[\prod_{i\in V_{w}(v)}X_{i}^{\mu_{i}}]\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\tau_{a_{i}}(\gamma_{i})|_{i\in{\sf I}_{v}}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}_{g_{v},|{\sf I}_{v}|,|V_{w}(v)|}^{\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1},T}.$
Then we have the following theorem.
Theorem 6.8.
Let $S$ be a pure connected moduli specification. Then we have
$\text{OGW}(\Gamma_{S},\vec{a},\vec{\gamma},\mathbf{t})=\sum_{\Gamma_{S^{\prime}}\text{ is pure}}\sum_{\Gamma_{M}\in\mathrm{Hom}(\Gamma_{S^{\prime}},\Gamma_{S})}\frac{1}{|\mathrm{Aut}(\Gamma_{M})|}\frac{\text{Vol}_{BC(\Gamma_{M})}}{(-\mathsf{v})^{|\widetilde{E}|}}\prod_{v\in V_{b}(M)}I(\vec{a}|_{v},\vec{\gamma}|_{v},\mathbf{t}).$
Proof.
The proof follows from the enumeration of $M\in\mathrm{Hom}(S^{\prime},S)$ and the Orbit-Stabilizer theorem.
In Step 1, we will sum over the extended $l$ marked labels.
When $S$ is pure, $M\in\mathrm{Hom}(S^{\prime},S)$ doesn’t have the contracted boundary half-edges $H^{CB}$. The combinatorial formula states that
$\text{OGW}(S^{[l]},\vec{a}^{[l]},\vec{\gamma}^{[l]})=\sum_{S^{\prime}_{l}\text{ is pure}}\sum_{M_{l}\in\mathrm{Hom}(S^{\prime}_{l},S^{[l]})}\frac{\text{Vol}_{BC(M_{l})}}{(-\mathsf{v})^{|\widetilde{E}|}}\frac{I(S^{\prime}_{l},\vec{a}^{[l]},\vec{\gamma}^{[l]})}{|\mathrm{Aut}(M_{l})|}.$
For $M_{l}\in\mathrm{Hom}(S^{\prime}_{l},S^{[l]})$, there exists a forgetful map
$\text{for}_{l}:\mathrm{Hom}(S^{\prime}_{l},S^{[l]})\rightarrow\mathrm{Hom}(S^{\prime},S)$
defined by forgetting $\{1,\dots,l\}$ in ${\sf I}^{[l]}$. Let $M=\text{for}_{l}(M_{l})$.
Let $V_{b}(M)$ be the set of black vertices of $M$, and let $v^{[l]}=(v_{1},\dots,v_{l})$ be the vector of black vertices decorated by $i\in\{1,\dots,l\}$. There is a canonical $\mathrm{Aut}(M)$-action on $V_{b}(M)^{l}$ (length-$l$ vectors of $V_{b}(M)$) defined as
$\sigma\cdot(v_{1},\dots,v_{l}):=(\sigma(v_{1}),\dots,\sigma(v_{l})),\quad\sigma\in\mathrm{Aut}(M).$
By definition, the automorphism group $\mathrm{Aut}(M_{l})$ is a subgroup of $\mathrm{Aut}(M)$ that fixes black vertices labeled by $\{1,\dots,l\}$. In other words, $\mathrm{Aut}(M_{l})$ is the stablizer of $v^{[l]}$. By the orbit-stablizer theorem, we have
$\frac{|\mathrm{Aut}(M)|}{|\mathrm{Aut}(M_{l})|}=|\text{Orb}(v^{[l]})|.$
Two assignments $v_{1}^{[l]},v_{2}^{[l]}$ on $M$ gives the same decoration $M_{l}\in\text{for}^{-1}_{l}(M)$ if and only if $v_{1}^{[l]},v_{2}^{[l]}$ are in the same orbit of $\mathrm{Aut}(M)$. In other words, there is a one-to-one correspondence between the set $\text{for}^{-1}_{l}(M)$ and the set of orbits $V_{b}(M)^{l}/\mathrm{Aut}(M)$.
[…]
$\text{OGW}(S^{[l]},\vec{a}^{[l]},\vec{\gamma}^{[l]})=\sum_{S^{\prime}\text{ is pure}\atop{M\in\mathrm{Hom}(S^{\prime},S)}}\frac{\text{Vol}_{BC(M)}}{(-\mathsf{v})^{|\widetilde{E}|}}\sum_{M_{l}\in\text{for}_{l}^{-1}(M)}\frac{I(S^{\prime}_{l}(M_{l}),\vec{a}^{[l]},\vec{\gamma}^{[l]})}{|\mathrm{Aut}(M_{l})|}$
$=\sum_{S^{\prime}\text{ is pure}\atop{M\in\mathrm{Hom}(S^{\prime},S)}}\frac{1}{|\mathrm{Aut}(M)|}\frac{\text{Vol}_{BC(M)}}{(-\mathsf{v})^{|\widetilde{E}|}}\sum_{v^{[l]}\in V_{b}(M)^{l}}I(S^{\prime}_{v^{[l]}},\vec{a}^{[l]},\vec{\gamma}^{[l]})$ ,
where $S^{\prime}_{v^{[l]}}$ is the morphism obtained by labelling $\{1,\dots,l\}$ to $v_{i}$’s in $S^{\prime}$. Moreover, we have
$\sum_{l\geq 0}\frac{1}{l!}\text{OGW}(S^{[l]},\vec{a}^{[l]},\vec{\gamma}^{[l]})$
$=\sum_{S^{\prime}\text{ is pure}\atop{M\in\mathrm{Hom}(S^{\prime},S)}}\frac{1}{|\mathrm{Aut}(M)|}\frac{\text{Vol}_{BC(M)}}{(-\mathsf{v})^{|\widetilde{E}|}}\prod_{v\in V_{b}(M)}I(\vec{a}|_{v},\vec{\gamma}|_{v},d_{v},\mathbf{t}),$
where
$I(\vec{a}|_{v},\vec{\gamma}|_{v},d_{v},\mathbf{t}):=\sum_{l\geq 0}\frac{1}{l!}\langle\tau_{a_{i}}(\gamma_{i})|_{i\in{\sf I}_{v}},\mathbf{t}^{l}\rangle_{g_{v},\vec{d}(v),\vec{\mu}(v)}^{\mathbb{P}^{1},\mathbb{R}\mathbb{P}^{1},T},$
$\vec{d}(v)$ is given by (13), and $\vec{\mu}(v)$ is the vector of the values of $\mu$ on the white vertices attached to $v$.
In Step 2, we sum over all curve degree $d$.
Let $\Gamma_{M}\in\mathrm{Hom}(\Gamma_{S^{\prime}},\Gamma_{S})$ and $(d_{v})$ be the vector of decoration of curve degree on the black vertex $V_{b}(M)$. We write $\Gamma_{M}[d_{v}]\in\mathrm{Hom}(\Gamma_{S^{\prime}}[d_{v}],\Gamma_{S}[d])$, where $d=\sum_{v\in V_{b}(\Gamma_{M})}d_{v}$.
Let $m=|V_{b}(\Gamma_{M})|$, and let $(v_{1},\dots,v_{m})$ be the vector of black vertices of $\Gamma_{M}$. There is an $\mathrm{Aut}(\Gamma_{M})$ action on $(d_{v_{1}},\dots,d_{v_{m}})$:
$\sigma\cdot(d_{v_{1}},\dots,d_{v_{m}}):=(d_{\sigma(v_{1})},\dots,d_{\sigma(v_{m})}),\quad\sigma\in\mathrm{Aut}(\Gamma_{M}).$
Then $\mathrm{Aut}(\Gamma_{M}[d_{v}])=\text{Stab}_{\mathrm{Aut}(\Gamma_{M})}(d_{v})$. Then the Orbit-Stabilizer theorem tells that
$\frac{|\mathrm{Aut}(\Gamma_{M})|}{|\mathrm{Aut}(\Gamma_{M}[d_{v}])|}=|\text{Orb}((d_{v}))|.$
Let $d_{v}]$ be the equivalent classes of $(d_{v})$ under the action $\mathrm{Aut}(\Gamma_{M})$, which are 1-1 corresponding to the morphisms $\Gamma_{M}[d_{v}]\in\mathrm{Hom}(\Gamma_{S^{\prime}}[d_{v}],\Gamma_{S}[d])$ with $d=\sum_{v\in V_{b}(\Gamma_{M})}d_{v}$. We have
$\text{OGW}(\Gamma_{S},\vec{a},\vec{\gamma},\mathbf{t})$
$=\sum_{\Gamma_{S^{\prime}}\text{ is pure}}\sum_{\Gamma_{M}\in\mathrm{Hom}(\Gamma_{S^{\prime}},\Gamma_{S})}\frac{\text{Vol}_{BC(\Gamma_{M})}}{(-\mathsf{v})^{|\widetilde{E}|}}\sum_{[d_{v}]}\frac{1}{|\mathrm{Aut}(\Gamma_{M}[d_{v}])|}\prod_{v\in V_{b}(M)}I(\vec{a}|_{v},\vec{\gamma}|_{v},d_{v},\mathbf{t})$
$=\sum_{\Gamma_{S^{\prime}}\text{ is pure}}\sum_{\Gamma_{M}\in\mathrm{Hom}(\Gamma_{S^{\prime}},\Gamma_{S})}\frac{1}{|\mathrm{Aut}(\Gamma_{M})|}\frac{\text{Vol}_{BC(\Gamma_{M})}}{(-\mathsf{v})^{|\widetilde{E}|}}\sum_{d_{v}}\prod_{v\in V_{b}(M)}I(\vec{a}|_{v},\vec{\gamma}|_{v},d_{v},\mathbf{t})$
$=\sum_{\Gamma_{S^{\prime}}\text{ is pure}}\sum_{\Gamma_{M}\in\mathrm{Hom}(\Gamma_{S^{\prime}},\Gamma_{S})}\frac{1}{|\mathrm{Aut}(\Gamma_{M})|}\frac{\text{Vol}_{BC(\Gamma_{M})}}{(-\mathsf{v})^{|\widetilde{E}|}}\prod_{v\in V_{b}(M)}I(\vec{a}|_{v},\vec{\gamma}|_{v},\mathbf{t}),$
which complete the proof. ∎
6.4. B-model geometric counting
Let $W_{g,m,\vec{\mu}}(\mathbf{t},\mathcal{E}_{i})$ be the coefficients of the Laurent series $W_{g,m,n}(t^{0},q;\mathcal{E}_{i};X_{j})$ (see (9)):
$W_{g,m,n}(t^{0},q;\mathcal{E}_{1},\dots,\mathcal{E}_{m};{X}_{1},\dots,{X}_{n})=\sum_{\vec{\mu}=(\mu_{1},\dots,\mu_{n})\atop{\mu_{i}\in\mathbb{Z}_{\neq 0}}}W_{g,m,\vec{\mu}}(\mathbf{t},\mathcal{E}_{i})X_{1}^{\mu_{1}}\dots X_{n}^{\mu_{n}}.$
Let $\Gamma_{S}=(V_{b}\sqcup V_{w},E,g,{\sf{I}},\mu)$ be a connected decorated graph with one black vertex, $n$ white vertices and $m$ marked labels $\sf I$}. We define a graph sum based on the summation over $\Gamma_{M}\in\mathrm{Hom}(\Gamma_{S^{\prime}},\Gamma_{S})$. On the marked label $\sf I$}, we assign the classes $\frac{\kappa_{z_{i}}(\mathcal{E}_{i})}{z_{i}-\psi_{i}}$, $i\in{\sf I}$. In A-model, we define $\text{OGW}_{g,m,n}(\Gamma_{S},\frac{\kappa_{z_{i}}(\mathcal{E}_{i})}{z_{i}-\psi_{i}},\mathbf{t})$ as a generating function of open geometric counting with $\vec{a},\vec{\gamma}$ corresponding to the class $\frac{\kappa_{z_{i}}(\mathcal{E}_{i})}{z_{i}-\psi_{i}}$:
$\text{OGW}_{g,m,n}(\Gamma_{S},\frac{\kappa_{z_{i}}(\mathcal{E}_{i})}{z_{i}-\psi_{i}},\mathbf{t}):=\sum_{\vec{a}=(a_{1},\dots,a_{m})\atop{a_{i}\in\mathbb{Z}_{\geq 0}}}\text{OGW}(\Gamma_{S},\vec{a},(\kappa_{z_{i}}(\mathcal{E}_{i}))_{i\in{\sf I}},\mathbf{t})\prod_{i=1}^{m}z_{i}^{-a_{i}-1}.$
In B-model, we define
$G_{g,m,n}(\Gamma_{S},\mathbf{t})=\sum_{\Gamma_{S^{\prime}}\text{ is pure}}$ $\sum_{\Gamma_{M}\in\mathrm{Hom}(\Gamma_{S^{\prime}},\Gamma_{S})}\frac{1}{|\mathrm{Aut}(\Gamma_{M})|}\frac{\text{Vol}_{BC(\Gamma_{M})}}{(-\mathsf{v})^{|\widetilde{E}|}}$
$\times\prod_{v\in V_{b}(M)}(-1)^{g_{v}-1}W_{g_{v},|{\sf I}_{v}|,\vec{\mu}(v)}(\mathbf{t},\mathcal{E}_{i}|i\in{\sf I}_{v}),$
By Theorem 5.8 and Theorem 6.8, we have
Theorem 6.9.
Let $S$ be a pure connected moduli specification with one black vertex, $n$ white vertices and $m$ marked labels $\sf I$}, it holds that
$\text{OGW}_{g,m,n}(\Gamma_{S},\frac{\kappa_{z_{i}}(\mathcal{E}_{i})}{z_{i}-\psi_{i}},\mathbf{t})=G_{g,m,n}(\Gamma_{S},\mathbf{t}).$
Appendix A Bessel functions
The special function $I_{\alpha}(x)$ in $J$-function is the modified Bessel function of the first kind. It is defined as
$I_{\alpha}(x)=\sum_{m=0}^{\infty}\frac{1}{m!\Gamma(m+\alpha+1)}(\frac{x}{2})^{2m+\alpha}.$
The derivative of $I_{\alpha}(x)$ is given by
$I^{\prime}_{\alpha}(x)=I_{\alpha+1}(x)+\frac{\alpha}{x}I_{\alpha}(x).$
Let $d\in\mathbb{Z}_{\neq 0}$ and $q,\mathsf{v}>0$, we get
$I^{\prime}_{d}\Big(\frac{2\sqrt{q}d}{\mathsf{v}}\Big)=I_{d+1}\Big(\frac{2\sqrt{q}d}{\mathsf{v}}\Big)+\frac{\mathsf{v}}{2\sqrt{q}}I_{d}\Big(\frac{2\sqrt{q}d}{\mathsf{v}}\Big).$
For $n\in\mathbb{N}$,
$I_{n}(x)=I_{-n}(x).$
We have
$I_{0}(0)=1,\quad I_{\alpha}(0)=0\quad(\alpha>0).$
Appendix B Residue Calculation
Let $\mu\in\mathbb{Z}$, $q,\mathsf{v}>0$, consider
$C_{\mu}:=\mathrm{Res}_{Y=0}e^{\mu(Y+\frac{q}{Y})/\mathsf{v}}\frac{dY}{Y^{\mu+1}},$
$D_{\mu}:=\mathrm{Res}_{Y=0}e^{\mu(Y+\frac{q}{Y})/\mathsf{v}}\frac{dY}{Y^{\mu+2}}.$
To calculate $C_{\mu}$ and $D_{\mu}$, we expand
$e^{\mu(Y+q/Y)/\mathsf{v}}=\sum_{m,n=0}^{\infty}\frac{(\frac{\mu}{\mathsf{v}})^{n}(\frac{\mu q}{\mathsf{v}})^{m}Y^{n-m}}{n!m!}.$
For $C_{\mu},\ \mu\geq 0$, let $n=m+\mu$,
$C_{\mu}$ $=\sum_{m=0}^{\infty}\frac{(\frac{\mu}{\mathsf{v}})^{m+\mu}(\frac{\mu q}{\mathsf{v}})^{m}}{(m+\mu)!m!}$
$=\frac{1}{\sqrt{q}^{\mu}}\sum_{m=0}^{\infty}\frac{1}{m!(m+\mu)!}\Big(\frac{\mu\sqrt{q}}{\mathsf{v}}\Big)^{2m+\mu}$
$=\frac{1}{\sqrt{q}^{\mu}}I_{\mu}\Big(\frac{2\sqrt{q}\mu}{\mathsf{v}}\Big).$
For $C_{\mu},\ \mu<0$, let $m=n-\mu$,
$C_{\mu}$ $=\sum_{n=0}^{\infty}\frac{(\frac{\mu}{\mathsf{v}})^{n}(\frac{\mu q}{\mathsf{v}})^{n-\mu}}{n!(n-\mu)!}$
$=\frac{1}{\sqrt{q}^{\mu}}\sum_{n=0}^{\infty}\frac{1}{n!(n-\mu)!}\Big(\frac{\mu\sqrt{q}}{\mathsf{v}}\Big)^{2n-\mu}$
$=\frac{1}{\sqrt{q}^{\mu}}I_{-\mu}\Big(\frac{2\sqrt{q}\mu}{\mathsf{v}}\Big)=\frac{1}{\sqrt{q}^{\mu}}I_{\mu}\Big(\frac{2\sqrt{q}\mu}{\mathsf{v}}\Big).$
For $D_{\mu},\ \mu\geq-1$, let $n=m+\mu+1$,
$D_{\mu}$ $=\sum_{m=0}^{\infty}\frac{(\frac{\mu}{\mathsf{v}})^{m+\mu+1}(\frac{\mu q}{\mathsf{v}})^{m}}{(m+\mu+1)!m!}$
$=\frac{1}{\sqrt{q}^{\mu+1}}\sum_{m=0}^{\infty}\frac{1}{(m+\mu+1)!m!}\Big(\frac{\mu\sqrt{q}}{\mathsf{v}}\Big)^{2m+\mu+1}$
$=\frac{1}{\sqrt{q}^{\mu+1}}I_{\mu+1}\Big(\frac{2\sqrt{q}\mu}{\mathsf{v}}\Big).$
For $D_{\mu},\ \mu<-1$, let $m=n-\mu-1$,
$D_{\mu}$ $=\sum_{n=0}^{\infty}\frac{(\frac{\mu}{\mathsf{v}})^{n}(\frac{\mu q}{\mathsf{v}})^{n-\mu-1}}{n!(n-\mu-1)!}$
$=\frac{1}{\sqrt{q}^{\mu+1}}\sum_{n=0}^{\infty}\frac{1}{n!(n-\mu-1)!}\Big(\frac{\mu\sqrt{q}}{\mathsf{v}}\Big)^{2n-\mu-1}$
$=\frac{1}{\sqrt{q}^{\mu+1}}I_{-\mu-1}\Big(\frac{2\sqrt{q}\mu}{\mathsf{v}}\Big)=\frac{1}{\sqrt{q}^{\mu+1}}I_{\mu+1}\Big(\frac{2\sqrt{q}\mu}{\mathsf{v}}\Big).$